### v1

# Optimal low-rank approximations of posteriors for linear Gaussian inverse problems on Hilbert spaces

Abstract
For linear inverse problems with Gaussian priors and observation noise, the posterior is Gaussian, with mean and covariance determined by the conditioning formula. We analyse measure approximation problems of finding the best approximation to the posterior in a family of Gaussians with approximate covariance or approximate mean, for Hilbert parameter spaces and finite-dimensional observations. We quantify the error of the approximating Gaussian either with the Kullback–Leibler divergence or the family of Rényi divergences. Using the Feldman–Hajek theorem and recent results on reduced-rank operator approximations, we identify optimal solutions to these measure approximation problems. Our results extend those of Spantini et al. (SIAM J. Sci. Comput. 2015) to Hilbertian parameter spaces. In addition, our results show that the posterior differs from the prior only on a subspace of dimension equal to the rank of the Hessian of the negative log-likelihood, and that this subspace is a subspace of the Cameron–Martin space of the prior.
Keywords: nonparametric linear Bayesian inverse problems, Gaussian measures, low-rank operator approximation, generalised operator eigendecomposition MSC codes: 28C20, 47A58, 60G15, 62F15, 62G05
1 Introduction
The class of Bayesian inverse problems with linear forward models and Gaussian priors plays a special role in the context of Bayesian statistical inference. For example, this class of linear Gaussian inverse problems appears naturally in the Laplace approximation of posteriors for nonlinear statistical inverse problems, and the classical Kalman filter can be understood as an iterative solution method for a sequence of linear Gaussian inverse problems. A particularly attractive feature of the class of linear Gaussian inverse problems is the availability of a closed-form solution, in the case where the parameter space is a separable Hilbert space. In this case, given a linear forward model $G$ with codomain $\mathbb{R}^{n}$ for some $n\in\mathbb{N}$, a realisation $y$ of the $\mathbb{R}^{n}$-valued data random variable
$Y=GX+\zeta,\quad\zeta\sim\mathcal{N}(0,\mathcal{C}_{\textup{obs}}),$
and given a Gaussian prior $\mu_{\textup{pr}}=\mathcal{N}(m_{\textup{pr}},\mathcal{C}_{\textup{pr}})$ for the unknown parameter $X$, the solution $\mu_{\textup{pos}}$ to the Bayesian inverse problem is a Gaussian measure $\mathcal{N}(m_{\textup{pos}},\mathcal{C}_{\textup{pos}})$. The posterior mean $m_{\textup{pos}}$ and the posterior covariance $\mathcal{C}_{\textup{pos}}$ can be computed explicitly:
$m_{\textup{pos}}=m_{\textup{pr}}+\mathcal{C}_{\textup{pr}}G^{*}\mathcal{C}_{\textup{obs}}^{-1}(y-Gm_{\textup{pr}}),\quad\mathcal{C}_{\textup{pos}}=\mathcal{C}_{\textup{pr}}-\mathcal{C}_{\textup{pr}}G^{*}(\mathcal{C}_{\textup{obs}}+G\mathcal{C}_{\textup{pr}}G^{*})^{-1}G\mathcal{C}_{\textup{pr}},$
see e.g. [31, Example 6.23].
The availability of closed-form solutions to linear Gaussian inverse problems endows these problems with structure that makes them interesting objects to study in the context of measure approximation problems. Measure approximation problems have become ubiquitous in modern statistical inference, often because one cannot sample exactly from the probability measure of interest, e.g. for computational cost reasons, or because one has only partial information about the measure of interest. In the context of Bayesian inverse problems, we can also consider measure approximation problems as a way to analyse the Bayesian prior-to-posterior update.
Computational studies of Bayesian inverse problems on high- but finite-dimensional parameter spaces show that the data is often ‘informative’ — i.e., that the posterior differs from the prior — only on a subspace of much lower dimension; see e.g. [14]. Since the posterior is obtained by reweighting the prior by the likelihood, this data-informed subspace is determined by how the concentration of the likelihood interacts with the concentration of prior. In the case of linear Gaussian inverse problems, the concentration of the likelihood and the concentration of the prior are described by the eigenpairs of the Hessian of the negative log-likelihood and the eigenpairs of the prior precision. These ideas were used in [14] to identify low-rank approximations of the posterior covariance matrix. In [30], the optimality of these posterior covariance approximations with respect to a family of spectral loss functions was shown. In particular, the leading generalised eigenvectors of the Hessian-prior precision matrix pencil build a hierarchy of nested low-dimensional subspaces on which the posterior differs from the prior.
The primary goal of this paper is to develop a mathematical analysis of the Bayesian prior-to-posterior update for linear inverse problems where the parameter space is a separable Hilbert space and the observation space is finite-dimensional, using low-rank measure approximation problems to identify optimal Gaussian approximations to the true posterior. The secondary goal is to extend the results of [30] that were developed for finite-dimensional parameter spaces, to the case where the parameter space is possibly infinite-dimensional.
The first main contribution of this paper is Lemma 3.6. In this result, we identify the generalised eigenpairs of three operator pencils: the first pencil is defined by the Hessian of the negative log-likelihood and the prior precision, the second is defined by the same Hessian and the posterior precision, and the third is defined by the posterior covariance and prior covariance. We use the generalised eigenpairs for the third operator pencil to show that the subspace on which the posterior differs from the prior has dimension equal to the rank of the Hessian of the negative log-likelihood, and that this subspace is spanned by vectors that belong to the Cameron–Martin space of the prior.
The second main contribution of this paper is to formulate and solve some families of Gaussian measure approximation problems, where we measure the error of the approximating Gaussian using either the Kullback–Leibler divergence or the family of Rényi divergences. The first family of measure approximation problems concerns approximating the Gaussian posterior by Gaussians with the true posterior mean but approximate posterior covariances that have the structure of being low-rank negative updates of the prior covariance. We identify optimal solutions for this family in Theorem 4.18. The second family concerns approximating the Gaussian posterior by Gaussians with the true posterior covariance but approximate posterior means. We consider two different types of approximate posterior means and identify optimal solutions for these two types in Theorem 5.6 and Theorem 5.7.
The third main contribution of this paper is to consider the family of measure approximation problems where both the posterior mean and posterior covariance are jointly approximated. In Section 6, we exploit the Pythagorean structure of the formula for the Kullback–Leibler divergence between Gaussians, to show that the optimal solutions for the covariance approximation and mean approximation problems from Theorems 4.18, 5.6 and 5.7 yield optimal solutions for the joint approximation problems.
1.1 Related literature
The problem of low-rank approximations of covariances of posteriors for linear Gaussian inverse problems posed on finite-dimensional parameter spaces was studied in [14]. In particular, [14, Equation (5)] presents a formula for a low-rank approximation of the posterior covariance that exploits spectral decay in the Hessian of the negative log-likelihood, and [14, Equation (4)] indicates that the error of this low-rank approximation is related to the tail of the spectrum of the so-called ‘prior-preconditioned Hessian’ of the negative log-likelihood.
In [30], a precise formulation of the low-rank posterior covariance approximation problem is given and rigorously analysed, for linear Gaussian inverse problems on finite-dimensional parameter spaces. The low-rank approximation proposed in [14] is shown to be an optimal solution for a family of spectral loss functions that include as special cases the Kullback–Leibler divergence and Hellinger distance between Gaussians with the same mean but different covariances. In addition, the low-rank posterior mean approximation problem was formulated, and two types of posterior mean approximations were studied. The Bayes risk after averaging over realisations of the data was used to quantify the approximation error. For this loss function, optimal solutions to the posterior mean approximation problems were obtained. This approach was further developed for goal-oriented linear Gaussian inverse problems in [29]. Dimension reduction methods for linear Gaussian inverse problems using projections of the data were studied using generalised eigenvalue problems in [15]. The Kullback–Leibler divergence and mutual information were used to quantify the error of the approximating measures.
Dimension reduction for Bayesian inverse problems with possibly nonlinear forward models and non-Gaussian priors appears to have been first analysed in [32]. Joint dimension reduction of parameter and data was studied in [3], also for possibly nonlinear forward models and non-Gaussian priors. The results of [32] were further improved in [21, 20], which derived error bounds in terms of Amari $\alpha$-divergences
$D_{\alpha}(\nu\|\mu)\coloneqq\frac{1}{\alpha(\alpha-1)}\biggr{(}\int\left(\frac{\mathrm{d}\nu}{\mathrm{d}\mu}\right)^{\alpha}\mathrm{d}\mu-1\biggr{)},$
for probability measures $\nu$ and $\mu$ such that $\nu\ll\mu$ and $0<\alpha\leq 1$; see [21, Equation (7)]. The above-cited works consider only the setting of finite-dimensional parameter spaces, and do not consider infinite-dimensional parameter spaces. While [30] provides explicit formulas for the approximation errors, [32, 3, 21, 20] provide only error bounds.
In [25], Kullback–Leibler approximation of probability measures on infinite-dimensional Polish spaces using Gaussians was studied from the calculus of variations perspective. The main results of this work concern existence of minimisers and convergence of a proposed minimisation scheme for identifying the best approximation in a class of approximating Gaussian measures. In our setting, the posterior is already Gaussian, and the approximation classes we consider differ from those in [25].
In [1, Section 3], importance sampling for linear Gaussian inverse problems posed on separable Hilbert spaces is considered. The main result is to identify two types of intrinsic dimension, such that if both dimensions are finite, then absolute continuity of the posterior with respect to the prior holds, and thus importance sampling may be possible. The focus of our work is not to determine whether importance sampling is possible, but rather to identify low-dimensional structure in the Bayesian prior-to-posterior update.
1.2 Outline
We introduce key notation in Section 1.3 below. In Section 2, we begin by formulating the posterior covariance and posterior mean approximation classes that define our measure approximation problems. In Section 3 we recall the Feldman–Hajek theorem and formulas for the Kullback–Leibler divergence and Rényi divergence of two equivalent Gaussians. We state the first main result of this paper, Lemma 3.6, which identifies the generalised eigenpairs of the three operator pencils mentioned in the introduction and identifies the finite-dimensional subspace on which the posterior differs from the prior. In Section 4, we consider measure approximation problems where the posterior covariance is approximated and identify solutions in Theorem 4.18. In Section 5, we consider the analogous problem for two families of posterior mean approximations, and identify the corresponding solutions in Theorem 5.6 and Theorem 5.7. We then consider joint approximation of the posterior mean and covariance in Section 6. Auxiliary results aare presented in Appendix A, and proofs of the results in this work can be found in Appendix B.
1.3 Notation
[…]
We denote equivalence of two measures $\mu$ and $\nu$ by $\mu\sim\nu$. We write $X\sim\mu$ to denote the fact that the law of a random variable $X$ is $\mu$, and $\mathcal{N}(m,\mathcal{C})$ to denote the Gaussian distribution with mean $m$ and covariance $\mathcal{C}$. The ‘precision’ of $\mathcal{N}(m,\mathcal{C})$ is $\mathcal{C}^{-1}$. We denote the support of a measure $\mu$ by $\operatorname{supp}\mu$. For $p>0$ and $I$ a non-empty interval in $\mathbb{R}$, $\ell^{p}(I)$ denotes the space of $p$-summable sequences, i.e. $\ell^{p}(I)=\{(x_{i})_{i\in\mathbb{N}}\subset I\ :\ \sum_{i\in\mathbb{N}}\lvert x_{i}\rvert^{p}<\infty\}$. We write ‘$a\leftarrow b$’ to denote the replacement of $a$ with $b$. We shall abbreviate ‘orthonormal basis’ by ‘ONB’.
2 Low-rank approximations to linear Gaussian inverse problems
Let $\mathcal{H}$ be a separable Hilbert space over $\mathbb{R}$ of dimension $\dim{H}\leq\infty$. Consider the observation model defined by a continuous linear forward model $G\in\mathcal{B}(\mathcal{H},\mathbb{R}^{n})$ and additive Gaussian observation error
$Y=Gx^{\dagger}+\zeta,\quad\zeta\sim\mathcal{N}(0,\mathcal{C}_{\textup{obs}}).$ (1)
The covariance $\mathcal{C}_{\textup{obs}}\in B(\mathbb{R}^{n})_{\mathbb{R}}$ of the observation noise $\zeta$ is positive, and from a frequentist nonparametric perspective, $x^{\dagger}\in\mathcal{H}$ is the unknown true data-generating parameter to recover after observing a realisation of $Y$. By the Gaussian assumption on the noise, it follows that for any fixed $x\in\mathcal{H}$, the likelihood of observing $y$ is proportional to $\exp(-\frac{1}{2}\lVert\mathcal{C}_{\textup{obs}}^{-1}(y-Gx)\rVert^{2})$. The Hessian of the negative log-likelihood is
$H=G^{*}\mathcal{C}_{\textup{obs}}^{-1}G\in\mathcal{B}_{00,n}(\mathcal{H})_{\mathbb{R}}.$ (2)
It follows from $H={G}^{*}\mathcal{C}_{\textup{obs}}^{-1/2}(G^{*}\mathcal{C}_{\textup{obs}}^{-1/2})^{*}$ that $H$ is self-adjoint and nonnegative.
We adopt the Bayesian perspective to the problem of inferring $x^{\dagger}$ given the observation $y$ of $Y$, by modeling the unknown $x^{\dagger}$ with an $\mathcal{H}$-valued random variable $X$. Its distribution, the prior distribution, is taken to be a nondegenerate Gaussian measure $\mu_{\textup{pr}}=\mathcal{N}(0,\mathcal{C}_{\textup{pr}})$ on $\mathcal{H}$ and we assume that $X$ and $\zeta$ are independent. As the covariance of a Gaussian measure, $\mathcal{C}_{\textup{pr}}$ lies in $L_{1}(\mathcal{H})_{\mathbb{R}}$ and $\mathcal{C}_{\textup{pr}}\geq 0$, hence $\mathcal{C}_{\textup{pr}}$ has a unique nonnegative square root $\mathcal{C}_{\textup{pr}}^{1/2}\in L_{2}(\mathcal{H})_{\mathbb{R}}$. Nondegeneracy of $\mu_{\textup{pr}}$ implies that $\operatorname{supp}(\mu_{\textup{pr}})=\mathcal{H}$, see e.g. [4, Definition 3.6.2] and that $\mathcal{C}_{\textup{pr}}$ is injective, hence $\mathcal{C}_{\textup{pr}}>0$ and $\mathcal{C}_{\textup{pr}}^{1/2}>0$, see Lemma A.19. Hence the inverse $\mathcal{C}_{\textup{pr}}^{-1}$ is a well-defined bijection $\operatorname{ran}{C}_{\textup{pr}}\rightarrow\mathcal{H}$ which is self-adjoint, c.f. Definition A.12 and Lemma A.16(ii), and is unbounded if $\dim{H}=\infty$. The Cameron–Martin space of $\mu_{\textup{pr}}$ is the Hilbert space $(\operatorname{ran}{C}_{\textup{pr}}^{1/2},\lVert\cdot\rVert_{\mathcal{C}_{\textup{pr}}^{-1}})$, see e.g. [4, p. 293], where the Cameron–Martin norm of an element $h\in\operatorname{ran}{\mathcal{C}_{\textup{pr}}^{1/2}}$ is defined by $\lVert h\rVert_{\mathcal{C}_{\textup{pr}}^{-1}}\coloneqq\lVert\mathcal{C}_{\textup{pr}}^{-1/2}h\rVert$. As $\mathcal{C}_{\textup{pr}}$ is injective and compact, $\operatorname{ran}{C}_{\textup{pr}}^{1/2}$ is dense in $\mathcal{H}$ and if $\dim{H}=\infty$, then $\operatorname{ran}{C}_{\textup{pr}}^{1/2}$ is a proper dense subspace of $\mathcal{H}$.
Given a realisation $y$ of the random variable $Y$ defined by the observation model, the posterior law $\mu_{\textup{pos}}=\mu_{\textup{pos}}(y)$ of $X$ given $Y=y$ is the Gaussian measure $\mathcal{N}(m_{\textup{pos}},\mathcal{C}_{\textup{pos}})$, where
$m_{\textup{pos}}=m_{\textup{pos}}(y)$ $=\mathcal{C}_{\textup{pos}}G^{*}\mathcal{C}_{\textup{obs}}^{-1}y\in\operatorname{ran}{\mathcal{C}_{\textup{pos}}},$ (3a)
$\mathcal{C}_{\textup{pos}}$ $=\mathcal{C}_{\textup{pr}}-\mathcal{C}_{\textup{pr}}G^{*}(\mathcal{C}_{\textup{obs}}+G\mathcal{C}_{\textup{pr}}G^{*})^{-1}G\mathcal{C}_{\textup{pr}},$ (3b)
$\mathcal{C}_{\textup{pos}}^{-1}$ $=\mathcal{C}_{\textup{pr}}^{-1}+G^{*}\mathcal{C}_{\textup{obs}}^{-1}G=\mathcal{C}_{\textup{pr}}^{-1}+H,$ (3c)
see e.g. [31, Example 6.23]. Equation (3c) should be understood to hold on $\operatorname{ran}{C}_{\textup{pos}}$. While all nondegenerate Gaussians are equivalent in finite-dimensionsional setting, this is no longer true in infinite-dimenional setting, where in fact it holds that nondegenerate Gaussians that are not equivalent must be mutually singular. By [31, Section 6.3], $\mu_{\textup{pos}}$ and $\mu_{\textup{pr}}$ are equivalent. We shall construct approximations to $\mu_{\textup{pos}}$ that are equivalent to $\mu_{\textup{pos}}$.
The equations in (3) motivate certain Gaussian approximations of $\mu_{\textup{pos}}$ that, as we shall see, retain equivalence to $\mu_{\textup{pos}}$. By (3b), $\mathcal{C}_{\textup{pos}}$ is an update of $\mathcal{C}_{\textup{pr}}$ by a nonpositive self-adjoint operator $-\mathcal{C}_{\textup{pr}}G^{*}(\mathcal{C}_{\textup{obs}}+G\mathcal{C}_{\textup{pr}}G^{*})^{-1}G\mathcal{C}_{\textup{pr}}$. The range of this update is contained in $\operatorname{ran}{\mathcal{C}_{\textup{pr}}}$ and the rank of this update is at most $n$ since $G\in\mathcal{B}(\mathcal{H},\mathbb{R}^{r})$. For $r\in\mathbb{N}$, this motivates the rank-constained approximation of $\mathcal{C}_{\textup{pos}}$ by updating $\mathcal{C}_{\textup{pr}}$ by nonpositive self-adjoint operators of the form $KK^{*}$, for $K\in\mathcal{B}(\mathbb{R}^{r},\mathcal{H})$ with $\operatorname{ran}{K}\subset\operatorname{ran}{\mathcal{C}_{\textup{pr}}}$ and $\mathcal{C}_{\textup{pr}}-KK^{*}\in\mathcal{E}$. That is, we consider
$\mathscr{C}_{r}\coloneqq\left\{\mathcal{C}_{\textup{pr}}-KK^{*}>0:\ K\in\mathcal{B}(\mathbb{R}^{r},\mathcal{H}),\operatorname{ran}{K}\subset\operatorname{ran}{\mathcal{C}_{\textup{pr}}}\right\},\quad r\in\mathbb{N},$ (4)
and we have $\mathcal{C}_{\textup{pos}}\in\mathscr{Cj_{r}}$ for some $r\leq n$.
Alternatively, we can consider approximations of $\mu_{\textup{pos}}$ by constructing rank-constrained updates of the prior precision $\mathcal{C}_{\textup{pr}}^{-1}$. By (3c), $\mathcal{C}_{\textup{pr}}^{-1}$ is an update of $\mathcal{C}_{\textup{pr}}^{-1}$ by the Hessian $H$, which is a self-adjoint, nonnegative, and has rank at most $n$. For $r\in\mathbb{N}$, we can therefore consider the class of approximations of $\mathcal{C}_{\textup{pos}}^{-1}$ of the form $\mathcal{C}_{\textup{pr}}^{-1}+UU^{*}$, for $U\in\mathcal{B}(\mathbb{R}^{r},\mathcal{H})$. That is, we consider
$\mathscr{P}_{r}\coloneqq\left\{\mathcal{C}_{\textup{pr}}^{-1}+UU^{*}:\ U\in\mathcal{B}(\mathbb{R}^{r},\mathcal{H})\right\},\quad r\in\mathbb{N},$ (5)
and hence $\mathcal{C}_{\textup{pos}}^{-1}\in\mathscr{P}_{r}$ for some $r\leq n$, by the fact that $H\in\mathcal{B}_{00,n}(\mathcal{H})$.
We note that every operator $SS^{*}$ for $S\in\mathcal{B}(\mathbb{R}^{r},\mathcal{H})$ is a positive, self-adjoint operator with rank at most $r$, and that every nonnegative operator $T\in\mathcal{B}_{00,r}(\mathcal{H})_{\mathbb{R}}$ can be written in this way. Therefore, we could write the above approximations as $\mathcal{C}_{\textup{pr}}-T$ or $\mathcal{C}_{\textup{pr}}^{-1}+T$ for nonnegative $T\in\mathcal{B}_{00,r}(\mathcal{H})$, such that $\mathcal{C}_{\textup{pr}}-T$ is positive and maps into $\operatorname{ran}{\mathcal{C}_{\textup{pr}}}$. However, the set of nonnegative elements of $\mathcal{B}_{00,r}(\mathcal{H})$ is not convex, since rank is not preserved by convex combinations. By replacing $T$ by $SS^{*}$, we avoid formulating an optimisation problem over a nonconvex set. Indeed, $\mathcal{B}(\mathbb{R}^{r},\mathcal{H})$ is not only convex but is also a Banach space.
For approximations of the posterior mean, we observe from (3a) that the posterior mean is the result of left-multiplying the data $y$ by an operator. We shall consider approximations of the posterior mean that retain equivalence with the exact posterior. This motivates the following classes of operators:
$\mathscr{M}^{(1)}_{r}\coloneqq$$\{(\mathcal{C}_{\textup{pr}}-B)G^{*}\mathcal{C}_{\textup{obs}}^{-1}:\ B\in\mathcal{B}_{00,r}(\mathcal{H}),\ N((\mathcal{C}_{\textup{pr}}-B)G^{*}\mathcal{C}_{\textup{obs}}^{-1}y,\mathcal{C}_{\textup{pos}})\sim\mu_{\textup{pos}}\},$ (6a)
$\mathscr{M}^{(2)}_{r}\coloneqq$$\{A\in\mathcal{B}_{00,r}(\mathbb{R}^{n},\mathcal{H}):\ N((\mathcal{C}_{\textup{pr}}-B)G^{*}\mathcal{C}_{\textup{obs}}^{-1}y,\mathcal{C}_{\textup{pos}})\sim\mu_{\textup{pos}}\}.$ (6b)
The class in (6a) takes properties of the posterior covariance that are implied by (3b) into account. In contrast, the class in (6b) ignores this structure and takes only into account that the resulting approximate posterior approximation is equivalent to the exact posterior. By (3a), it follows that there exists $r\leq n$ such that $m_{\textup{pos}}\in\mathscr{M}^{(1)}_{r}\cap\mathscr{M}^{(2)}_{r}$. In fact, if $r\geq\operatorname{rank}\left({G^{*}}\right)=\operatorname{rank}\left({G}\right)$, then $(\mathcal{C}_{\textup{pr}}-B)G^{*}\mathcal{C}_{\textup{obs}}^{-1}\in\mathcal{B}_{00,r}(\mathbb{R}^{n},\mathcal{H})$ for every $B\in\mathcal{B}_{00,r}(\mathcal{H})_{\mathbb{R}}$. The classes $\mathscr{C}_{r}$, $\mathscr{P}_{r}$, $\mathscr{M_{r}}^{(1)}$ and $\mathscr{M}_{r}^{(2)}$ generalise those considered in [30]. It is a natural to search for low-rank approximations of the objects in (3), where ‘low-rank’ refers to the fact that we consider approximations in the classes $\mathscr{C}_{r}$, $\mathscr{P}_{r}$, $\mathscr{M}_{r}^{(1)}$ and $\mathscr{M}_{r}^{(2)}$ for $r<n$ respectively.
3 Equivalence and Divergences between Gaussian measures
Since our approximation problems are formulated in the context of statistical inverse problems, and since absolute continuity of the posterior with respect to the prior is important for statistical inference, we ensure that the approximate posteriors we consider are equivalent to $\mu_{\textup{pos}}$. In Section 3.1, we recall the Feldman–Hajek theorem which gives necessary and sufficient conditions for Gaussian measures to be equivalent, and apply this theorem to the setting described in Section 2. Then, in Section 3.2, we consider certain divergences between equivalent Gaussian measures, which we can use to measure approximation quality of our constructed low-rank posterior approximations.
3.1 Equivalence between Gaussian measures
Given a fixed nondegenerate reference Gaussian measure, the set of equivalent Gaussian measures is described by the Feldman–Hajek theorem, see e.g. [4, Corollary 6.4.11] or [11, Theorem 2.25].
Theorem 3.1 (Feldman–Hajek).
Let $\mathcal{H}$ be a Hilbert space and $\mu=N(m_{1},\mathcal{C}_{1})$ and $\nu=(m_{2},\mathcal{C}_{2})$ be Gaussian measures on $\mathcal{H}$. Then $\mu$ and $\nu$ are singular or equivalent, and $\mu$ and $\nu$ are equivalent if and only if the following conditions hold:
  1. (i)
    $\operatorname{ran}{\mathcal{C}_{1}^{1/2}}=\operatorname{ran}{\mathcal{C}_{2}^{1/2}}$,
  2. (ii)
    $m_{1}-m_{2}\in\operatorname{ran}{\mathcal{C}_{1}^{1/2}}$ and,
  3. (iii)
    $(\mathcal{C}_{1}^{-1/2}\mathcal{C}_{2}^{1/2})(\mathcal{C}_{1}^{-1/2}\mathcal{C}_{2}^{1/2})^{*}-I\in L_{2}(\mathcal{H})$.
The operator appearing in Theorem 3.1(iii) quantifies the amount of similarity between Gaussian measures. If it does not have square summable eigenvalues, then the Gaussian measures are mutually singular. In the other extreme, if the Gaussian measures are equal, then this operator is equal to 0 and the squared eigenvalues sum to 0.
Let us define
$\mathcal{E}\coloneqq\{\mathcal{C}\in L_{1}(\mathcal{H})_{\mathbb{R}}:\ \mathcal{N}(m_{\textup{pos}},\mathcal{C})\sim\mu_{\textup{pos}}\}.$ (7)
That is, $\mathcal{E}$ contains those covariances $\mathcal{C}$ such that $\mathcal{N}(m_{\textup{pos}},\mathcal{C})$ is equivalent to $\mu_{\textup{pos}}$. Since $\mu_{\textup{pos}}$ and $\mu_{\textup{pr}}$ are equivalent, we have $\mathcal{C}_{\textup{pr}}\in\mathcal{E}$.
In order to characterise the set $\mathcal{E}$ in (7), we introduce the following definition, which is closely related to item (iii) of Theorem 3.1. This definition appears in [4, Section 6.3].
Definition 3.2.
If $A\in\mathcal{B}(H)$ is invertible and $AA^{*}-I\in L_{2}(\mathcal{H})$, then we say that $A$ satisfies ‘property E’.
By [4, Lemma 6.3.1(ii)] the set of operators that satisfy property E is closed under taking inverses, adjoints and compositions. Furthermore, since $\mathcal{\mu}_{\textup{pos}}\sim\mu_{\textup{pr}}$, $\mathcal{C}_{\textup{pr}}^{-1/2}\mathcal{C}_{\textup{pos}}^{1/2}$ satisfies property E. One can now use Theorem 3.1 to describe the set $\mathcal{E}$ in (7) explicitly, see Lemma A.20.
=$\begin{split}\mathcal{E}&=\left\{C\in L_{1}(\mathcal{H})_{\mathbb{R}}:\ C\geq 0,\ \mathcal{C}^{-1/2}\mathcal{C}_{\textup{pos}}^{1/2}\text{ satisfies property E}\right\}\\ &=\left\{C\in L_{1}(\mathcal{H})_{\mathbb{R}}:\ C\geq 0,\ \mathcal{C}^{-1/2}\mathcal{C}_{\textup{pr}}^{1/2}\text{ satisfies property E}\right\}\end{split}$ (8)
For $\mathcal{C}_{1},\mathcal{C}_{2}\in\mathcal{E}$, we now define
$R(\mathcal{C}_{2}\|\mathcal{C}_{1})\coloneqq\mathcal{C}_{1}^{-1/2}\mathcal{C}_{2}^{1/2}(\mathcal{C}_{1}^{-1/2}\mathcal{C}_{2}^{1/2})^{*}-I.$ (9)
By Theorem 3.1(iii), $R(\mathcal{C}_{2}\|\mathcal{C}_{1})\in L_{2}(\mathcal{H})$. Since $R(\mathcal{C}_{2}\|\mathcal{C}_{1})$ is a self-adjoint compact operator, there exists an ONB of $\mathcal{H}$ that diagonalises $R(\mathcal{C}_{2}\|\mathcal{C}_{1})$, see Lemma A.6. We note that $R(\cdot\|\cdot)$ is in general not symmetric in its arguments. The result below will be used frequently in our analysis of low-rank approximations of the posterior covariance operator.
Lemma 3.3.
Let $\mathcal{C}_{1},\mathcal{C}_{2}$ be injective covariances of equivalent Gaussian measures. Then there exists a sequence $(\lambda_{i})_{i}\in\ell^{2}((-1,\infty))$ and ONBs $(w_{i})_{i}$ and $(v_{i})_{i}$ of $\mathcal{H}$ such that $v_{i}=\sqrt{1+\lambda_{i}}\mathcal{C}_{2}^{-1/2}\mathcal{C}_{1}^{1/2}w_{i}$ and the following hold
  1. (i)
    $\mathcal{C}_{1}^{-1/2}\mathcal{C}_{2}\mathcal{C}_{1}^{-1/2}-I\subset(\mathcal{C}_{1}^{-1/2}\mathcal{C}_{2}^{1/2})(\mathcal{C}_{1}^{-1/2}\mathcal{C}_{2}^{1/2})^{*}-I=\sum_{i=1}\lambda_{i}w_{i}\otimes w_{i}\in L_{2}(\mathcal{H})$,
  2. (ii)
    $\mathcal{C}_{2}^{1/2}\mathcal{C}_{1}^{-1}\mathcal{C}_{2}^{1/2}-I\subset(\mathcal{C}_{1}^{-1/2}\mathcal{C}_{2}^{1/2})^{*}(\mathcal{C}_{1}^{-1/2}\mathcal{C}_{2}^{1/2})-I=\sum_{i}\lambda_{i}v_{i}\otimes v_{i}\in L_{2}(\mathcal{H})$,
  3. (iii)
    $\mathcal{C}_{2}^{-1/2}\mathcal{C}_{1}\mathcal{C}_{2}^{-1/2}-I\subset(\mathcal{C}_{2}^{-1/2}\mathcal{C}_{1}^{1/2})(\mathcal{C}_{2}^{-1/2}\mathcal{C}_{1}^{1/2})^{*}-I=\sum_{i}\frac{-\lambda_{i}}{1+\lambda_{i}}v_{i}\otimes v_{i}\in L_{2}(\mathcal{H})$,
  4. (iv)
    $\mathcal{C}_{1}^{1/2}\mathcal{C}_{2}^{-1}\mathcal{C}_{1}^{1/2}-I\subset(\mathcal{C}_{2}^{-1/2}\mathcal{C}_{1}^{1/2})^{*}(\mathcal{C}_{2}^{-1/2}\mathcal{C}_{1}^{1/2})-I=\sum_{i}\frac{-\lambda_{i}}{1+\lambda_{i}}w_{i}\otimes w_{i}\in L_{2}(\mathcal{H})$,
where the leftmost operators are defined on $\operatorname{ran}{\mathcal{C}_{1}^{1/2}}=\operatorname{ran}{\mathcal{C}_{2}^{1/2}}$.
Lemma 3.3 shows that the operator in Theorem 3.1(iii) can equivalently described by the operators in items (ii), (iii) and (iv). The operators in Lemma 3.3 can be seen as generalisation of the notion of an operator pencil, which we formally define below.
Definition 3.4.
For possibly unbounded operators $T,S:\mathcal{H}\rightarrow\mathcal{H}$, the operator pencil $(T,S)$ is defined by the collection of operators $\{T-\lambda S,\ \lambda\in\mathbb{R}\}$. A ‘generalised eigenvalue’ of $(T,S)$ is a value $\lambda\in\mathbb{R}$ for which $T-\lambda S$ is not injective. For such $\lambda$ there exists a nonzero $v\in\operatorname{dom}{T}\cap\operatorname{dom}{S}$ such that $Tv=\lambda Sv$, which is called a ‘generalised eigenvector’, and we say that $(\lambda,v)$ is a ‘generalised eigenpair’ of $(T,S)$.
Remark 3.5 (Generalised eigenpairs).
If $w_{i}\in\operatorname{dom}{\mathcal{C}_{2}^{1/2}\mathcal{C}_{1}^{-1/2}}=\operatorname{ran}{\mathcal{C}_{1}^{1/2}}$ for some $i$, then the statement of item (i) implies $\mathcal{C}_{2}\mathcal{C}_{1}^{-1/2}w_{i}=(1+\lambda_{i})\mathcal{C}_{1}^{1/2}w_{i}$. In other words, $\mathcal{C}_{2}(\mathcal{C}_{1}^{-1/2}w_{i})=\mathcal{C}_{1}(\mathcal{C}_{1}^{-1/2}w_{i})$, showing that $(1+\lambda_{i},\mathcal{C}_{1}^{-1/2}w_{i})$ is a generalised eigenpair of the generalised operator pencil $(\mathcal{C}_{\textup{pos}},\mathcal{C}_{\textup{pr}})$. Furthermore, by the relation $v_{i}=\sqrt{1+\lambda_{i}}\mathcal{C}_{2}^{-1/2}\mathcal{C}_{1}^{1/2}w_{i}$, it follows that $v_{i}\in\operatorname{ran}{\mathcal{C}_{1}^{1/2}}=\operatorname{ran}{\mathcal{C}_{2}^{1/2}}$, and the relation item (ii) implies $\mathcal{C}_{1}^{-1}\mathcal{C}_{2}^{1/2}v_{i}=(1+\lambda_{i})\mathcal{C}_{2}^{-1}\mathcal{C}_{2}^{1/2}v_{i}$ showing that $(1+\lambda_{i},\mathcal{C}_{2}^{1/2}v_{i})=(1+\lambda_{i},\sqrt{1+\lambda_{i}}\mathcal{C}_{1}^{1/2}w_{i})$ is a generalised eigenpair of $(\mathcal{C}_{1}^{-1},\mathcal{C}_{2}^{-1})$. Thus, in the case $(w_{i})_{i}$ and $(v_{i})_{i}$ lie in the Cameron–Martin space $\operatorname{ran}{\mathcal{C}_{1}^{1/2}}=\operatorname{ran}{\mathcal{C}_{2}^{1/2}}$, items (i), (ii), (iii) and (iv) in Lemma 3.3 can be interpreted as statements on operator pencils. The statements in Lemma 3.3 are, however, more general.
Theorem 3.1 and Lemma 3.3 hold for any equivalent Gaussian measures. In the specific case of the linear Bayesian Inverse Problem (1), in which case the posterior precision is a finite-rank update $H$ of the prior by (3c), more can be said about the eigenvectors and eigenvalues given by Lemma 3.3 of the operators $R(\mathcal{C}_{\textup{pr}}\|\mathcal{C}_{\textup{pos}})$ and $R(\mathcal{C}_{\textup{pos}}\|\mathcal{C}_{\textup{pr}})$. We remind the reader of the definition of the Hessian ${H}$ in (2).
Lemma 3.6.
There exists an increasing sequence $(\lambda_{i})_{i}\in\ell^{2}((-1,0])$ consisting of exactly $\operatorname{rank}\left({H}\right)$ nonzero elements and ONBs $(w_{i})_{i}$ and $(v_{i})_{i}$ of $\mathcal{H}$ such that $w_{i},v_{i}\in\operatorname{ran}{\mathcal{C}_{\textup{pr}}^{1/2}}$ and $v_{i}=\sqrt{1+\lambda_{i}}\mathcal{C}_{\textup{pos}}^{-1/2}\mathcal{C}_{\textup{pr}}^{1/2}w_{i}$ for every $i\in\mathbb{N}$, and
$\mathcal{C}_{\textup{pr}}^{1/2}H\mathcal{C}_{\textup{pr}}^{1/2}$ $=(\mathcal{C}_{\textup{pos}}^{-1/2}\mathcal{C}_{\textup{pr}}^{1/2})^{*}(\mathcal{C}_{\textup{pos}}^{-1/2}\mathcal{C}_{\textup{pr}}^{1/2})-I=\sum_{i}\frac{-\lambda_{i}}{1+\lambda_{i}}w_{i}\otimes w_{i},$ (10a)
$\mathcal{C}_{\textup{pos}}^{1/2}H\mathcal{C}_{\textup{pos}}^{1/2}$ $=I-(\mathcal{C}_{\textup{pr}}^{-1/2}\mathcal{C}_{\textup{pos}}^{1/2})^{*}(\mathcal{C}_{\textup{pr}}^{-1/2}\mathcal{C}_{\textup{pos}}^{1/2})=\sum_{i}(-\lambda_{i})v_{i}\otimes v_{i},$ (10b)
$\mathcal{C}_{\textup{pos}}^{1/2}\mathcal{C}_{\textup{pr}}^{-1/2}w_{i}$ $=(1+\lambda_{i})\mathcal{C}_{\textup{pos}}^{-1/2}\mathcal{C}_{\textup{pr}}^{1/2}w_{i},\quad\forall i\in\mathbb{N}.$ (10c)
In particular, it holds that
$\frac{\textup{Var}_{X\sim\mu_{\textup{pos}}}(\langle X,\mathcal{C}_{\textup{pr}}^{-1/2}w_{i}\rangle)}{\textup{Var}_{X\sim\mu_{\textup{pr}}}(\langle X,\mathcal{C}_{\textup{pr}}^{-1/2}w_{i}\rangle)}=1+\lambda_{i}=\frac{1}{1+\tfrac{-\lambda_{i}}{1+\lambda_{i}}},\quad\forall i\in\mathbb{N}.$ (11)
Remark 3.7 (Variance reduction).
The significance of (11) is that the posterior variance along the span of $\mathcal{C}_{\textup{pr}}^{-1/2}w_{i}$ is smaller than the prior variance along the same subspace by a factor of $1+\tfrac{-\lambda_{i}}{1+\lambda_{i}}$, for $i\in\mathbb{N}$. This was observed in the finite-dimensional case by [30, Equation (3.4)]. Thus, Lemma 3.6 implies that finite-dimensional data can only inform finitely many directions in parameter space, in the sense that posterior variance is reduced relative to prior variance only over a finite-dimensional subspace. The directions $(\mathcal{C}_{\textup{pr}}^{-1/2}w_{i})_{i\leq\operatorname{rank}\left({H}\right)}$ are orthogonal with respect to the $\mathcal{C}_{\textup{pr}}$-weighted inner product $\langle h_{1},h_{2}\rangle_{\mathcal{C}_{\textup{pr}}}\coloneqq\langle\mathcal{C}_{\textup{pr}}h_{1},h_{2}\rangle$, and not the unweighted inner product of $\mathcal{H}$. If, for some $r<\operatorname{rank}\left({H}\right)$, there exists an $r$-dimensional subspace and the variance reduction on the complement of this subspace is sufficiently small, then this $r$-dimensional subspace is also called the ‘likelihood-informed-subspace’ in literature, see e.g. [8, 9, 10].
In Lemma 3.6, $w_{i},v_{i}\in\operatorname{ran}{\mathcal{C}_{\textup{pr}}^{1/2}}$ for all $i$, hence by Remark 3.5, the equations (10a) and (10b) can be interpreted as statements on operator pencils. More specifically, for any $i$, (10a) states that $(\frac{-\lambda_{i}}{1+\lambda_{i}},\mathcal{C}_{\textup{pr}}^{1/2}w_{i})$ is a generalised eigenpair of $(H,\mathcal{C}_{\textup{pr}}^{-1})$ and (10b) states that $(-\lambda_{i},\mathcal{C}_{\textup{pos}}^{1/2}v_{i})=(-\lambda_{i},\sqrt{1+\lambda_{i}}\mathcal{C}_{\textup{pr}}^{-1/2}w_{i})$ is a generalised eigenpair of $(H,\mathcal{C}_{\textup{pos}}^{-1})$. Furthermore, (10c) can be interpreted as a statement on the operator pencil $(\mathcal{C}_{\textup{pos}},\mathcal{C}_{\textup{pr}})$. The prior-preconditioned Hessian $\mathcal{C}_{\textup{pr}}^{1/2}H\mathcal{C}_{\textup{pr}}^{1/2}$ has been found to be the central object of study in the reduction of finite-dimensional linear Gaussian inverse problems, see [30, 8]. We observe that this operator is directly related to $R(\mathcal{C}_{\textup{pos}}\|\mathcal{C}_{\textup{pr}})$ via the equivalent characterisations given by Lemma 3.3 Items (i), (ii), (iii) and (iv) and hence to the function $R(\cdot\|\cdot)$ which quantifies the similarity of Gaussian measures by Theorem 3.1(iii).
3.2 Divergences between equivalent Gaussian measures
To measure the quality of an approximation $\tilde{\mathcal{C}}\in\mathcal{E}$ of $\mathcal{C}_{\textup{pos}}$ and $\tilde{m}_{\textup{pos}}$ to $m_{\textup{pos}}$, we shall use the Rényi divergences of order $\rho\in(0,1)$ and the forward an reverse Kullback-Leibler divergence (KL-divergence). The KL-divergence from a measure $\nu_{1}$ to a measure $\nu_{2}$ is defined as
$D_{\textup{KL}}(\nu_{2}\|\nu_{1})=\int_{\mathcal{H}}\log{\frac{\operatorname{d}\!{\nu}_{2}}{\operatorname{d}\!{\nu}_{1}}}\operatorname{d}\!{\nu}_{2}.$
If $\nu_{1}$ is an approximation of $\nu_{2}$, then the forward KL-divergence of the approximation is $D_{\textup{KL}}(\nu_{2}\|\nu_{1})$ and the reverse KL-divergence is $D_{\textup{KL}}(\nu_{1}\|\nu_{2})$. The Rényi divergence of order $\rho\in(0,1)$ is defined by, c.f. [22, Equation (130)],
$D_{\textup{Ren},\rho}(\nu_{2}\|\nu_{1})=-\frac{1}{\rho(1-\rho)}\log{\int_{\mathcal{H}}\left(\frac{\operatorname{d}\!{\nu}_{2}}{\operatorname{d}\!{\nu}_{1}}\right)^{\rho}\operatorname{d}\!{\nu}_{1}}$
In the Gaussian case, an explicit representation of these divergences holds, as shown in [22]. For this, we need a generalisation of the determinant to infinite-dimensional Hilbert spaces. Because in infinite dimensions the eigenvalues of a compact operator accumulate at 0, direct extension of the finite-dimensional definition of the determinant as the product of the eigenvalues to infinite dimensions will result in the determinant function being equal to the constant 00. A generalisation of the concept of the determinant for trace-class and Hilbert–Schmidt operators is given by the Fredholm determinant and Hilbert–Carleman determinant respectively. These are defined on respectively trace-classs and Hilbert–Schmidt perturbations of the identity, and are indicated with $\det(I+A)$, $A\in L_{1}(\mathcal{H})$, and respectively $\det_{2}{(I+A)}$, $A\in L_{2}(\mathcal{H})$. We refer to [27, Theorem 3.2, Theorem 6.2] or [28, Lemma 3.3, Theorem 9.2]. For $A\in L_{1}(\mathcal{H})$, we have $\det_{2}{(I+A)}=\det(I+A)\exp(-\operatorname{tr}\,\left({A}\right))$, and for $A\in L_{2}(\mathcal{H})$ the determinants are related via $\det_{2}(I+A)=\det(I+(I+A)\exp(-A)).$ By [27, Theorem 4.2, Theorem 6.2] or [28, Theorem 3.7, Theorem 9.2] for each $\mu\in\mathbb{R}$, we have the expression
$\det(1+\mu A)$ $=\prod_{i}(1+\mu\lambda_{i}),\quad A\in L_{1}(\mathcal{H})$ (12)
$\det_{2}(I+\mu A)$ $=\prod_{i}(1+\mu\lambda_{i})\exp{(-\mu\lambda_{i})},\quad A\in L_{2}(\mathcal{H}),$ (13)
where $(\lambda_{i})_{i}$ denotes the eigenvalue sequence of $A$. In the case that $\dim{\mathcal{H}}<\infty$ we note that $A-I\in L_{1}(\mathcal{H})$ and $\det(A)=\det(I+(A-I))=\prod_{i}\lambda_{i}$ and thus $\det(\cdot)$ indeed extends the finite-dimensional definition of the determinant. We can now formulate the explicit expression of the KL and Rényi divergences for equivalent Gaussian measures.
Theorem 3.8.
Let $m_{1},m_{2}\in\mathcal{H}$ and $\mathcal{C}_{1},\mathcal{C}_{2}\in L_{2}(\mathcal{H})_{\mathbb{R}}$ be nonnegative. If $m_{1}-m_{2}\in\operatorname{ran}{\mathcal{C}_{1}^{1/2}}$ and if $\mathcal{C}_{1}^{-1/2}\mathcal{C}_{2}^{1/2}$ satisfies property E, then
$D_{\textup{KL}}(\mathcal{N}(m_{2},\mathcal{C}_{2})\|\mathcal{N}(m_{1},\mathcal{C}_{1}))$ $\coloneqq\frac{1}{2}\left\lVert\mathcal{C}_{1}^{-1/2}(m_{2}-m_{1})\right\rVert^{2}-\frac{1}{2}\log\det_{2}(I+R(\mathcal{C}_{2}\|\mathcal{C}_{1})),$ (14a)
$\begin{split}D_{\textup{Ren},\rho}(\mathcal{N}(m_{2},\mathcal{C}_{2})\|\mathcal{N}(m_{1},\mathcal{C}_{1}))&\coloneqq\frac{1}{2}\left\lVert\bigr{(}\rho I+(1-\rho)(I+R(\mathcal{C}_{2}\|\mathcal{C}_{1}))\bigr{)}^{-1/2}\mathcal{C}_{1}^{-1/2}(m_{2}-m_{1})\right\rVert^{2}\\ &+\frac{\log\det\left[\bigl{(}I+R(\mathcal{C}_{2}\|\mathcal{C}_{1})\bigr{)}^{\rho-1}\bigl{(}\rho I+(1-\rho)(I+R(\mathcal{C}_{2}\|\mathcal{C}_{1}))\bigr{)}\right]}{2\rho(1-\rho)}.\end{split}$ (14b)
Furthermore,
$\lim_{\rho\rightarrow 1}D_{\textup{Ren},\rho}(\mathcal{N}(m_{2},\mathcal{C}_{2})\|\mathcal{N}(m_{1},\mathcal{C}_{1}))=D_{\textup{KL}}(\mathcal{N}(m_{2},\mathcal{C}_{2})\|\mathcal{N}(m_{1},\mathcal{C}_{1})),$
$\lim_{\rho\rightarrow 0}D_{\textup{Ren},\rho}(\mathcal{N}(m_{2},\mathcal{C}_{2})\|\mathcal{N}(m_{1},\mathcal{C}_{1}))=D_{\textup{KL}}(\mathcal{N}(m_{1},\mathcal{C}_{1})\|\mathcal{N}(m_{2},\mathcal{C}_{2})).$
4 Optimal approximations of covariance operators
In this section, we formulate a minimisation problem that aims at finding low-rank approximations of $\mathcal{C}_{\textup{pos}}$ that are optimal simultaneously with respect to all members of a class of spectral loss functions. This class includes the forward and reverse Kullback–Leibler divergence and the Rényi divergences as special cases.
4.1 Spectral loss functions and approximation classes
To measure the quality of a given approximation of the exact posterior covariance $\mathcal{C}_{\textup{pos}}$, we define a class of loss functions on $\mathcal{E}^{2}$ in the following way. Define
$\mathscr{F}$$\coloneqq\left\{f\in C^{1}((-1,\infty)):\ f(0)=0,\ xf^{\prime}(x)>0\text{ for }x\not=0,\ \lim_{x\rightarrow\infty}f(x)=\infty,\ f^{\prime}\text{ Lipschitz at }0\right\},$ (15a)
$\mathscr{L}$×$\coloneqq\left\{\mathcal{E}\times\mathcal{E}\ni(\mathcal{C}_{2},\mathcal{C}_{1})\mapsto\mathcal{L}_{f}(\mathcal{C}_{2}\|\mathcal{C}_{1})\coloneqq\sum_{i}f(\Lambda_{i}(R(\mathcal{C}_{2}\|\mathcal{C}_{1})))\right\}.$ (15b)
As we show in Lemma 4.1 below, the conditions $f\in C^{1}$ and $xf^{\prime}(x)>0$ for $x\not=0$ ensure that 00 is the unique minimiser of every $f\in\mathscr{F}$. The Lipshitz continuity of $f^{\prime}$ at 00 implies that every $\mathcal{L}\in\mathscr{L}$ is takes only finite values and, as will be shown later in Lemma 4.9, that $\mathcal{L}(\mathcal{C}_{\textup{pos}}\|\cdot)$ is differentiable on a suitable subspace of $\mathcal{E}$. The blowup at infinity condition is used to prove coercivity of $\mathcal{L}(\mathcal{C}_{\textup{pos}}\|\cdot)$ on a suitable subspace of $\mathcal{E}$, as we show in Lemma 4.16.
Lemma 4.1.
Let $f\in\mathscr{F}$. Then
  1. (i)
    $f^{\prime}(x)=0$ if and only if $x=0$, the image of $f$ lies in $[0,\infty)$ and for every $x\in\ell^{2}((-1,\infty))$ it holds that $\sum_{i}f(x_{i})<\infty$. In particular, the image of every $\mathcal{L}_{f}\in\mathscr{L}$ lies in $[0,\infty)$.
  2. (ii)
    Let $\alpha:(-1,\infty)\rightarrow(-1,\infty)$ be defined by $\alpha(x)=\frac{-x}{1+x}$. If $f\in\mathscr{F}$ satisfies $\lim_{x\rightarrow-1}f(x)=\infty$, then $f\circ\alpha\in\mathscr{F}$.
The class of loss functions considered in the finite-dimensional of [30, Definition 2.1] differs from the class (15) in two aspects. First, the domain of the former functions $f$ is $(0,\infty)$ and $f(0)$ need not be 0, while the domain of the latter is $(-1,\infty)$ and we require $f(0)=0$. That the natural class to consider involves the horizontal shift of $-1$ and the vertical shift becomes apparent as the fundamental object governing the losses is given by (9), which is a compact operator and therefore has an eigenvalue sequence accumulating at 0. Second, there is an additional Lipschitz condition in (15a), which implies that $\mathcal{L}$ is finite on $\mathcal{E}^{2}$ for every $\mathcal{L}\in\mathscr{L}$. Note that the condition that $f^{\prime}$ is Lipschitz continuous at 00 is not implied by the other conditions in (15a), as the function $f$ with $f^{\prime}(x)=sgn(x)\lvert x\rvert^{\alpha}$, $\alpha\in(0,1)$, and $f(0)=0$ shows. Here, $sgn(x)$ denotes the function assigning $1$ to $x\geq 0$ and $-1$ otherwise. This function satisfies all conditions of (15a) except the Lipschitz condition of $f^{\prime}$ at 0.
While restricted compared to [30], the loss class (15) is still rich enough to include the forward and reverse KL divergences and Réyni divergence between equivalent Gaussian measures with the same mean. This partially extends [30, Lemma 2.2], in which this is shown for the forward KL and Hellinger distance in the finite-dimensional setting.
Lemma 4.2.
Let $m\in\mathcal{H}$. Let $\mu_{i}=N(m_{i},\mathcal{C}_{i})$ be nondegenerate and $\mathcal{C}_{i}\in\mathcal{E}$ for $i=1,2$.
  1. (i)
    Let $f_{\textup{KL}}(x)\coloneqq\frac{1}{2}(x-\log(1+x))$. Then $f_{\textup{KL}}\in\mathscr{F}$ and
    $D_{\textup{KL}}(\mu_{2}\|\mu_{1})=-\frac{1}{2}\log\det_{2}(I+R(\mathcal{C}_{2}\|\mathcal{C}_{1}))=\mathcal{L}_{f_{\textup{KL}}}(\mathcal{C}_{2}\|\mathcal{C}_{1}).$
  2. (ii)
    Let $\rho\in(0,1)$ and $f_{\textup{Ren},\rho}(x)\coloneqq\frac{\rho-1}{2\rho(1-\rho)}\log(1+x)+\frac{1}{2\rho(1-\rho)}\log{(\rho+(1-\rho)(1+x))}.$ Then $f_{\textup{Ren},\rho}\in\mathscr{F}$ and
    $D_{\textup{Ren},\rho}(\mu_{2}\|\mu_{1})=\frac{\log\det\left[\bigl{(}I+R(\mathcal{C}_{2}\|\mathcal{C}_{1})\bigr{)}^{\rho-1}\bigl{(}\rho I+(1-\rho)(I+R(\mathcal{C}_{2}\|\mathcal{C}_{1}))\bigr{)}\right]}{2\rho(1-\rho)}=\mathcal{L}_{f_{\textup{Ren},\rho}}(\mathcal{C}_{2}\|\mathcal{C}_{1}).$
  3. (iii)
    For the reverse distances, we have $f_{\textup{KL}}\circ\alpha,f_{\textup{Ren},\rho}\circ\alpha\in\mathscr{F}$ with $\alpha(x)=\frac{-x}{1+x}$ on $(-1,\infty)$, and
    $D_{\textup{KL}}(\mu_{1}\|\mu_{2})=\mathcal{L}_{f_{\textup{KL}}\circ\alpha}(\mathcal{C}_{2}\|\mathcal{C}_{1}),\quad D_{\textup{Ren},\rho}(\mu_{1}\|\mu_{2})=\mathcal{L}_{f_{\textup{Ren},\rho}\circ\alpha}(\mathcal{C}_{2}\|\mathcal{C}_{1}).$
Given the approximation classes (4) and (5) and given the covariance loss functions in (15b), we can define the following low-rank approximation problem, for every $r\leq n$.
Problem 4.3 (Rank-$r$ nonpositive covariance updates).
Find $\mathcal{C}^{\textup{opt}}_{r}\in\mathscr{C}_{r}$ such that for every $\mathcal{L}\in\mathscr{L}$, $\mathcal{L}(\mathcal{C}_{\textup{pos}}\|\mathcal{C}^{\textup{opt}}_{r})=\min\{\mathcal{L}(\mathcal{C}_{\textup{pos}}\|\mathcal{C}):\ \mathcal{C}\in\mathscr{C}_{r}\}$.
Problem 4.4 (Inverses of rank-$r$ nonnegative precision updates).
Find $\mathcal{P}^{\textup{opt}}_{r}\in\mathscr{P}_{r}$ such that for every $\mathcal{L}\in\mathscr{L}$, $\mathcal{L}(\mathcal{C}_{\textup{pos}}\|(\mathcal{P}^{\textup{opt}}_{r})^{-1})=\min\{\mathcal{L}(\mathcal{C}_{\textup{pos}}\|\mathcal{P}^{-1}):\ \mathcal{P}\in\mathscr{P}_{r}\}$.
We note that even if an optimal covariance and precision can be found for given $\mathcal{L}\in\mathscr{L}$, it is not a priori clear that they are in fact independent of $\mathcal{L}$.
We also emphasise the following. Since the inverse of a self-adjoint positive matrix is again self-adjoint and positive, inverses of covariance operators are covariance operators in the case $\dim{\mathcal{H}}<\infty$. If $\dim{H}=\infty$, then this no longer holds because the trace-class property is not preserved under inversion in infinite dimensions. In fact, if $\dim{H}=\infty$ and $T$ is trace class, then $T^{-1}$ is an unbounded operator and its eigenvalue sequence is not summable since it is not bounded. In order to define a loss on precisions, as is done in the finite-dimensional case of [30, Corollary 3.1], one approach is to extend the domain of $\mathcal{L}\in\mathscr{L}$ to $\mathcal{E}^{2}\cup(\mathcal{E}^{-1})^{2}$ via $\mathcal{L}_{ext}(\mathcal{C}_{1}^{-1}\|\mathcal{C}_{2}^{-1})\coloneqq\mathcal{L}(\mathcal{C}_{1}\|\mathcal{C}_{2})$ and $\mathcal{L}_{ext}(\mathcal{C}_{1}\|\mathcal{C}_{2})\coloneqq\mathcal{L}(\mathcal{C}_{1}\|\mathcal{C}_{2})$ for $\mathcal{C}_{1},\mathcal{C}_{2}\in\mathcal{E}$. If $(\mathcal{C}_{1},\mathcal{C}_{2})\in\mathcal{E}^{2}\cap(\mathcal{E}^{-1})^{2}$, then $\mathcal{L}_{ext}(\mathcal{C}_{1}^{-1}\|\mathcal{C}_{2}^{-1})=\mathcal{L}(\mathcal{C}_{1}\|\mathcal{C}_{2})=\mathcal{L}_{ext}(\mathcal{C}_{1}\|\mathcal{C}_{2})$, showing that $\mathcal{L}_{ext}$ is well-defined. By the definitions (15), (9) and Lemma 3.3, $\mathcal{L}$ depends on $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$ only via the set of eigenvalues of the extensions of the densely defined and bounded operator $\mathcal{C}_{1}^{-1/2}\mathcal{C}_{2}\mathcal{C}_{1}^{-1/2}-I$. Lemma 3.3(i)-(ii) shows that eigenvalues of the latter operator remain unchanged when replacing $\mathcal{C}_{1}$ by $\mathcal{C}_{2}^{-1}$ and $\mathcal{C}_{2}$ by $\mathcal{C}_{1}^{-1}$. Thus, $\mathcal{L}_{ext}(\mathcal{C}_{2}^{-1}\|\mathcal{C}_{1}^{-1})=\mathcal{L}_{ext}(\mathcal{C}_{1}\|\mathcal{C}_{2})=\mathcal{L}(\mathcal{C}_{1}\|\mathcal{C}_{2})$. This equation firstly implies that $\mathcal{L}(\mathcal{C}_{\textup{pos}}\|\mathcal{P}^{-1})=\mathcal{L}_{ext}(\mathcal{P}\|\mathcal{C}_{\textup{pos}}^{-1})$ for $\mathcal{P}\in\mathscr{P}_{r}$, and we can reformulate 4.4 accordingly in terms of a loss $\mathcal{L}_{ext}$ on precisions. Secondly, it shows that there is no need to explicitly define a loss on precisions, as we can just use $\mathcal{L}$ on the corresponding covariances in reversed order instead.
A natural question is what the precisions corresponding to approximate covariances in $\mathscr{C}_{r}$ are, and what the covariances corresponding approximate precisions in $\mathscr{P}_{r}$ are. The next result shows that $\mathscr{C}_{r}$ and $\mathscr{P}_{r}$ are in fact each others inverses and that Problems 4.3 and 4.4 are in fact equivalent. Item (i) can be seen as a generalisation of the finite-dimensional result [30, Lemma A.2] to infinite-dimensional Hilbert spaces.
Lemma 4.5.
Let $r\in\mathbb{N}$.
  1. (i)
    For every $K\in\mathcal{B}(\mathbb{R}^{r},\mathcal{H})$ such that $\mathcal{C}_{\textup{pr}}-KK^{*}\in\mathscr{C}_{r}$, there exists $U\in\mathcal{B}(\mathbb{R}^{r},\mathcal{H})$ of the same rank as $K$, such that $(\mathcal{C}_{\textup{pr}}-KK^{*})^{-1}=\mathcal{C}_{\textup{pr}}^{-1}+UU^{*}\in\mathscr{P}_{r}$. The reverse correspondence also holds: for every $U\in\mathcal{B}(\mathbb{R}^{r},\mathcal{H})$ such that $\mathcal{C}_{\textup{pr}}^{-1}+UU^{*}\in\mathscr{P}_{r}$, there exists $K\in\mathcal{B}(\mathbb{R}^{r},\mathcal{H})$ of the same rank as $U$, such that $(\mathcal{C}_{\textup{pr}}^{-1}+UU^{*})^{-1}=\mathcal{C}_{\textup{pr}}-KK^{*}\in\mathscr{C}_{r}$. In particular, $\mathscr{C}_{r}^{-1}\coloneqq\{\mathcal{C}^{-1}:\ \mathcal{C}\in\mathscr{C}_{r}\}=\mathscr{P}_{r}$ and $\mathscr{P}_{r}^{-1}\coloneqq\{\mathcal{P}^{-1}:\ \mathcal{P}\in\mathscr{P}_{r}\}=\mathscr{C}_{r}$
  2. (ii)
    An approximation $\mathcal{C}^{\textup{opt}}_{r}\in\mathscr{C}_{r}$ solves 4.3 if and only if $(\mathcal{C}^{\textup{opt}}_{r})^{-1}\in\mathscr{P}_{r}$ solves 4.4. Furthermore, $\mathcal{L}(\mathcal{C}_{\textup{pos}}\|\mathcal{C}_{r}^{\textup{opt}})=\mathcal{L}(\mathcal{C}_{\textup{pos}}\|(\mathcal{P}_{r}^{\textup{opt}})^{-1})$.
Finally, we show that the approximation of $\mathcal{C}_{\textup{pos}}$ or $\mathcal{C}_{\textup{pos}}^{-1}$ by elements of $\mathscr{C}_{r}$ and $\mathscr{P}_{r}$ respectively results in approximate posterior measures that are equivalent to the exact posterior, when keeping the mean fixed. We also shows that 4.3 is well-defined in the sense that $\mathcal{L}(\mathcal{C}_{\textup{pos}}\|\cdot)<\infty$ is finite on $\mathscr{C}_{r}$ for any $\mathcal{L}\in\mathscr{L}$. By Lemma 4.5(ii), it follows that 4.4 is analogously well-defined. Recall the definition of $\mathcal{E}$ in (7).
Lemma 4.6.
It holds that $\mathcal{C}_{r}\subset\mathcal{E}$. Thus, for any $\mathcal{L}\in\mathscr{L}$, the map $\mathcal{C}\mapsto\mathcal{L}(\mathcal{C}_{\textup{pos}}\|\mathcal{C})$ is finite on $\mathscr{C}_{r}$ and the map $\mathcal{P}\mapsto\mathcal{L}(\mathcal{C}_{\textup{pos}}\|\mathcal{P}^{-1})$ is finite on $\mathscr{P}_{r}$.
4.2 Differentiability and minimisers of covariance loss function
In this section we solve 4.4, and by virtue of Lemma 4.5(ii) then also 4.3, by formulating it as a minimisation problem over the set of $U\in\mathcal{B}(\mathbb{R}^{r},\mathcal{H})$. We want to find the minimiser of the function
$J_{f}:\mathcal{B}(\mathbb{R}^{r},\mathcal{H})\rightarrow\mathbb{R},\quad U\mapsto L_{f}(\mathcal{C}_{\textup{pos}}\|(\mathcal{C}_{\textup{pr}}^{-1}+UU^{*})^{-1}),$
for any $f\in\mathscr{F}$, which we shall express as a composition of functions. This composition will facilitate the analysis of its differentiability and thereby the identification of its stationary points.
Let $\Lambda:L_{2}(\mathcal{H})_{\mathbb{R}}\rightarrow\ell^{2}(\mathbb{R})$ be a function that sends a self-adjoint Hilbert–Schmidt operator to its square summable eigenvalue sequence. We recall that $\ell^{2}(\mathbb{R})$ denotes the space of square integrable sequences. One ordering could be the one that labels the negative eigenvalues with the even integers and the positive eigenvalues with the odd integers, both ordered decreasingly in absolute value. A different choice is to order the eigenvalues in order of decreasing absolute value. We note that the choice of ordering of $\Lambda$ in two operators $T,S\in L_{2}(\mathcal{H})_{\mathbb{R}}$ is allowed to be different. The precise ordering of eigenvalues that $\Lambda$ assigns to an operator is not important, as we shall only look at compositions of $\Lambda$ with a function on $\ell^{2}(\mathbb{R})$ that is permutation invariant.
The function that we shall apply to the image of $\Lambda$ is
$F_{f}:\ell^{2}((-1,\infty))\rightarrow[0,\infty),\quad F_{f}((x_{i})_{i})=\sum_{i}f(x_{i}),$ (16)
which is well-defined by Lemma 4.1(i). Note that if $P\in\mathcal{B}(\ell_{2}((-1,\infty))$ is a permutation, i.e. $((Px)_{i})_{i}=(x_{\alpha(i)})_{i}$ for some bijection $\alpha$ of $\mathbb{N}$, then for every $x\in\ell^{2}((-1,\infty))$ it holds that $F_{f}(Px)=F_{f}(x)$.
We finally define the function $g:\mathcal{B}(\mathbb{R}^{r},\mathcal{H})\rightarrow\mathcal{B}(\mathcal{H})_{\mathbb{R}}$ by
$g(U)=\left((\mathcal{C}_{\textup{pr}}^{-1}+UU^{*})^{1/2}\mathcal{C}_{\textup{pos}}^{1/2}\right)^{*}(\mathcal{C}_{\textup{pr}}^{-1}+UU^{*})^{1/2}\mathcal{C}_{\textup{pos}}^{1/2}-I.$ (17)
Then the image of $g$ in fact consists of Hilbert–Schmidt operators and they can be diagonalised in the Cameron–Martin space, as is shown next. This result also motives the definition of $g$, as it shows that $g(U)$ has the same eigenvalues as $R(\mathcal{C}_{\textup{pos}}\|(\mathcal{C}_{\textup{pr}}^{-1}+UU^{*})^{-1})$. By eq. 10b, it also shows that $g(U)$ is a nonnegative, self–adjoint update of rank-$r$ of the negative of the posterior-preconditioned Hessian.
Lemma 4.7.
Let $r\in\mathbb{N}$, $U\in\mathcal{B}(\mathbb{R}^{r},\mathcal{H})$ and $g$ as in (17). Then $\operatorname{rank}\left({g(U)}\right)\leq r+\operatorname{rank}\left({H}\right)$ and there exists a sequence $(e_{i})_{i}\subset\operatorname{ran}{\mathcal{C}_{\textup{pr}}^{1/2}}$ which forms an ONB of $\mathcal{H}$ and a sequence $(\gamma_{i})_{i}\in\ell^{2}((-1,\infty))$ satisfying $g(U)=\sum_{i}\gamma_{i}e_{i}\otimes e_{i}$. Furthermore, it holds that
$g(U)=(\mathcal{C}_{\textup{pr}}^{-1/2}\mathcal{C}_{\textup{pos}}^{1/2})^{*}\mathcal{C}_{\textup{pr}}^{-1/2}\mathcal{C}_{\textup{pos}}^{1/2}-I+\mathcal{C}_{\textup{pos}}^{1/2}UU^{*}\mathcal{C}_{\textup{pos}}^{1/2}.$ (18)
Finally, the eigenvalues of $g(U)$ and $R(\mathcal{C}_{\textup{pos}}\|(\mathcal{C}_{\textup{pr}}^{-1}+UU^{*})^{-1})$ agree, counting multiplicities.
Remark 4.8.
In the finite-dimensional setting, $\operatorname{ran}{C}_{\textup{pr}}^{1/2}=\mathcal{H}$ and thus one can simply define $g(U)$ on all of $\mathcal{H}$ as
$\mathcal{C}_{\textup{pos}}^{1/2}(\mathcal{C}_{\textup{pr}}^{-1}+UU^{*})\mathcal{C}_{\textup{pos}}^{1/2}-I$ $=\mathcal{C}_{\textup{pos}}^{1/2}\mathcal{C}_{\textup{pr}}^{-1/2}(I+\mathcal{C}_{\textup{pr}}^{1/2}UU^{*}\mathcal{C}_{\textup{pr}}^{1/2})\mathcal{C}_{\textup{pr}}^{-1/2}\mathcal{C}_{\textup{pos}}^{1/2}-I.$
Of this operator, $(\gamma,v)\in\mathbb{R}_{+}\times\mathcal{H}$ is an eigenpair if and only if $w:=\mathcal{C}_{\textup{pr}}^{-1/2}\mathcal{C}_{\textup{pos}}^{1/2}v$ satisfies
$(I+\mathcal{C}_{\textup{pr}}^{1/2}UU^{*}\mathcal{C}_{\textup{pr}}^{1/2})w=(1+\gamma)\mathcal{C}_{\textup{pr}}^{1/2}\mathcal{C}_{\textup{pos}}^{-1/2}v=(1+\gamma)(\mathcal{C}_{\textup{pr}}^{1/2}\mathcal{C}_{\textup{pos}}^{-1}\mathcal{C}_{\textup{pr}}^{1/2})w.$
Defining $Z:=\mathcal{C}_{\textup{pr}}^{1/2}U$ and using (3b) and (2), the previous equation yields
$(I+ZZ^{*})w=(1+\gamma)(\mathcal{C}_{\textup{pr}}^{1/2}H\mathcal{C}_{\textup{pr}}^{1/2}+I)w.$
In this way, we recover the formulation that [30] takes as a starting point to solve the optimization problem. See the proof of [30, Theorem 2.3].
As a consequence of Lemma 4.7, we can write $J_{f}$ as
$J_{f}(U)=F_{f}(\Lambda(R(\mathcal{C}_{\textup{pos}}\|(\mathcal{C}_{\textup{pr}}^{-1}+UU^{*})^{-1})))=F_{f}\circ\Lambda\circ g\ (U).$ (19)
which yields the desired reformulation of the loss as a composition of functions. We search for a solution of 4.4 in the set of stationary points of $J_{f}$. For this, we need to show that $J_{f}$ is Gateaux differentiable. To do so, we use the following result, which states that $g$ and $F_{f}$ Fréchet differentiable and gives an explicit form of the derivatives.
Lemma 4.9.
The functions $g$ and $F_{f}$ defined in (17) and (16) respectively are Fréchet differentiable, with derivatives
$g^{\prime}(U)(V)$ $=\mathcal{C}_{\textup{pos}}^{1/2}(UV^{*}+VU^{*})\mathcal{C}_{\textup{pos}}^{1/2},$ $U,V\in\mathcal{B}(\mathbb{R}^{r},\mathcal{H}),$
$F_{f}^{\prime}(x)(y)$ $=\sum_{i}f^{\prime}(x_{i})y_{i},$ $x\in\ell_{2}((-1,\infty)),\ y\in\ell^{2}(\mathbb{R}).$
Remark 4.10.
For a finite set of indices $i\in\{1,\ldots,l\}$, $l\in\mathbb{N}$, the convergence $(f(x_{i}+y_{i})-f(x_{i})-f^{\prime}(x_{i})y_{i})/y_{i}\rightarrow 0$ is uniform $i$. This implies that $\frac{1}{\lVert y\rVert}(\sum_{i}^{l}f(x_{i}+y_{i})-f(x_{i})-f^{\prime}(x_{i})y_{i})\rightarrow 0$, which implies differentiability of $F_{f}$ for finite-dimensional $\mathcal{H}$. In infinite dimensions, the convergence of each term is no longer uniform in $i$, as now $i\in\mathbb{N}$, and the previous sum need not converge to 0. Compared to a finite-dimensional setting, in the infinite-dimensional setting we therefore need to more assumptions of $f$ to assert its convergence. Hence we must restrict the class of spectral functions $\mathscr{F}$. In particular, we require additionally that $f$ has minimum 0 and a derivative which is Lipschitz at 0.
Let $U,V\in\mathcal{B}(\mathbb{R}^{r},\mathcal{H})$. If $\mathcal{W}\subset L_{2}(\mathcal{H})_{\mathbb{R}}$ is a subspace of finite dimension that contains $g(U+tV)$ for all $t\in\mathbb{R}$, then the restriction ${\left.\kern-1.2pt(F_{f}\circ\Lambda)\vphantom{\big{|}}\right|_{\mathcal{W}}}:\mathcal{W}\rightarrow\mathbb{R}$ of $F_{f}\circ\Lambda$ to $\mathcal{W}$ satisfies ${\left.\kern-1.2pt(F_{f}\circ\Lambda)\vphantom{\big{|}}\right|_{\mathcal{W}}}\circ g(U+tV)=F_{f}\circ\Lambda\circ g(U+tV)$ for all $t\in\mathbb{R}$. Thus, $F_{f}\circ\Lambda\circ g$ is Gateaux differentiable in $U$ in direction $V$ if and only if ${\left.\kern-1.2pt(F_{f}\circ\Lambda)\vphantom{\big{|}}\right|_{\mathcal{W}}}\circ g$ is. Hence, by the chain rule, e.g. [17, Section 1.4.1] or [5, Theorem 12.2.2], it suffices to show that ${\left.\kern-1.2pt(F_{f}\circ\Lambda)\vphantom{\big{|}}\right|_{\mathcal{W}}}$ is Fréchet differentiable on all of $(\mathcal{W},\lVert\cdot\rVert_{L_{2}(\mathcal{H})})$ and that $g$ is Gateaux differentiable in $U$ in direction $V$ to be able to conclude the Gateaux differentiability of $J_{f}=F_{f}\circ\Lambda\circ g$ in $U$ in direction $V$.
We now use the finite-dimensional result of [19, Theorem 1.1] on differentiability of spectral functions to deduce Fréchet differentiability of ${\left.\kern-1.2pt(G\circ\Lambda)\vphantom{\big{|}}\right|_{\mathcal{W}}}$ for certain $\mathcal{W}\subset L_{2}(\mathcal{H})_{\mathbb{R}}$ and $G:\mathbb{R}^{\dim{\mathcal{H}}}\rightarrow\mathbb{R}$. This is done in Proposition 4.12 and Proposition 4.14. For this purpose, we introduce the following definition.
Definition 4.11.
Let $m\in\mathbb{N}\cup\{\infty\}$. A set $\Omega\subset\mathbb{R}^{m}$ is symmetric if $Px\in\Omega$ for every $x\in\Omega$ and every permutation $P:\mathbb{R}^{m}\rightarrow\mathbb{R}^{m}$. If $\Omega\subset\mathbb{R}^{\mathbb{N}}$ is symmetric, then a function $G:\Omega\rightarrow\mathbb{R}$ is symmetric if $G(Px)=G(x)$ for every $x\in\Omega$ and every permutation $P:\mathbb{R}^{m}\rightarrow\mathbb{R}^{m}$.
As an example of a symmetric set and symmetric function, consider respectively $\ell^{2}((-1,\infty])$ and $F_{f}$ from (16) for any $f\in\mathscr{F}$.
In analogy to the eigenvalue map $\Lambda:L_{2}(\mathcal{H})_{\mathbb{R}}\rightarrow\ell^{2}(\mathbb{R})$, we define $\Lambda^{m}:L_{2}(\mathcal{Z})_{\mathbb{R}}\rightarrow\mathbb{R}^{m}$ for any $m$-dimensional subspace $\mathcal{Z}\subset\mathcal{H}$ to be any map that sends $X\in L_{2}(\mathcal{Z})_{\mathbb{R}}$ to its eigenvalue sequence, which we order in a nonincreasing way.
Proposition 4.12.
Let $m\in\mathbb{N}$ and let the set $\Omega\subset\mathbb{R}^{m}$ be open and symmetric, and suppose that $G:\Omega\rightarrow\mathbb{R}$ is symmetric. Let $\mathcal{Z}\subset\mathcal{H}$ be finite-dimensional and let $X\in L_{2}(\mathcal{Z})_{\mathbb{R}}$ be such that $\Lambda^{m}(X)\in\Omega$. Then the function $G\circ\Lambda^{m}:L_{2}(\mathcal{Z})_{\mathbb{R}}\rightarrow\mathbb{R}$ is Fréchet differentiable at $X$ if and only if $G$ is Fréchet differentiable at $\Lambda^{m}(X)\in\mathbb{R}^{m}$. In this case the Fréchet derivative of $G\circ\Lambda^{m}$ at $X$ is
$(G\circ\Lambda^{m})^{\prime}(X)=\sum_{i}G^{\prime}(\Lambda^{m}(X))_{i}e_{i}\otimes e_{i}\in L_{2}(\mathcal{Z}),$
where $(e_{i})_{i}$ is an orthonormal sequence in $\mathcal{Z}$ satisfying $X=\sum_{i}(\Lambda^{m}(X))_{i}e_{i}\otimes e_{i}$.
Remark 4.13.
By definition of the Fréchet derivative, $(G\circ\Lambda^{m})^{\prime}(X)\in L_{2}(\mathcal{Z})_{\mathbb{R}}^{*}$. By the Riesz representation theorem, $L_{2}(\mathcal{Z})_{\mathbb{R}}^{*}\simeq L_{2}(\mathcal{Z})_{\mathbb{R}}$, and we consider $(G\circ\Lambda^{m})^{\prime}(X)$ as an element of $L_{2}(\mathcal{Z})_{\mathbb{R}}$.
As a consequence of Proposition 4.12, we can now deduce the Fréchet differentiability of ${\left.\kern-1.2pt(F\circ\Lambda)\vphantom{\big{|}}\right|_{\mathcal{W}}}$ for symmetric functions $F$ and for a suitable subspace $\mathcal{W}$ of $L_{2}(\mathcal{H})_{\mathbb{R}}$.
Proposition 4.14.
Let $\mathcal{Z}\subset\mathcal{H}$ be a finite-dimensional subspace. Let $\mathcal{W}\coloneqq\{X\in L_{2}(\mathcal{H})_{\mathbb{R}}:\ \operatorname{ran}{X}\subset\mathcal{Z}\}\subset L_{2}(\mathcal{H})_{\mathbb{R}}$. Let $F:\ell^{2}(\mathbb{R})\rightarrow\mathbb{R}$ be a symmetric function and let $X\in\mathcal{W}$. Then $\ker{X}^{\perp}=\operatorname{ran}{X}$, and if $F$ is Fréchet differentiable in $\Lambda(X)$, then ${\left.\kern-1.2pt(F\circ\Lambda)\vphantom{\big{|}}\right|_{\mathcal{W}}}$ is Fréchet differentiable in $X\in\mathcal{W}$. In this case, the Fréchet derivative is given by
${\left.\kern-1.2pt(F\circ\Lambda)\vphantom{\big{|}}\right|_{\mathcal{W}}}^{\prime}(X)=\sum_{i}F^{\prime}(\Lambda(X))_{i}e_{i}\otimes e_{i}\in L_{2}(\mathcal{H})_{\mathbb{R}},$
where $(e_{i})_{i}$ is an orthonormal sequence in $\mathcal{Z}$ satisfying $X=\sum_{i}\Lambda(X)_{i}e_{i}\otimes e_{i}.$
Recall that $J_{f}=F_{f}\circ\Lambda\circ g$ by (19). We can now prove Gateaux differentiability of $J_{f}$.
Proposition 4.15.
Let $F_{f}$ and $g$ be defined in (16) and (17). Then $J_{f}$ is Gateaux differentiable on $\mathcal{B}(\mathbb{R}^{r},\mathcal{H})$, and for any $U,V\in\mathcal{B}(\mathbb{R}^{r},\mathcal{H})$, the Gateaux derivative at $U$ in direction $V$ is given by
$J_{f}^{\prime}(U)(V)=2\sum_{i}f^{\prime}(\Lambda_{i}(g(U)))\langle\mathcal{C}_{\textup{pos}}^{1/2}e_{i},VU^{*}\mathcal{C}_{\textup{pos}}^{1/2}e_{i}\rangle,$
where $(e_{i})_{i}$ is an ONB of $\mathcal{H}$ satisfying $g(U)=\sum_{i}\Lambda_{i}(g(U))e_{i}\otimes e_{i}$.
A coercive and Gateaux differentiable function has a global minimum, that can be found among its stationary points. The following lemma establishes the coercivity of $J_{f}$ over every finite-dimensional subspace of $\mathcal{H}$.
Lemma 4.16.
Let $f\in\mathscr{F}$ and $\mathcal{V}\subset\mathcal{H}$ be finite-dimensional. Then $J_{f}$ is coercive over $\mathcal{B}(\mathbb{R}^{r},\mathcal{V})$, i.e. $J_{f}(U_{n})\rightarrow\infty$ whenever $U_{n}\rightarrow\infty$ in $\mathcal{B}(\mathbb{R}^{r},\mathcal{V})$. In particular, $J_{f}$ has a global minimum on $\mathcal{B}(\mathbb{R}^{r},\mathcal{V})$, which can be found among the stationary points of $J_{f}$.
Unless $\mathcal{H}$ is finite-dimensional, the function $J_{f}$ is no longer coercive on all of $\mathcal{B}(\mathbb{R}^{r},\mathcal{H})$. To show this, we exploit the property that the finite-dimensional ranges of a sequence $U_{n}\in\mathcal{B}(\mathbb{R}^{r},\mathcal{H})$ do not need to lie in the same finite-dimensional subspace of $\mathcal{H}$.
Example 4.17.
Let $\mathcal{H}=\ell^{2}(\mathbb{R})$ and $r=1$. Furthermore, let $(e_{k})_{k}$ be the standard basis of $\mathcal{H}$ and put $\mathcal{C}_{\textup{pos}}e_{k}=k^{-\alpha}e_{k}$ for $\alpha>1$. Let $U_{n}t=n^{\beta}te_{n}$ for $t\in\mathbb{R}$, with $\beta>0$. Then $U_{n}U_{n}^{*}e_{k}=\delta_{n,k}n^{2\beta}e_{n}$. It follows that $\mathcal{C}_{\textup{pos}}^{1/2}U_{n}U_{n}^{*}\mathcal{C}_{\textup{pos}}^{1/2}e_{k}=\delta_{n,k}n^{-\alpha+2\beta}e_{n}$. Hence $\lVert\mathcal{C}_{\textup{pos}}^{1/2}U_{n}U_{n}^{*}\mathcal{C}_{\textup{pos}}^{1/2}\rVert_{2}^{2}=\langle\mathcal{C}_{\textup{pos}}^{1/2}U_{n}U_{n}^{*}\mathcal{C}_{\textup{pos}}^{1/2}e_{n},e_{n}\rangle=n^{-\alpha+2\beta}$ which is bounded from above for $\alpha>2\beta$. Therefore $J_{f}(U_{n})$ is bounded in $n$, while $\lVert U_{n}\rVert=n^{\beta}\rightarrow\infty$.
In the proof of Theorem 4.18, coercivity on finite-dimensional subspaces of $\mathcal{B}(\mathbb{R}^{r},\mathcal{H})$ of the form $\mathcal{B}(\mathbb{R}^{r},\mathcal{V})$, $\mathcal{V}\subset\mathcal{H}$ finite-dimensional, is enough to show the existence of a global minimizer of $J_{f}$, because all the stationary points lie in one such finite-dimensional subspace.
We can now state the main result of this section.
Theorem 4.18.
Let $r\leq n$ and let $(\lambda_{i})_{i}\in\ell^{2}((-1,0])$ and $(w_{i})_{i}\subset\operatorname{ran}{\mathcal{C}_{\textup{pr}}^{1/2}}$ be as given in Lemma 3.6. Define
$\mathcal{P}^{\textup{opt}}_{r}$ $\coloneqq\mathcal{C}_{\textup{pr}}^{-1}+\sum_{i=1}^{r}\frac{-\lambda_{i}}{1+\lambda_{i}}\mathcal{C}_{\textup{pr}}^{-1/2}w_{i}\otimes\mathcal{C}_{\textup{pr}}^{-1/2}w_{i},$ (20)
$\mathcal{C}^{\textup{opt}}_{r}$ $\coloneqq\mathcal{C}_{\textup{pr}}-\sum_{i=1}^{r}-\lambda_{i}\mathcal{C}_{\textup{pr}}^{1/2}w_{i}\otimes\mathcal{C}_{\textup{pr}}^{1/2}w_{i}.$ (21)
Then $\mathcal{P}^{\textup{opt}}_{r}$ and $\mathcal{C}^{\textup{opt}}_{r}$ are solutions to 4.4 and 4.3 respectively and $\mathcal{P}^{\textup{opt}}_{r}$ and $\mathcal{C}^{\textup{opt}}_{r}$ are inverses of each other. For every $f\in\mathscr{F}$, the associated minimal loss is $\mathcal{L}_{f}(\mathcal{C}_{\textup{pos}}\|\mathcal{C}^{\textup{opt}}_{r})=\sum_{i>r}f(\lambda_{i})$. The solutions $\mathcal{P}^{\textup{opt}}_{r}$ and $\mathcal{C}^{\textup{opt}}_{r}$ are unique if and only if either $\lambda_{r}=0$ or $\lambda_{r}<\lambda_{r+1}$.
Theorem 4.18 shows that $\mathcal{C}^{\textup{opt}}_{r}$ and $\mathcal{P}^{\textup{opt}}_{r}$ are the optimal low-rank updates of the $\mathcal{C}_{\textup{pr}}$ and $\mathcal{C}_{\textup{pr}}^{-1}$ respectively, for all $\mathcal{L}\in\mathscr{L}$ simultaneously. By Lemma 4.2, this optimality includes optimality with respect to the forward and reverse KL-divergence and the Rényi divergences when keeping the posterior mean fixed. This also holds for the Hellinger distance. Indeed, if we denothe Hellinger distance between measures $\nu_{1}$ and $\nu_{2}$ by $d_{\textup{H}}(\nu_{1},\nu_{2})$, it holds that $d_{\textup{H}}(\nu_{1},\nu_{2})^{2}=2(1-\exp(-D_{\textup{Ren},1/2}(\nu_{1}\|\nu_{2})))$ by e.g. [22, Equations (134)–(135)], it follows that minimising the Hellinger distance $d_{\textup{H}}(\nu_{1},\nu_{2})$ is equivalent to minimising the Bhattacharyya distance $D_{\textup{Ren},1/2}(\nu_{1}\|\nu_{2})$.
5 Optimal approximation of the mean
In this section, we discuss an optimal low-rank approximation procedure for the posterior mean $m_{\textup{pos}}(y)=\mathcal{C}_{\textup{pos}}G^{*}\mathcal{C}_{\textup{obs}}^{-1}y$, see (3a). Given the data $y$, the approximations considered are of the form $Ay$, with $A\in\mathscr{M}^{(i)}$ for $i=1,2$. Because $\operatorname{ran}{G^{*}}=\ker{G}^{\perp}$, the classes in (6) with $r\leq n$ can be described by Theorem 3.1 as
$\mathscr{M}^{(1)}_{r}=$$\{(\mathcal{C}_{\textup{pr}}-B)G^{*}\mathcal{C}_{\textup{obs}}^{-1}:\ B\in\mathcal{B}_{00,r}(\mathcal{H})_{\mathbb{R}},\ B(\ker{G}^{\perp})\subset\operatorname{ran}{\mathcal{C}_{\textup{pr}}^{1/2}}\},$ (22a)
$\mathscr{M}^{(2)}_{r}=$$\{A\in\mathcal{B}_{00,r}(\mathbb{R}^{n},\mathcal{H}):\ \operatorname{ran}{A}\subset\mathcal{C}_{\textup{pr}}^{1/2}\}.$ (22b)
We shall assess the approximation quality of an approximate posterior mean by averaging the mean dependent term for the Rényi divergence and the reverse KL-divergence (14a) over all possible realisation $y$ of $Y$. By averaging over $Y$, the optimal operator $A$ will be data-independent, i.e. will not depend on a specific realisation $y$. While averaging over $Y$ implies that the resulting posterior mean approximations are not optimal in general for a specific realisation $y$ of $Y$, this approach has the benefit that $A$ can be constructed prior to observing the data. In this way, it facilitates an offline-online approach to posterior mean approximation: the preliminary ‘offline’ stage computes one operator, which can then can be applied in the subsequent ‘online’ stage to any realisations of the data. This is in analogy to the finite-dimensional case studied in [30, Section 4.1]. Furthermore, averaging over $Y$ enables us to exploit recent work on reduced rank operator approximation [6].
Recall that we use the observation model $Y=Gx^{\dagger}+\zeta$ for $\zeta\sim\mathcal{N}(0,\mathcal{C}_{\textup{obs}})$ for $G\in\mathcal{B}(\mathcal{H},\mathbb{R}^{n})$ and positive $\mathcal{C}_{\textup{obs}}\in\mathcal{B}(\mathbb{R}^{n})_{\mathbb{R}}$, and that our prior model for $x^{\dagger}$ is $X\sim\mathcal{N}(0,\mathcal{C}_{\textup{pr}})$, with $X$ and $\zeta$ independent. These assumptions imply that $Y\sim\mathcal{N}(0,\mathcal{C}_{\textup{y}})$, where $\mathcal{C}_{\textup{y}}\coloneqq G\mathcal{C}_{\textup{pr}}G^{*}+\mathcal{C}_{\textup{obs}}\in\mathcal{B}(\mathbb{R}^{n})$.
Since $R(\mathcal{C}\|\mathcal{C})=0$ for any positive $\mathcal{C}\in L_{1}(\mathcal{H})_{\mathbb{R}}$, by Theorem 3.8, the Rényi divergence and reverse KL divergence of approximating $\mathcal{N}(m_{\textup{pos}},\mathcal{C})$ by $\mathcal{N}(m,\mathcal{C})$ for any $m\in\mathcal{H}$ satisfying $m-m_{\textup{pos}}\in\operatorname{ran}{\mathcal{C}^{1/2}}$ is given by, for any $\rho\in(0,1)$,
$D_{\textup{KL}}(\mathcal{N}(m,\mathcal{C})\|\mathcal{N}(m_{\textup{pos}},\mathcal{C}))=\lVert m-m_{\textup{pos}}\rVert^{2}_{\mathcal{C}^{-1}}=D_{\textup{Ren},\rho}(\mathcal{N}(m,\mathcal{C})\|\mathcal{N}(m_{\textup{pos}},\mathcal{C})).$
The optimal low-rank posterior mean is therefore the solution to the following problem.
Problem 5.1.
Let $r\leq n$ and $i\in\{1,2\}$. Find $A^{\textup{opt},(i)}_{r}\in\mathscr{M}_{r}^{(i)}$ such that
$\mathbb{E}\left[\lVert A^{\textup{opt},(i)}_{r}Y-m_{\textup{pos}}(Y)\rVert_{\mathcal{C}_{\textup{pos}}^{-1}}^{2}\right]=\min\left\{\mathbb{E}\left[\lVert\mathcal{C}_{\textup{pos}}^{-1/2}(AY-m_{\textup{pos}}(Y))\rVert^{2}\right]:\ A\in\mathscr{M}_{r}^{(i)}\right\}.$
The term on the inside of the expectation on the left hand side above corresponds to the mean dependent term in (14a), and has the interpretation that it penalises errors in approximation of the posterior mean more in those directions in which the posterior covariance is small.
In our derivation of the optimal $A^{\textup{opt},(i)}_{r}$, we shall make use of the specific non self-adjoint square roots $S_{\textup{pos}}\in L_{2}(\mathcal{H})$ and $S_{\textup{y}}\in\mathcal{B}(\mathbb{R}^{n})$ of the covariances $\mathcal{C}_{\textup{pos}}$ and $\mathcal{C}_{y}$ respectively. Since $n<\infty$, $\mathcal{C}_{\textup{obs}}^{-1}$ is bounded and self-adjoint and we can decompose $\mathcal{C}_{\textup{obs}}^{-1}=\mathcal{C}_{\textup{obs}}^{-1/2}(\mathcal{C}_{\textup{obs}}^{-1/2})^{*}$ by Lemma A.13. Therefore, by (10a) in Lemma 3.6.
$(\mathcal{C}_{\textup{pr}}^{1/2}G^{*}\mathcal{C}_{\textup{obs}}^{-1/2})(\mathcal{C}_{\textup{pr}}^{1/2}G^{*}\mathcal{C}_{\textup{obs}}^{-1/2})^{*}=\mathcal{C}_{\textup{pr}}^{1/2}H\mathcal{C}_{\textup{pr}}^{1/2}=\sum_{i=1}^{n}\frac{-\lambda_{i}}{1+\lambda_{i}}w_{i}\otimes w_{i},$ (23)
with $(w_{i})_{i}$ and $(\lambda_{i})_{i}$ as in Lemma 3.6. Thus, there exists an orthonormal sequence $(\varphi_{i})_{i}$ in $\mathbb{R}^{n}$ such that
$\mathcal{C}_{\textup{pr}}^{1/2}G^{*}\mathcal{C}_{\textup{obs}}^{-1/2}=\sum_{i=1}^{n}\sqrt{\frac{-\lambda_{i}}{1+\lambda_{i}}}w_{i}\otimes\varphi_{i}.$ (24)
We now define,
$\begin{split}S_{\textup{pos}}&=\mathcal{C}_{\textup{pr}}^{1/2}\left(I+\sum_{i\in\mathbb{N}}\frac{-\lambda_{i}}{1+\lambda_{i}}w_{i}\otimes w_{i}\right)^{-1/2},\\ S_{\textup{y}}&=\mathcal{C}_{\textup{obs}}^{1/2}\left(I+\sum_{i=1}^{n}\frac{-\lambda_{i}}{1+\lambda_{i}}\varphi_{i}\otimes\varphi_{i}\right)^{1/2}.\end{split}$ (25)
Note that $\sum_{i=1}^{m}(1+\frac{-\lambda_{i}}{1+\lambda_{i}})w_{i}\otimes w_{i}$ does not converge in $\mathcal{B}(\mathcal{H})$ as $m\rightarrow\infty$, when $\mathcal{H}$ is infinite-dimensional. Indeed, if $\sum_{i=1}^{m}(1+\frac{-\lambda_{i}}{1+\lambda_{i}})w_{i}\otimes w_{i}$ converges, then $\sum_{i=1}^{m}(1+\frac{-\lambda_{i}}{1+\lambda_{i}})w_{i}\otimes w_{i}-\sum_{i=1}^{n}\frac{-\lambda_{i}}{1+\lambda_{i}}w_{i}\otimes w_{i}$ is a sequence of finite rank operators converging to the identity. Since the identity in $\mathcal{B}(\mathcal{H})$ is not compact when $\mathcal{H}$ is infinite-dimensional, the series $\sum_{i=1}^{m}(1+\frac{-\lambda_{i}}{1+\lambda_{i}})w_{i}\otimes w_{i}$ does not converge as $m\rightarrow\infty$. However, there is in fact pointwise convergence: for $h\in\mathcal{H}$, we may compute,
$\left(I+\sum_{i}\frac{-\lambda_{i}}{1+\lambda_{i}}w_{i}\otimes w_{i}\right)h=\sum_{i}\left(1+\frac{-\lambda_{i}}{1+\lambda_{i}}\right)\langle h,w_{i}\rangle w_{i}=\sum_{i}\frac{1}{1+\lambda_{i}}\langle h,w_{i}\rangle w_{i}.$
Similarly, a direct computation shows that for $h\in\mathcal{H}$ and $x\in\mathbb{R}^{n}$,
$\left(I+\sum_{i}\frac{-\lambda_{i}}{1+\lambda_{i}}w_{i}\otimes w_{i}\right)^{-1/2}h$ $=\sum_{i}(1+\lambda_{i})^{1/2}\langle h,w_{i}\rangle w_{i},$ (26a)
$\left(I+\sum_{i}\frac{-\lambda_{i}}{1+\lambda_{i}}\varphi_{i}\otimes\varphi_{i}\right)^{1/2}x$ $=\sum_{i}(1+\lambda_{i})^{-1/2}\langle x,\varphi_{i}\rangle\varphi_{i},$ (26b)
We first note that $S_{\textup{pos}},S_{\textup{y}}$ are indeed square roots, and that they have well-defined inverses.
Lemma 5.2.
Let $S_{\textup{pos}},S_{\textup{y}}$ be as in (25). It holds that $\mathcal{C}_{\textup{pos}}=S_{\textup{pos}}S_{\textup{pos}}^{*}$ and $\mathcal{C}_{\textup{y}}=S_{\textup{y}}S_{\textup{y}}^{*}$ and $S_{\textup{pos}}^{-1}:\operatorname{ran}{C}_{\textup{pr}}^{1/2}\rightarrow\mathcal{H}$ and $S_{\textup{y}}^{-1}\in\mathcal{B}(\mathbb{R}^{n})$ exist.
Let $S_{\textup{pos}}^{-*}:=(S_{\textup{pos}}^{-1})^{*}=(S_{\textup{pos}}^{*})^{-1}$. Given Lemma 5.2, we can write
$\lVert h\rVert_{\mathcal{C}_{\textup{pos}}^{-1}}=\lVert\mathcal{C}_{\textup{pos}}^{-1/2}h\rVert=\langle\mathcal{C}_{\textup{pos}}^{-1}h,h\rangle=\langle S_{\textup{pos}}^{-*}\mathcal{S}_{\textup{pos}}^{-1}h,h\rangle=\lVert S_{\textup{pos}}^{-1}h\rVert^{2}.$ (27)
The equation (27) can be used to evaluate the norms in 5.1 by replacing $\mathcal{C}_{\textup{pos}}^{-1/2}$ by $S_{\textup{pos}}^{-1}$. Furthermore, for each $A\in\mathcal{B}(\mathbb{R}^{n},\mathcal{H})$ with $\operatorname{ran}{A}\subset\operatorname{ran}{\mathcal{C}_{\textup{pr}}^{1/2}}$, there exists a unique $\widetilde{A}\in\mathcal{B}(\mathbb{R}^{n},\mathcal{H})$ such that $A=S_{\textup{pos}}\widetilde{A}$, i.e. $\widetilde{A}\coloneqq S_{\textup{pos}}^{-1}A$. Let us now define,
~$\widetilde{\mathscr{M}}^{(1)}_{r}$ $:=S_{\textup{pos}}^{-1}\mathscr{M}_{r}^{(2)}=\{(S_{\textup{pos}}^{-1}\mathcal{C}_{\textup{pr}}-\widetilde{B})G^{*}\mathcal{C}_{\textup{obs}}^{-1}:\ \widetilde{B}\in\mathcal{B}_{00,r}(\mathcal{H})\},$
~$\widetilde{\mathscr{M}}^{(2)}_{r}$ $:=S_{\textup{pos}}^{-1}\mathscr{M}_{r}^{(1)}=\mathcal{B}_{00,r}(\mathbb{R}^{n},\mathcal{H}).$
Then we can reformulate 5.1 as follows.
Problem 5.3.
Let $r\leq n$ and $i\in\{1,2\}$. Find $\widetilde{A}^{\textup{opt},(i)}_{r}\in\widetilde{\mathscr{M}}_{r}^{(i)}$ such that
$\mathbb{E}\left[\left\lVert\widetilde{A}^{\textup{opt}}Y-S_{\textup{pos}}^{-1}m_{\textup{pos}}(Y)\right\rVert^{2}\right]=\min\left\{\mathbb{E}\left[\left\lVert\widetilde{A}Y-S_{\textup{pos}}^{-1}m_{\textup{pos}}(Y)\right\rVert^{2}\right],\ \widetilde{A}\in\widetilde{\mathscr{M}}^{(i)}\right\}.$
The solutions to 5.1 and 5.3 are related via $S_{\textup{pos}}\widetilde{A}^{\textup{opt},(i)}_{r}=A^{\textup{opt},(i)}_{r}$.
Lemma 5.4.
It holds that
$\mathbb{E}\left\lVert\widetilde{A}Y-S_{\textup{pos}}^{-1}m_{\textup{pos}}(Y)\right\rVert^{2}$ $=\left\lVert\widetilde{A}S_{\textup{y}}-\mathcal{C}_{\textup{pr}}^{1/2}G^{*}\mathcal{C}_{\textup{obs}}^{-1/2}\right\rVert_{L_{2}(\mathcal{H})}^{2},\quad\widetilde{A}\in\mathcal{B}(\mathcal{H}).$ (28)
In order to solve 5.3, we use a result on reduced rank operator approximation in $L_{2}(\mathcal{H})$ norm, proven in [6]. It is a generalised version of the Eckhart–Young theorem. For $C\in\mathcal{B}(\mathcal{H})$ we denote by $C^{\dagger}$ the Moore–Penrose, or pseudo-inverse, of $C$. If $C$ has closed range, then $C^{\dagger}$ is bounded, c.f. [13, Proposition 2.4]. The following is an application of [6, Theorem 3.2, Remark 3.5] to the case where the operators $B$ and $C$ occurring in the theorem have closed range.
Theorem 5.5.
Let $\mathcal{H}_{1},\mathcal{H}_{2},\mathcal{H}_{3},\mathcal{H}_{4}$ be Hilbert spaces and let $T\in\mathcal{B}(\mathcal{H}_{3},\mathcal{H}_{4})$, $S\in\mathcal{B}(\mathcal{H}_{1},\mathcal{H}_{2})$ both have closed range and let $M\in L_{2}(\mathcal{H}_{1},\mathcal{H}_{4})$. Then, for each rank-$r$ truncated SVD $(P_{\operatorname{ran}{T}}MP_{\ker{S}^{\perp}})_{r}$ of $P_{\operatorname{ran}{T}}MP_{\ker{S}^{\perp}}$,
$\hat{N}\coloneqq T^{\dagger}(P_{\operatorname{ran}{T}}MP_{\ker{S}^{\perp}})_{r}S^{\dagger},$ (29)
is a solution to the problem,
$\min\{\lVert M-TNS\rVert_{L_{2}(\mathcal{H}_{1},\mathcal{H}_{4})},\ N\in\mathcal{B}_{00,r}(\mathcal{H}_{2},\mathcal{H}_{3})\},$ (30)
such that
$N=P_{\ker{T}^{\perp}}NP_{\operatorname{ran}{S}}.$ (31)
Furthermore, (29) is the only solution of (30) satisfying (31) if and only if one of the following conditions holds:
  1. (i)
    $r\geq\operatorname{rank}\left({P_{\operatorname{ran}{T}}MP_{\ker{S}^{\perp}}}\right),$
  2. (ii)
    $\sigma_{r}(P_{\operatorname{ran}{T}}MP_{\ker{S}^{\perp}})>\sigma_{r+1}(P_{\operatorname{ran}{T}}MP_{\ker{S}^{\perp}}).$
With Theorem 5.5, we can now identify solutions of 5.1, by solving 5.3 for $\widetilde{A}^{\textup{opt},(i)}\in\widetilde{\mathscr{M}}^{\textup{opt},(i)}$ and letting ${A}^{\textup{opt},(i)}=S_{\textup{pos}}\widetilde{A}^{\textup{opt},(i)}$. We first consider the low-rank posterior mean approximation problem for the approximation class ${\mathscr{M}}_{r}^{(2)}$ given in (22b), compute the corresponding minimal loss, and show that the solution $A^{\textup{opt},(2)}$ not only satisfies $\operatorname{ran}{A^{\textup{opt},(2)}}\subset\operatorname{ran}{\mathcal{C}}_{\textup{pr}}^{1/2}$, but also $\operatorname{ran}{A^{\textup{opt},(2)}}\subset\operatorname{ran}{\mathcal{C}}_{\textup{pr}}$. The latter condition is also satisfied by the exact posterior mean: $\mathcal{C}_{\textup{pos}}G^{*}\mathcal{C}_{\textup{obs}}^{-1}\subset\operatorname{ran}{\mathcal{C}_{\textup{pr}}}$.
Theorem 5.6.
Fix $r\leq n$. Let $(\lambda_{i},w_{i})_{i}$ be as in Lemma 3.6 and $(\varphi_{i})_{i=1}^{n}$ be as in (24). Then a solution of 5.1 for $i=2$ is given by $A_{r}^{\textup{opt},(2)}=\mathcal{C}_{\textup{pr}}^{1/2}\sum_{i=1}^{r}\sqrt{-\lambda_{i}(1+\lambda_{i})}w_{i}\otimes\varphi_{i}\mathcal{C}_{\textup{obs}}^{-1/2}\in\mathscr{M}_{r}^{(2)}.$ Furthermore, $\operatorname{ran}{A_{r}^{\textup{opt},(2)}}\subset\operatorname{ran}{\mathcal{C}_{\textup{pos}}}$. The corresponding loss is $\sum_{i>r}\frac{-\lambda_{i}}{1+\lambda_{i}}.$
Next, we solve 5.1 for the approximation class $\mathcal{\mathscr{M}}_{r}^{(1)}$, and show that the solutions in fact satisfy $\operatorname{ran}{A^{\textup{opt},(1)}}\subset\operatorname{ran}{\mathcal{C}}_{\textup{pr}}$.
Theorem 5.7.
Fix $r\leq n$. Let $(\lambda_{i})_{i}$ be as in Lemma 3.6 and $\mathcal{C}^{\textup{opt}}_{r}$ be the optimal rank-$r$ approximation of $\mathcal{C}_{\textup{pos}}$ from (21) in Theorem 4.18. Then a solution of 5.1 for $i=1$ is given by $A_{r}^{\textup{opt},(1)}=\mathcal{C}^{\textup{opt}}_{r}G^{*}\mathcal{C}_{\textup{obs}}^{-1}\in\mathscr{M}_{r}^{(1)}$. Furthermore, $\operatorname{ran}{A_{r}^{\textup{opt},(1)}}\subset\operatorname{ran}{\mathcal{C}_{\textup{pos}}}$. The corresponding loss is $\sum_{i>r}\left(\frac{-\lambda_{i}}{1+\lambda_{i}}\right)^{3}.$
Theorem 5.6 and Theorem 5.7 generalise the results of [30, Theorem 4.1 and Theorem 4.2] to an infinite-dimensional setting. We note that in both considered approximation classes $\mathscr{M}_{r}^{(i)}$, $i\in\{1,2\}$, the optimal operator $A^{\textup{opt},(i)}_{r}$ maps into $\operatorname{ran}{\mathcal{C}_{\textup{pr}}}$, just like the exact operator $\mathcal{C}_{\textup{pos}}G^{*}\mathcal{C}_{\textup{obs}}^{-1}$ does.
Furthermore, as was noted in [30], a comparison between the minimal losses of Theorem 5.6 and Theorem 5.7 gives us insight as to which approximation procedure is preferable in a specific setting. As the theorems show, the decay of the eigenvalues $(\lambda_{i})_{i}$ of $R(\mathcal{C}_{\textup{pos}}\|\mathcal{C}_{\textup{pr}})$ governs this choice. The loss of the optimal approximation in Theorem 5.6 and in Theorem 5.7 is $\sum_{i>r}(\tfrac{-\lambda_{i}}{1+\lambda_{i}})$ and $\sum_{i>r}(\tfrac{-\lambda_{i}}{1+\lambda_{i}})^{3}$ respectively. Recall that the $(\lambda_{i})_{i}$ in Lemma 3.6 are ordered in a nondecreasing way. If $\tfrac{-\lambda_{i}}{1+\lambda_{i}}\leq 1$ or equivalently $-\lambda_{i}\leq\tfrac{1}{2}$ for every $i>r$, then we have $\sum_{i>r}(\tfrac{-\lambda_{i}}{1+\lambda_{i}})\geq\sum_{i>r}(\tfrac{-\lambda_{i}}{1+\lambda_{i}})^{3}$. Since the sequence $(\lambda_{i})_{i}\subset(-1,0]$ increases to zero by Lemma 3.6, and since $(\lambda_{i})_{i}$ have the interpretation of variance reduction by Remark 3.7, it follows that if there exists some $r<n$ such that the relative variance reduction along $\mathcal{C}_{\textup{pr}}^{-1/2}w_{i}$ is smaller than $\tfrac{1}{2}$ for $i>r$, then the loss $\sum_{i>r}(\tfrac{-\lambda_{i}}{1+\lambda_{i}})^{3}$ that arises from exploiting the structure (3a) of the posterior mean is smaller than the loss that ignores this structure. In other words, one can achieve on average a smaller loss in the posterior mean approximation that exploits the structure (3a) of the posterior mean, if the ratio of the posterior variance to the prior variance along $\mathcal{C}_{\textup{pr}}^{-1/2}w_{i}$ decays below the threshold of $\tfrac{1}{2}$ for sufficiently large $i$. If for example $\lambda_{i}>-\tfrac{1}{2}$ for every $i\in\mathbb{N}$, then this decay does not occur, and one can obtain a smaller loss by ignoring the structure.
6 Optimal joint approximation of the mean and covariance
In Section 4, we considered the optimal rank-$r$ approximation of the posterior covariance given the same mean, while in Section 5 we considered the optimal rank-$r$ approximation of the posterior mean given the same posterior covariance. In this section, we consider jointly approximating the posterior mean and covariance in the reverse KL divergence defined in Section 3.2.
Let $y\in\mathbb{R}^{n}$ be an arbitrary data vector and $m_{\textup{pos}}(y)$ be as in (3a). Let $\widetilde{m}_{\textup{pos}}(y)$ be an approximation of $m_{\textup{pos}}(y)$ and $\widetilde{\mathcal{C}}_{\textup{pos}}$ be an approximation of $\mathcal{C}_{\textup{pos}}$ such that $N(\widetilde{m}_{\textup{pos}}(y),\mathcal{C}_{\textup{pos}})\sim\mu_{\textup{pos}}$, and let $m\in\mathcal{H}$ be arbitrary. Then, by (14a) and Lemma 4.2(i),
$D_{\textup{KL}}(\mathcal{N}(\widetilde{m}_{\textup{pos}}(y),\widetilde{\mathcal{C}}_{\textup{pos}})\|\mu_{\textup{pos}})$ $=\frac{1}{2}\left\lVert\mathcal{C}_{\textup{pos}}^{-1/2}(\widetilde{m}_{\textup{pos}}(y)-m_{\textup{pos}}(y))\right\rVert^{2}-\frac{1}{2}\log\det_{2}\left(I+R(\widetilde{\mathcal{C}}_{\textup{pos}}\|\mathcal{C}_{\textup{pos}})\right)$
$=\frac{1}{2}\left\lVert\mathcal{C}_{\textup{pos}}^{-1/2}(\widetilde{m}_{\textup{pos}}(y)-m_{\textup{pos}}(y))\right\rVert^{2}+D_{\textup{KL}}(\mathcal{N}(m,\widetilde{\mathcal{C}}_{\textup{pos}})\|\mathcal{N}(m,\mathcal{C}_{\textup{pos}}))$
$=D_{\textup{KL}}(\mathcal{N}(\widetilde{m}_{\textup{pos}}(y),\mathcal{C}_{\textup{pos}})\|\mathcal{N}(m_{\textup{pos}}(y),\mathcal{C}_{\textup{pos}}))$
$+D_{\textup{KL}}(\mathcal{N}(m,\widetilde{\mathcal{C}}_{\textup{pos}})\|\mathcal{N}(m,\mathcal{C}_{\textup{pos}})),$
which constitutes a Pythagorean identity for the Kullback–Leibler divergence between two Gaussians. The identity above is reasonable, since the Kullback–Leibler divergence is a Bregman divergence, which are known to satisfy generalised Pythagorean theorems. See e.g. [2, Section 1.6] or [24] for the information geometry perspective on Pythagorean identities and [21, Theorem 2.1] for a Pythagorean theorem in the context of dimension reduction for Bayesian inverse problems.
In our context, the Pythagorean identity above implies that, in order to solve the joint approximation problem, it suffices to solve the posterior mean approximation problem and the posterior covariance approximation problems separately. Let $r\in\mathbb{N}$. Suppose we search for $\widetilde{m}_{\textup{pos}}(y)$ of the form $Ay$ for $A$ in one of the approximation classes $\mathscr{M}^{(i)}_{r}$ defined in (6), and that we search for $\widetilde{\mathcal{C}}_{\textup{pos}}$ of the form $(\mathcal{C}_{\textup{pr}}^{-1}+UU^{*})^{-1}$ from $\mathscr{C}_{r}$ defined in (4). Then for $i=1,2$,
$\min\left\{\mathbb{E}\left[D_{\textup{KL}}(\mathcal{N}(AY,(\mathcal{C}_{\textup{pr}}^{-1}+UU^{*})^{-1})\|\mathcal{N}(m_{\textup{pos}}(Y),\mathcal{C}_{\textup{pos}}))\right]:\ A\in\mathscr{M}^{(i)}_{r},(\mathcal{C}_{\textup{pr}}^{-1}+UU^{*})^{-1}\in\mathscr{C}_{r}\right\}$
$=$ $\min\left\{\mathbb{E}\left[D_{\textup{KL}}(\mathcal{N}(AY,\mathcal{C}_{\textup{pos}})\|\mathcal{N}(m_{\textup{pos}}(Y),\mathcal{C}_{\textup{pos}})\right]:\ A\in\mathscr{M}^{(i)}_{r}\right\}$
$+$ $\min\left\{D_{\textup{KL}}(\mathcal{N}(m,(\mathcal{C}_{\textup{pr}}^{-1}+UU^{*})^{-1})\|\mathcal{N}(m,\widetilde{\mathcal{C}}_{\textup{pos}})):\ (\mathcal{C}_{\textup{pr}}^{-1}+UU^{*})^{-1})\in\mathscr{C}_{r}\right\}.$
The two minimisation problems can then be solved using the Theorem 4.18 and either Theorem 5.6 or Theorem 5.7, and the corresponding minimal loss is found by summing the respective losses:
$\min\left\{\mathbb{E}\left[D_{\textup{KL}}(\mathcal{N}(AY,(\mathcal{C}_{\textup{pr}}^{-1}+UU^{*})^{-1})\|\mathcal{N}(m_{\textup{pos}}(Y),\mathcal{C}_{\textup{pos}}))\right]:\ A\in\mathscr{M}^{(i)}_{r},(\mathcal{C}_{\textup{pr}}^{-1}+UU^{*})^{-1}\in\mathscr{C}_{r}\right\}$
$=$ $\sum_{i>r}f_{\textup{KL}}\left(\frac{-\lambda_{i}}{1+\lambda_{i}}\right)+\left(\frac{-\lambda_{i}}{1+\lambda_{i}}\right)^{\alpha(i)},$
where $\alpha(1)=3$ as in Theorem 5.6 and $\alpha(2)=1$ as in Theorem 5.7, and where $f_{\textup{KL}}$ is defined in Lemma 4.2(i).
7 Conclusion
Linear Gaussian inverse problems on possibly infinite-dimensional Hilbert spaces are an important kind of nonparametric inverse problem. For example, they can be used to approximate nonlinear nonparametric problems using the Laplace approximation, and they often serve as the native infinite-dimensional formulation of linear inverse problems before the parameter space $\mathcal{H}$ is discretised and they are in this sense ‘dimension independent’.
Optimal low-rank approximation of the posterior covariance for a class of losses that includes the KL divergence and the Hellinger metric, and of the posterior mean for the Bayes Risk, was studied in [30]. The analysis showed that certain matrix pencils, namely the ones defined by the Hessian and prior covariance and the prior and posterior covariance, form the central objects of study.
So far, these results applied to finite-dimensional parameter spaces only. In this work we have derived optimal low-rank approximations to the posterior covariance, posterior mean, and entire posterior distribution, on possibly infinite-dimensional parameter spaces, thereby generalising the conclusions of [30]. The resulting posterior approximations are equivalent with the exact posterior distribution. Moreover, the optimal low-rank updates for the covariance and mean share the property with the exact updates that their range lies in $\operatorname{ran}{\mathcal{C}_{\textup{pos}}}$. We have shown that the posterior covariance approximations are optimal for a class of losses which includes the KL divergence, the reverse KL divergence, the Hellinger metric and the Rényi divergences. The constructed low-rank approximation of the mean is optimal for the reverse KL divergence and the Rényi divergences, and the low-rank approximation of the mean and covariance jointly is optimal for the reverse KL divergence. Furthermore, we have shown that the operator pencils which proved central in the finite-dimensional analysis, are both equivalent representations of the Hilbert–Schmidt operator appearing in the Feldman–Hajek theorem which quantifies similarity of Gaussian measures. It is therefore this operator that is central to the reduction of the posterior distribution for linear Gaussian inverse problems.
Finally, the low-rank approximations constructed in this work, which have been proven to be optimal in various measures of divergence, may provide a basis for the development of computationally efficient approximation of certain linear Gaussian problems, such as Gaussian Process Regression, and for optimal approximation of nonlinear nonparametric inverse problems.
8 Acknowledgements
The research of the authors has been partially funded by the Deutsche Forschungsgemeinschaft (DFG) Project-ID 318763901 – SFB1294. The authors thank Ricardo Baptista and Youssef Marzouk for helpful remarks concerning the joint approximation problem, and Bernhard Stankewitz for helpful discussions.
Appendix A Auxiliary results
Lemma A.1.
Let $X$ be a Banach space, $S\subset X$ a subspace, and $T\in\mathcal{B}(X)$ such that $T(S)$ is closed. Then $T(\overline{S})=T(S)$.
Proof.
By continuity of $T$, $T(\overline{S})\subset\overline{T(S)}$. So $T(S)\subset T(\overline{S})\subset\overline{T(S)}=T(S)$. ∎
Lemma A.2.
Let $\mathcal{H}$ be a separable Hilbert space and $\mathcal{D}\subset\mathcal{H}$ be a dense subspace and $(e_{i})_{i=1}^{m}$ be an orthonormal sequence in $\mathcal{D}$ for $m\in\mathbb{N}$. Then there exists a countable sequence $(d_{i})_{i}\subset\mathcal{D}$ such that $(d_{i})_{i}$ is an ONB of $\mathcal{H}$ and $d_{i}=e_{i}$ for $i\leq m$.
Proof.
The proof is a slight modification of the argument of [12, Lemma A.2]. By separability of $\mathcal{H}$ we may a countable and dense sequence $(h_{i})_{i}$ of $\mathcal{H}$. By density of $\mathcal{D}$ we can construct a countable sequence $(d_{i}^{\prime})_{i}\subset\mathcal{D}$ dense in $\mathcal{H}$ by taking an element of $\mathcal{D}$ from the ball $B(h_{i},1/j)$, for all $i$ and $j\in\mathbb{N}$. Now, we apply Gram-Schmidt to the countable sequence $(e_{1},\ldots,e_{m},d_{1}^{\prime},d_{2}^{\prime},\ldots)\subset\mathcal{D}$ to obtain a countable orthonormal sequence $(d_{i})_{i}\subset\mathcal{D}$. Since $(e_{i})_{i=1}^{m}$ was already orthonormal, $d_{i}=e_{i}$ for $i\leq m$. Furthermore, $d_{i}^{\prime}\in\operatorname{span}{\left(d_{j},j\leq m+i\right)}$. It follows that $(d_{i}^{\prime})_{i}\subset\operatorname{span}{\left(d_{i}\right)}$ and since $(d_{i}^{\prime})_{i}$ is dense, so is $(d_{i})_{i}$. ∎
Lemma A.3.
[7, Proposition II.2.7] Let $\mathcal{H}$ be a Hilbert space. If $A\in\mathcal{B}(\mathcal{H})$, then $\lVert A\rVert=\lVert A^{*}\rVert=\lVert AA^{*}\rVert^{1/2}$.
Lemma A.4.
[18, Theorem 4.2.6] Let $\mathcal{H}$ be a Hilbert space and $A\in\mathcal{B}(\mathcal{H})$ be compact and self-adjoint. Then $\lVert A\rVert=\max\{\lvert\lambda\rvert:\ \lambda\in\sigma_{p}(A)\}$ where $\sigma_{p}(A)\coloneqq\{\lambda:\ Av=\lambda v\text{ for some }v\not=0\}$ is the point spectrum of $A$.
Lemma A.5.
Let $\mathcal{H}$ be a Hilbert space and $A\in\mathcal{B}(\mathcal{H})$. Then $A>0$ if and only if $A\geq 0$ and $A$ is injective.
Proof.
Assume $A$ is positive. If $h\in\ker{A}$, then $\langle Ah,h\rangle_{H}=0$, so $h=0$. Now assume $A$ is nonnegative and injective. If $\langle Ah,h\rangle=\lVert A^{1/2}h\rVert^{2}=0$ for $h\not=0$, then $h\in\ker{A}^{1/2}\subset\ker{A}$, so $h=0$. ∎
Lemma A.6.
[18, Theorem 4.3.1] Let $\mathcal{H},\mathcal{K}$ be Hilbert spaces, and $A\in\mathcal{B}(\mathcal{H},\mathcal{K})$ be compact and normal. Then $A$ is diagonalisable, that is, there exists an ONB $(h_{i})_{i}$ of $\mathcal{H}$ and an orthonormal sequence $(k_{i})_{i}$ of $\mathcal{K}$ and a sequence $(\lambda_{i})_{i}$ in $\mathbb{C}$ such that $A=\sum_{i}\lambda_{i}k_{i}\otimes h_{i}$.
Lemma A.7.
[7, Proposition VI.1.8] Let $\mathcal{H}$, $\mathcal{K}$ be Hilbert spaces and $A\in\mathcal{B}(\mathcal{H},\mathcal{K})$. Then $\ker{A}=\operatorname{ran}{A^{*}}^{\perp}$ and $\ker{A}^{\perp}=\overline{\operatorname{ran}{A}}$.
Lemma A.8.
Let $\mathcal{H}$, $\mathcal{K}$ be Hilbert spaces and $A\in\mathcal{B}_{00}(\mathcal{H},\mathcal{K})$. Then $\operatorname{rank}\left({AA^{*}}\right)=\operatorname{rank}\left({A}\right)$.
Proof.
Since $\operatorname{ran}{A^{*}}$ is closed, by Lemma A.7, $\operatorname{ran}{AA^{*}}=\operatorname{ran}{AP_{\operatorname{ran}{A^{*}}}}=\operatorname{ran}{AP_{\overline{\operatorname{ran}{A^{*}}}}}=\operatorname{ran}{AP_{\ker{A}^{\perp}}}=A$, where $P_{V}$ denotes the projection onto closed $V\subset\mathcal{H}$. ∎
Definition A.9.
Let $\mathcal{H}$ be Hilbert. A linear operator $A:\operatorname{dom}{A}\subset\mathcal{H}\rightarrow\mathcal{H}$ is said to be closed if its graph $\{(h,Ah):\ h\in\operatorname{dom}{A}\}$ is closed in $\mathcal{H}\oplus\mathcal{H}$.
Lemma A.10.
Let $\mathcal{H}$ be Hilbert, $A:\operatorname{dom}{A}\subset\mathcal{H}\rightarrow\mathcal{H}$ be closed and $B\in\mathcal{B}(\mathcal{H})$, then $AB$ is closed, $A+B$ is closed, and if $A$ is bijective, then $A^{-1}$ is closed.
Proof.
If $(h_{n},ABh_{n})\rightarrow(h,k)\in\mathcal{H}\oplus\mathcal{H}$, then $(Bh_{n},ABh_{n})\rightarrow(Bh,k)$ by continuity of $B$. Since $A$ is closed, $Bh\in\operatorname{dom}{A}$, that is, $h\in\operatorname{dom}{AB}$, and $k=ABh$. Next, if $(h_{n},Ah_{n}+Bh_{n})\rightarrow(h,k)\in\mathcal{H}\oplus\mathcal{H}$, then $Bh_{n}\rightarrow z$ for some $z\in\mathcal{H}$ by continuity of $B$, and $(h_{n},Ah_{n})\rightarrow(h,k-z)$. Since $A$ is closed, $h\in\operatorname{dom}{A}=\operatorname{dom}{A+B}$ and $Ah=k-z=k-Bh$. Finally, we have $\{(h,Ah):\ h\in\operatorname{dom}{A}\}=\{(A^{-1}k,k):\ k\in\mathcal{H}\}$, and this set is closed if and only if $\{(k,A^{-1}k):\ k\in\mathcal{H}\}$ is closed. ∎
Lemma A.11 (closed graph theorem).
Let $\mathcal{H}$ be a Banach space. If $A:\operatorname{dom}{A}\subset\mathcal{H}\rightarrow\mathcal{H}$ and $\operatorname{dom}{A}=\mathcal{H}$, then $A$ is continuous if and only if $A$ is closed.
Proof.
It follows by definition of continuity that $A$ is closed for $A\in\mathcal{B}(\mathcal{H})$. For the converse, see [7, Theorem III.12.6]. ∎
Definition A.12.
Let $\mathcal{H},\mathcal{K}$ be a separable Hilbert space and $A:\mathcal{H}\rightarrow\mathcal{K}$ a densely defined linear operator on $\mathcal{H}$. Then we define
$\operatorname{dom}{A^{*}}:=\{k\in K:\ h\mapsto\langle Ah,k\rangle\text{ is a bounded linear functional on }\operatorname{dom}{A}\}.$
As $\operatorname{dom}{A}\subset\mathcal{H}$ dense, if $k\in\mathcal{K}$, there exists by the Riesz representation theorem a $f\in\mathcal{H}$ such that $\langle Ah,k\rangle=\langle h,f\rangle$ for all $h\in\mathcal{H}$. We define $A^{*}:\operatorname{dom}{A^{*}}\rightarrow\mathcal{H}$ by setting $A^{*}k=f$.
Lemma A.13.
Let $\mathcal{H}$ be separable Hilbert. If $A,B:\mathcal{H}\rightarrow\mathcal{H}$ are densely defined, then
  1. (i)
    $(AB)^{*}\supset B^{*}A^{*}$
  2. (ii)
    If $B^{*}A^{*}$ is bounded, then $(AB)^{*}=B^{*}A^{*}$.
Proof.
Let $k\in\operatorname{dom}{B^{*}A^{*}}$ and $h\in\operatorname{dom}{AB}$. Since $k\in\operatorname{dom}{A^{*}}$ and $Bh\in\operatorname{dom}{A}$, $\langle ABh,k\rangle=\langle Bh,A^{*}k\rangle.$ Since $A^{*}k\in\operatorname{dom}{B^{*}}$ and $h\in\operatorname{dom}{B}$, $\langle Bh,A^{*}k\rangle=\langle h,B^{*}A^{*}k\rangle.$ Thus $\langle ABh,k\rangle=\langle h,B^{*}A^{*}k\rangle$. Hence $h\mapsto\langle ABh,k\rangle$ is bounded and $(AB)^{*}k=B^{*}A^{*}k$, proving $(i)$. If $B^{*}A^{*}\in\mathcal{B}(\mathcal{H})$, then $\operatorname{dom}{(AB)^{*}}\subset\mathcal{H}=\operatorname{dom}{B^{*}A^{*}}$, showing $(ii)$. ∎
Definition A.14.
Let $\mathcal{H}$ be a separable Hilbert. A densely defined operator $A:\mathcal{H}\rightarrow\mathcal{H}$ is said to be symmetric if $\langle Ah,k\rangle=\langle h,Ak\rangle$ for all $h,k\in\operatorname{dom}{A}$. If $A=A^{*}$, then $A$ is said to be self-adjoint.
Remark A.15.
Note that $A=A^{*}$ if and only if $A$ is symmetric and additionally $\operatorname{dom}{A}=\operatorname{dom}{A^{*}}$ holds.
Lemma A.16.
[7, Proposition X.2.4] Let $H$ be a separable Hilbert space and $A$ a symmetric operator on $\mathcal{H}$.
  1. (i)
    If $\operatorname{ran}{A}$ is dense, then $A$ is injective.
  2. (ii)
    If $A=A^{*}$ and $A$ is injective, then $\operatorname{ran}{A}$ is dense and $A^{-1}$ is well-defined on $\operatorname{ran}{A}$ and self-adjoint.
  3. (iii)
    If $\operatorname{dom}{A}=\mathcal{H}$, then $A=A^{*}$ and $A\in\mathcal{B}(\mathcal{H})$.
  4. (iv)
    If $\operatorname{ran}{A}=\mathcal{H}$, then $A=A^{*}$ and $A^{-1}\in\mathcal{B}(\mathcal{H})$.
Lemma A.17.
Let $\mathcal{H}$ be a Hilbert space and $A:\mathcal{H}\rightarrow\mathcal{H}$ have a bounded inverse $A^{-1}\in\mathcal{B}(\mathcal{H})$. Then $\sigma(A)=\sigma(A^{-1})^{-1}$.
Proof.
Let $\lambda\in\mathbb{C}$. Then $A-\lambda^{-1}I=\lambda^{-1}(\lambda A-I)=-\lambda^{-1}A(A^{-1}-\lambda I).$ Thus $A-\lambda^{-1}I$ has a bounded inverse if and only if $A^{-1}-\lambda I$ has a bounded inverse. ∎
Lemma A.18.
Let $\mathcal{H}$ be a Hilbert space and $A:\mathcal{H}\rightarrow\mathcal{H}$ injective. Then the point spectrum $\sigma_{p}(A)\coloneqq\{\lambda:\ Av=\lambda v\text{ for some }v\not=0\}$ satisfies $\sigma_{p}(A)^{-1}\coloneqq\{\lambda^{-1}:\ \lambda\in\sigma_{p}(A)\}=\sigma_{p}(A^{-1})$.
Proof.
Since $A$ is injective, $0\not\in\sigma_{p}(A)$ and $A^{-1}:\operatorname{ran}{A}\rightarrow\mathcal{H}$ is well-defined. Furthermore, $(A-\lambda)v=0$ if and only if $(A^{-1}-\lambda^{-1})v=0$ for $\lambda\in\sigma_{p}(A)$. Thus, $\lambda\in\sigma_{p}(A)$ if and only if $\lambda^{-1}\in\sigma_{p}(A^{-1})$. ∎
Lemma A.19.
Let $\mathcal{H}$ be a separable Hilbert space and $\mathcal{C}_{1},\mathcal{C}_{2}\in L_{1}(\mathcal{H})_{\mathbb{R}}$ be nonnegative. If $\operatorname{ran}{\mathcal{C}_{1}^{1/2}}\subset\mathcal{H}$ densely, then the following holds.
  1. (i)
    $\mathcal{C}_{1}>0$ and $\mathcal{C}_{1}^{1/2}>0$.
  2. (ii)
    $\mathcal{C}_{1}^{-1/2}:\operatorname{ran}{\mathcal{C}_{1}^{1/2}}\rightarrow\mathcal{H}$ and $\mathcal{C}_{1}^{-1}:\operatorname{ran}{\mathcal{C}_{1}}\rightarrow\mathcal{H}$ are bijective and self-adjoint operators that are unbounded if $\dim{\mathcal{H}}$ is unbounded.
  3. (iii)
    If $\operatorname{ran}{\mathcal{C}_{1}^{1/2}}=\operatorname{ran}{\mathcal{C}_{2}^{1/2}}$, then $\mathcal{C}_{2}^{-1/2}\mathcal{C}_{1}^{1/2}\in\mathcal{B}(\mathcal{H})$.
Proof.
By Lemma A.16 part (i), $\mathcal{C}_{1}^{1/2}$ and hence $\mathcal{C}_{1}$ are injective. So (i) holds. By part (ii) of the same lemma, $\mathcal{C}_{1}^{-1/2}$ and $\mathcal{C}_{1}^{-1}$ are bijective and self-adjoint. The inverse of a compact operator on its range in an infinite-dimensional space is unbounded, hence (ii) holds. If $\operatorname{ran}{\mathcal{C}_{1}^{1/2}}=\operatorname{ran}{\mathcal{C}_{2}^{1/2}}$, then $\mathcal{C}_{2}^{-1/2}\mathcal{C}_{1}^{1/2}$ is well-defined, and by Lemma A.10, $\mathcal{C}_{2}^{-1/2}\mathcal{C}_{1}^{1/2}$ is a closed operator, and hence bounded by Lemma A.11, showing (iii). ∎
Condition $(i)$ of the Feldman–Hajek theorem Theorem 3.1 can be stated equivalently as follows.
Lemma A.20.
Let $\mathcal{H}$ be a Hilbert space and $\mu=N(m_{1},\mathcal{C}_{1})$ and $\nu=(m_{2},\mathcal{C}_{2})$ be Gaussian measures on $\mathcal{H}$ such that $\ker{\mathcal{C}_{1}^{1/2}}=\{0\}=\ker{\mathcal{C}_{2}^{1/2}}$. Then $\operatorname{ran}{\mathcal{C}_{1}^{1/2}}=\operatorname{ran}{\mathcal{C}_{2}^{1/2}}$ if and only if $\mathcal{C}_{2}^{-1/2}\mathcal{C}_{1}^{1/2}$ is a well-defined invertible operator in $\mathcal{B}(\mathcal{H})$.
Proof.
Suppose that $\operatorname{ran}{\mathcal{C}_{1}^{1/2}}=\operatorname{ran}{\mathcal{C}_{2}^{1/2}}$. Then $\mathcal{C}_{1}^{-1/2}\mathcal{C}_{2}^{1/2}$ is well-defined and bijective. By Lemma A.16 part $(ii)$, $\mathcal{C}_{1}^{-1/2}$ is self-adjoint, hence closed. By Lemma A.10, $\mathcal{C}_{1}^{-1/2}\mathcal{C}_{2}^{1/2}$ is closed, and by Lemma A.11, it is bounded. Conversely, if $\mathcal{C}_{1}^{-1/2}\mathcal{C}_{2}^{1/2}\in\mathcal{B}(H)$ is invertible, then,
$\operatorname{ran}{\mathcal{C}_{1}^{1/2}}$ =$=\{\mathcal{C}_{1}^{1/2}h:h\in\mathcal{H}\}=\{\mathcal{C}_{1}^{1/2}\mathcal{C}_{1}^{-1/2}\mathcal{C}_{2}^{1/2}h:h\in\mathcal{H}\}$
=$=\{\mathcal{C}_{2}^{1/2}h:h\in\mathcal{H}\}=\operatorname{ran}{\mathcal{C}_{2}^{1/2}}.$
Lemma A.21.
Let $\mathcal{H}$ be a seperable Hilbert space and $\mu=N(0,\mathcal{C})$ a Gaussian measure on $\mathcal{H}$. If $X\sim\mu$ and $\mathcal{C}=SS^{*}$ for $S\in L_{2}(\mathcal{H})$. Then $\mathbb{E}\lVert X\rVert^{2}=\lVert S^{*}\rVert_{L_{2}(\mathcal{H})}=\lVert S\rVert_{L_{2}(\mathcal{H})}^{2}.$
Proof.
Let $(e_{i})_{i}$ an ONB of $\mathcal{H}$ and $X=\sum_{i}\langle X,e_{i}\rangle e_{i}$. Then by Tonelli’s theorem, the definition of the covariance operator, the hypothesis that $\mathcal{C}=SS^{*}$, and the invariance of the Hilbert–Schmidt norm under adjoints,
$\mathbb{E}\lVert X\rVert^{2}$ $=\mathbb{E}\sum_{i}\lvert\langle X,e_{i}\rangle\rvert^{2}=\sum_{i}\mathbb{E}\lvert\langle X,e_{i}\rangle\rvert^{2}=\sum_{i}\langle\mathcal{C}e_{i},e_{i}\rangle=\sum_{i}\lVert S^{*}e_{i}\rVert^{2}=\lVert S^{*}\rVert_{L_{2}(\mathcal{H})}^{2}=\lVert S\rVert_{L_{2}(\mathcal{H})}^{2}.$
Lemma A.22.
If $\mathcal{H}_{1},\mathcal{H}_{2}$ are seperable Hilbert spaces, $X\sim\mu=N(\mu,\mathcal{C})$ Gaussian distribution on $\mathcal{H}_{1}$ and $A\in\mathcal{B}(\mathcal{H}_{1},\mathcal{H}_{2})$, then the distribution of $AX$ is $N(Am,ACA^{*})$
Lemma A.23.
Let $\mathcal{H}$ be a Hilbert space, $(e_{i})_{i}$ an orthonormal sequence and $(\delta_{i})_{i}\in\ell^{2}((-1,\infty))$. Then $I+\sum_{i}\delta_{i}e_{i}\otimes e_{i}$ is invertible in $\mathcal{B}(\mathcal{H})$, with inverse $I-\sum_{i}\frac{\delta_{i}}{1+\delta_{i}}e_{i}\otimes e_{i}$.
Proof.
Note that $\frac{\delta_{i}}{1+\delta_{i}}\leq c\cdot\delta_{i}$, with $c^{-1}=\inf_{i}{1+\delta_{i}}>0$, by the hypothesis that $(\delta_{i})_{i}\in\ell^{2}((-1,\infty))$. Thus, $\frac{\delta_{i}}{1+\delta_{i}}\rightarrow 0$ and $\sum_{i}\frac{\delta_{i}}{1+\delta_{i}}e_{i}\otimes e_{i}$ is well-defined. For $h\in\mathcal{H}$, we have by direct computation,
$\left(I+\sum_{i}\delta_{i}e_{i}\otimes e_{i}\right)\left(I-\sum_{i}\frac{\delta_{i}}{1+\delta_{i}}e_{i}\otimes e_{i}\right)h$ $=\left(I+\sum_{i}\delta_{i}e_{i}\otimes e_{i}\right)\sum_{i}\left(1-\frac{\delta_{i}}{1+\delta_{i}}\right)\langle h,e_{i}\rangle e_{i}$
$=\sum_{i}\left(1-\frac{\delta_{i}}{1+\delta_{i}}+\delta_{i}-\frac{\delta_{i}^{4}}{1+\delta_{i}}\right)\langle h,e_{i}\rangle e_{i}$
$=\sum_{i}\langle h,e_{i}\rangle e_{i}=h.$
Similarly,
$\left(I-\sum_{i\in\mathbb{N}}\frac{\delta_{i}}{1+\delta_{i}}e_{i}\otimes e_{i}\right)\left(I+\sum_{j\in\mathbb{N}}\delta_{i}e_{i}\otimes e_{i}\right)h=$ $\left(I-\sum_{i\in\mathbb{N}}\frac{\delta_{i}}{1+\delta_{i}}e_{i}\otimes e_{i}\right)\sum_{j\in\mathbb{N}}\left(1+\delta_{i}\right)\langle h,e_{j}\rangle e_{j}$
$=$ $\sum_{i\in\mathbb{N}}\left(1+\delta_{i}-\frac{\delta_{i}}{1+\delta_{i}}(1+\delta_{i})\right)\langle h,e_{i}\rangle e_{i}.$
Appendix B Proofs of results
B.1 Proofs of Section 3
See 3.3
Proof of Lemma 3.3.
The statements about extensions of operators follow from Lemma A.13.
By the Feldman–Hajek theorem, Theorem 3.1, $\operatorname{ran}{\mathcal{C}_{1}^{1/2}}=\operatorname{ran}{\mathcal{C}_{2}^{1/2}}$. Thus, by Lemma A.20, $A:=\mathcal{C}_{1}^{-1/2}\mathcal{C}_{2}^{1/2}$ is a well-defined bounded and invertible operator, and by Theorem 3.1(iii), $AA^{*}-I$ is Hilbert–Schmidt. That is, there exists a sequence $(\lambda_{i})_{i}\subset\ell^{2}(\mathbb{R})$ and ONB $(w_{i})_{i}$ of $\mathcal{H}$ such that,
$AA^{*}-I=\sum_{i}\lambda_{i}w_{i}\otimes w_{i},$
i.e.,
$AA^{*}w_{i}=(1+\lambda_{i})w_{i}.$ (32)
As $A$ is invertible in $\mathcal{B}(H)$, so are $A^{*}$ and $AA^{*}$. But $AA^{*}>0$, so $AA^{*}-I>-1$, showing that $\lambda_{i}>-1$ for all $i$, which proves item (i) holds.
By multiplying (32) by $A^{-1}$, $A^{-1}(A^{-1})^{*}A^{-1}$ and $(A^{-1})^{*}A^{-1}$ and rearranging, we obtain respectively,
$A^{*}AA^{-1}w_{i}$ $=(1+\lambda_{i})A^{-1}w_{i},$ (33)
$A^{-1}(A^{-1})^{*}A^{-1}w_{i}$ $=\frac{1}{1+\lambda_{i}}A^{-1}w_{i},$ (34)
$(A^{-1})^{*}A^{-1}w_{i}$ $=\frac{1}{1+\lambda_{i}}w_{i}.$ (35)
By (35), $v_{i}:=(1+\lambda_{i})^{1/2}A^{-1}w_{i}$ satisfies,
$\langle v_{i},v_{j}\rangle=(1+\lambda_{i})^{1/2}(1+\lambda_{j})^{1/2}\langle(A^{-1})^{*}A^{-1}w_{i},w_{j}\rangle=\delta_{ij},$
and, for all $h\in\mathcal{H}$,
$\sum_{i}\langle A^{-1}h,v_{i}\rangle v_{i}$ $=\sum_{i}(1+\lambda_{i})\langle h,(A^{-1})^{*}A^{-1}w_{i}\rangle A^{-1}w_{i}$
$=A^{-1}\sum_{i}\langle h,w_{i}\rangle w_{i}$
$=A^{-1}h,$
where we used that $A^{-1}$ is continuous and $(w_{i})_{i}$ is an ONB. Hence, $(v_{i})_{i}$ is an ONB. Now, (33), (34) and (35) become
$(A^{*}A-I)v_{i}$ $=\lambda_{i}v_{i},$
$(A^{-1}(A^{-1})^{*}-I)v_{i}$ $=\frac{-\lambda_{i}}{1+\lambda_{i}}v_{i},$
$((A^{-1})^{*}A^{-1}-I)w_{i}$ $=\frac{-\lambda_{i}}{1+\lambda_{i}}w_{i}.$
Notice that $\frac{-\lambda_{i}}{1+\lambda_{i}}\in\ell^{2}((-1,\infty))$, since $1+\lambda_{i}\rightarrow 1$ and $(\lambda_{i})_{i}\in\ell^{2}((-1,\infty))$. This proves items (ii), (iii) and (iv). ∎
See 3.6
Proof of Lemma 3.6.
By Lemma 3.3 with $\mathcal{C}_{1}\leftarrow\mathcal{C}_{\textup{pr}}$ and $\mathcal{C}_{2}\leftarrow\mathcal{C}_{\textup{pos}}$, there exists an eigenvalue sequence $(\lambda_{i})_{i\in\mathbb{N}}\subset\ell^{2}((-1,\infty))$ and ONBs $(w_{i})_{i\in\mathbb{N}}$ and $(v_{i})_{i}$ of $\mathcal{H}$ such that $v_{i}=\sqrt{1+\lambda_{i}}\mathcal{C}_{\textup{pos}}^{-1/2}\mathcal{C}_{\textup{pr}}^{1/2}w_{i}$ and items (i), (ii), (iii) and (iv) of Lemma 3.3 hold. By (3c), it holds on $\operatorname{ran}{\mathcal{C}_{\textup{pr}}^{1/2}}$,
$\mathcal{C}_{\textup{pr}}^{1/2}\mathcal{C}_{\textup{pos}}^{-1}\mathcal{C}_{\textup{pr}}^{1/2}-I=\mathcal{C}_{\textup{pr}}^{1/2}(\mathcal{C}_{\textup{pr}}^{-1}+H)\mathcal{C}_{\textup{pr}}^{1/2}-I=\mathcal{C}_{\textup{pr}}^{1/2}\mathcal{C}_{\textup{pr}}^{-1}\mathcal{C}_{\textup{pr}}^{1/2}+\mathcal{C}_{\textup{pr}}^{1/2}H\mathcal{C}_{\textup{pr}}^{1/2}-I.$
Now $\mathcal{C}_{\textup{pr}}^{1/2}H\mathcal{C}_{\textup{pr}}^{1/2}-I\in\mathcal{B}(\mathcal{H})$, hence it is defined on all of $\mathcal{H}$. The operator $\mathcal{C}_{\textup{pr}}^{1/2}\mathcal{C}_{\textup{pr}}^{-1}\mathcal{C}_{\textup{pr}}^{1/2}$ is extended by the identity operator $I$. Thus, $\mathcal{C}_{\textup{pr}}^{1/2}\mathcal{C}_{\textup{pos}}^{-1}\mathcal{C}_{\textup{pr}}^{1/2}-I\subset\mathcal{C}_{\textup{pr}}^{1/2}H\mathcal{C}_{\textup{pr}}^{1/2}.$ By the uniqueness of extensions of continuous functions on the dense set $\operatorname{ran}{\mathcal{C}_{\textup{pr}}}^{1/2}$, this implies together with item (iv) of Lemma 3.3 that (10a) holds. The proof of (10b) is similar: by (3c), it holds on $\operatorname{ran}{\mathcal{C}_{\textup{pos}}^{1/2}}=\operatorname{ran}{\mathcal{C}_{\textup{pr}}^{1/2}}$,
$\mathcal{C}_{\textup{pos}}^{1/2}\mathcal{C}_{\textup{pr}}^{-1}\mathcal{C}_{\textup{pos}}^{1/2}-I=\mathcal{C}_{\textup{pos}}^{1/2}(\mathcal{C}_{\textup{pr}}^{-1}-\mathcal{C}_{\textup{pos}}^{-1})\mathcal{C}_{\textup{pos}}^{1/2}=\mathcal{C}_{\textup{pos}}^{1/2}(-H)\mathcal{C}_{\textup{pos}}^{1/2}.$
and combining this with item (ii) of Lemma 3.3, (10b) follows by uniqueness of the extension.
We now prove the stated properties of the eigenvalues and eigenvectors. Recall that by (2), $H\in\mathcal{B}_{00,n}(\mathcal{H})$ is self-adjoint and non-negative. Hence $\mathcal{C}_{\textup{pr}}^{1/2}H\mathcal{C}_{\textup{pr}}^{1/2}=(\mathcal{C}_{\textup{pr}}^{1/2}H^{1/2})(\mathcal{C}_{\textup{pr}}^{1/2}H^{1/2})^{*}$ is also self-adjoint and non-negative, which implies that $(\tfrac{-\lambda_{i}}{1+\lambda_{i}})_{i\in\mathbb{N}}\subset\ell^{2}((-1,0])$, and thus that $(\lambda_{i})_{i\in\mathbb{N}}\in\ell^{2}((-1,0])$. Since $\mathcal{C}_{\textup{pr}}$ is injective on $\mathcal{H}$, it follows that $\operatorname{rank}\left({\mathcal{C}_{\textup{pr}}^{1/2}H\mathcal{C}_{\textup{pr}}^{1/2}}\right)=\operatorname{rank}\left({H}\right)$, whence $(\tfrac{-\lambda_{i}}{1+\lambda_{i}})_{i\in\mathbb{N}}$ contains exactly $\operatorname{rank}\left({H}\right)\leq n$ many nonzero entries. It follows directly from (10a), (10b) and the fact that $\lambda_{i}\not=0$ for $i\leq\operatorname{rank}\left({H}\right)$, that $w_{i}\in\operatorname{ran}{\mathcal{C}_{\textup{pr}}^{1/2}H\mathcal{C}_{\textup{pr}}^{1/2}}=\operatorname{ran}{\mathcal{C}_{\textup{pr}}^{1/2}}\subset\operatorname{ran}{\mathcal{C}_{\textup{pr}}^{1/2}}$ and $v_{i}\in\operatorname{ran}{\mathcal{C}_{\textup{pos}}^{1/2}H\mathcal{C}_{\textup{pos}}^{1/2}}\subset\operatorname{ran}{\mathcal{C}_{\textup{pos}}^{1/2}}=\operatorname{ran}{\mathcal{C}_{\textup{pr}}^{1/2}}$ for $i\leq\operatorname{rank}\left({H}\right)$. By Lemma A.2, we can extend $(w_{i})_{i=1}^{\operatorname{rank}\left({H}\right)}$ to an ONB $(w_{i}^{\prime})_{i}$ of $\mathcal{H}$ with $(w_{i}^{\prime})_{i}\subset\operatorname{ran}{\mathcal{C}_{\textup{pr}}^{1/2}}$ and $w_{i}^{\prime}=w_{i}$ for $i\leq\operatorname{rank}\left({H}\right)$. We now replace $w_{i}$ by $w_{i}^{\prime}$ and $v_{i}$ by $\mathcal{C}_{\textup{pos}}^{-1/2}\mathcal{C}_{\textup{pr}}^{1/2}w_{i}^{\prime}$ for $i>\operatorname{rank}\left({H}\right)$. The equations (10a) and (10b) and $v_{i}=\sqrt{1+\lambda_{i}}\mathcal{C}_{\textup{pos}}^{-1/2}\mathcal{C}_{\textup{pr}}^{1/2}w_{i}$ for all $i$ remain to hold, and we now have $w_{i},v_{i}\in\operatorname{ran}{\mathcal{C}_{\textup{pr}}^{1/2}}$ for all $i$.
By (i) of Lemma 3.3 and the fact that $(w_{i})_{i}$ lies in the Cameron–Martin space, it follows that
$\mathcal{C}_{\textup{pr}}^{-1/2}\mathcal{C}_{\textup{pos}}\mathcal{C}_{\textup{pr}}^{-1/2}w_{i}=(1+\lambda_{i})w_{i},\quad i\in\mathbb{N}.$
Multiplying both sides of the equation from the left by $\mathcal{C}_{\textup{pos}}^{-1/2}\mathcal{C}_{\textup{pr}}^{1/2}$ yields (10c).
Finally, we prove (11). Multiplying the equation (10c) from the left by $\mathcal{C}_{\textup{pos}}^{1/2}$ implies that $\mathcal{C}_{\textup{pos}}\mathcal{C}_{\textup{pr}}^{-1/2}w_{i}=(1+\lambda_{i})\mathcal{C}_{\textup{pr}}^{1/2}w_{i}=(1+\lambda_{i})\mathcal{C}_{\textup{pr}}(\mathcal{C}_{\textup{pr}}^{-1/2}w_{i})$. Taking the inner product of both sides of the last equation with $\mathcal{C}_{\textup{pr}}^{-1/2}w_{i}$, we obtain $\langle\mathcal{C}_{\textup{pos}}\mathcal{C}_{\textup{pr}}^{-1/2}w_{i},\mathcal{C}_{\textup{pr}}^{-1/2}w_{i}\rangle=(1+\lambda_{i})\langle\mathcal{C}_{\textup{pr}}\mathcal{C}_{\textup{pr}}^{-1/2}w_{i},\mathcal{C}_{\textup{pr}}^{-1/2}w_{i}\rangle$. By Lemma A.22, $\textup{Var}_{X\sim\mu_{\textup{pos}}}(\langle X,z\rangle)=\langle\mathcal{C}_{\textup{pos}}z,z\rangle$ and $\textup{Var}_{X\sim\mu_{\textup{pr}}}(\langle X,z\rangle)=\langle\mathcal{C}_{\textup{pr}}z,z\rangle$ for any $z\in\mathcal{H}$. Thus we obtain (11). ∎
See 3.8
Proof of Theorem 3.8.
We use the expressions for the KL and Rényi divergence of [22, Theorem 14, Theorem 15]. While they are stated for infinite-dimensional Hilbert spaces only, it is noted in [23, after Theorem 3] that these expressions also hold for finite-dimensional Hilbert spaces. The statements in the theorem now follow immediately from the expressions in [22, Theorem 14, Theorem 15] after noting that, since $(\mathcal{C}_{1}^{-1/2}\mathcal{C}_{2}^{1/2})^{*}=\mathcal{C}_{2}^{1/2}\mathcal{C}_{1}^{-1/2}$ on $\operatorname{ran}{\mathcal{C}_{1}^{1/2}}$ by Lemma A.13(i), we have for $S\coloneqq-R(\mathcal{C}_{2}\|\mathcal{C}_{1})\in L_{2}(\mathcal{H})_{\mathbb{R}}$, where $R(\cdot\|\cdot)$ is defined in (9), that
$\mathcal{C}_{1}^{1/2}(I-S)\mathcal{C}_{1}^{1/2}=\mathcal{C}_{1}^{1/2}(\mathcal{C}_{1}^{-1/2}\mathcal{C}_{2}^{1/2})(\mathcal{C}_{1}^{-1/2}\mathcal{C}_{2}^{1/2})^{*}\mathcal{C}_{1}^{1/2}=\mathcal{C}_{1}^{1/2}\mathcal{C}_{1}^{-1/2}\mathcal{C}_{2}\mathcal{C}_{1}^{-1/2}\mathcal{C}_{1}^{1/2}=\mathcal{C}_{2},$
and $I-(1-\rho)S=\rho I+(1-\rho)(I+R(\mathcal{C}_{2}\|\mathcal{C}_{1}))$ for $10\leq\rho<$ 1. ∎
B.2 Proofs of Section 4
See 4.1
Proof of Lemma 4.1.
Given that $xf^{\prime}(x)>0$ for $x\neq 0$, it follows that $f^{\prime}(x)<0$ for $x<0$ and $f^{\prime}(x)>0$ for $x>0$. This implies, by continuity of $f$, that $f^{\prime}(0)=0$. Hence the $f$ has a global minimum only at $x=0$ and $f\geq 0$. Thus, also $\mathcal{L}_{f}\geq 0$. By the Lipschitz assumption on $f^{\prime}$ at 0, there exists $\varepsilon\in(0,1)$ and $M_{0}>0$ such that $f^{\prime}(x)=f^{\prime}(x)-f^{\prime}(0)\leq M_{0}\lvert x\rvert$ for $\lvert x\rvert\leq\varepsilon$. For $\lvert y\rvert\leq\varepsilon$,
$f(y)=f(y)-f(0)=\int_{0}^{y}f^{\prime}(x)\operatorname{d}\!{x}\leq\int_{0}^{y}M_{0}\lvert x\rvert\operatorname{d}\!{x}=\frac{1}{2}M_{0}y^{2}.$
Let $(x_{i})_{i}\in\ell^{2}((-1,\infty))$. For $N$ large enough, its tail $(x_{i})_{i>N}$ lies in $(-\varepsilon,\varepsilon)$, so that the inequality above implies $\sum_{i>N}f(x_{i})\leq\tfrac{1}{2}M\sum_{i>N}x_{i}^{2}\leq\tfrac{1}{2}M\lVert x\rVert_{\ell^{2}}^{2}$. For $\mathcal{C}_{1}$, $\mathcal{C}_{2}\in\mathcal{E}$ we have $R(\mathcal{C}_{2}\|\mathcal{C}_{1})\in L_{2}(\mathcal{H})$ and its eigenvalue sequence is square summable. Hence $\mathcal{L}_{f}<\infty$ by the definition of $\mathcal{L}_{f}$ in (15a). This proves item (i). For item (ii), we note $f(\alpha(0))=f(0)=0$. Furthermore, we compute $\alpha^{\prime}(x)=-(1+x)^{-2}>0$ and, by the fact that $f\in\mathscr{F}$, $x(f\circ\alpha)^{\prime}(x)=\frac{1}{1+x}\frac{-x}{1+x}f^{\prime}(\frac{1}{1+x})>0$ for $x\not=0$. By the assumption on $f$, $\lim_{x\rightarrow\infty}f(\alpha(x))=\lim_{x\rightarrow-1}f(x)=\infty$. Finally, $\alpha$ is smooth, so $\alpha$ and $\alpha^{\prime}$ are Lipschitz at 0. Therefore, $f^{\prime}\circ\alpha$ is Lipschitz at 0 as the composition Lipschitz functions at 0 and $(f\circ\alpha)^{\prime}=(f^{\prime}\circ\alpha)\alpha^{\prime}$ is Lipschitz at 0 as the product of two Lipschitz functions at 0. ∎
See 4.2
Proof of Lemma 4.2.
Notice that $f_{\textup{KL}}$, $f_{\textup{Ren},\rho}\in\mathcal{C}^{\infty}(\mathbb{R})$, which implies $f_{\textup{KL}}$ and $f_{\textup{Ren},\rho}$ are locally Lipschitz on $(-1,\infty)$. We compute $xf_{\textup{KL}}^{\prime}(x)=\frac{x^{2}}{2(1+x)}>0$ for $x\not=0$ and
$f^{\prime}_{\textup{Ren},\rho}(x)=$ $\frac{1}{2\rho(1-\rho)}\left[\frac{\rho-1}{1+x}+\frac{1-\rho}{\rho+(1-\rho)(1+x)}\right]=\frac{x}{2(1+x)[\rho+(1-\rho)(1+x)]}.$
Hence $xf^{\prime}_{\textup{Ren},\rho}(x)>0$ for $x\not=0$. Furthermore, $f_{\textup{KL}}(0)=0=f_{\textup{Ren},\rho}(0)$ and $\lim_{x\rightarrow\infty}f_{\textup{KL}}(x)=\infty=\lim_{x\rightarrow\infty}f_{\textup{Ren},\rho}(x)$. So $f_{\textup{KL}},f_{\textup{Ren},\rho}\in\mathscr{F}$.
The first equations in items (i) and (ii) follows from Theorem 3.8. With $(\lambda_{i})_{i\in\mathbb{N}}$ the eigenvalues of $R(\mathcal{C}_{2}\|\mathcal{C}_{1})\in L_{2}(\mathcal{H})$, it holds by (13) that
$\det_{2}(I+R(\mathcal{C}_{2}\|\mathcal{C}_{1}))=\prod_{i\in\mathbb{N}}(1+\lambda_{i})\exp(-\lambda_{i})=\prod_{i\in\mathbb{N}}\exp(-2f_{\textup{KL}}(\lambda_{i}))=\exp\biggr{(}-2\sum_{i\in\mathbb{N}}f_{\textup{KL}}(\lambda_{i})\biggr{)},$
which proves that $D_{\textup{KL}}(\nu||\mu)=\sum_{i\in\mathbb{N}}f_{\textup{KL}}(\lambda_{i})=\mathcal{L}_{f_{\textup{KL}}}(\mathcal{C}_{2}\|\mathcal{C}_{1})$. Hence item (i) holds.
By the spectral mapping theorem, see e.g. [26, Theorem VII.1(e)] for a version that does not assume that $\mathcal{H}$ is defined over the complex field $\mathbb{C}$, the eigenvalues of $I+R(\mathcal{C}_{2}\|\mathcal{C}_{1})$ are $(1+\lambda_{i})_{i}$, and the eigenvalues of $A_{\rho}\coloneqq\bigl{(}I+R(\mathcal{C}_{2}\|\mathcal{C}_{1})\bigr{)}^{\rho-1}\bigl{(}\rho I+(1-\rho)(I+R(\mathcal{C}_{2}\|\mathcal{C}_{1}))\bigr{)}$ are $(\gamma_{i})_{i}$, with $\gamma_{i}\coloneqq(1+\lambda_{i})^{\rho-1}(\rho+(1-\rho)(1+\lambda_{i}))$, $i\in\mathbb{N}$. The eigenvalues of $A_{\rho}-I$ are then $(\gamma_{i}-1)_{i}$ and by (12),
$\det(A_{\rho})=\det(I+(A_{\rho}-I))=\prod_{i}(1+(\gamma_{i}-1))=\exp(\sum_{i}\log(\gamma_{i}))=\exp\left(2\rho(1-\rho)\sum_{i}f_{\textup{Ren},\rho}(\lambda_{i})\right).$
This shows item (ii) holds. Item (iii) now follows directly from items (i) and (ii) and Lemma 4.1(ii). ∎
See 4.5
Proof of Lemma 4.5.
Let $r\in\mathbb{N}$. Item (ii) follows from item (i): since $\mathscr{C}_{r}^{-1}\coloneqq\{\mathcal{C}^{-1}:\ \mathcal{C}\in\mathscr{C}_{r}\}=\mathscr{P}_{r}$, we have
$\min\{\mathcal{L}(\mathcal{C}_{\textup{pos}}\|\mathcal{P}^{-1}):\ \mathcal{P}\in\mathscr{P}_{r}\}=\min\{\mathcal{L}\bigr{(}\mathcal{C}_{\textup{pos}}\|(\mathcal{C}^{-1})^{-1}\bigr{)}:\ \mathcal{C}^{-1}\in\mathscr{P}_{r}\}=\min\{\mathcal{L}(\mathcal{C}_{\textup{pos}}\|\mathcal{C}):\ \mathcal{C}\in\mathscr{C}_{r}\}.$
To prove item (i), we first prove the forward correspondence and show that $\mathscr{C}_{r}^{-1}\subset\mathscr{P}_{r}$. Let $K\in\mathcal{B}(\mathbb{R}^{r},\mathcal{H})$ such that $\mathscr{C}_{\textup{pr}}-KK^{*}\in\mathcal{C}_{r}$. This means that $\operatorname{ran}{K}\subset\operatorname{ran}{\mathcal{C}_{\textup{pr}}}$ and $\mathcal{C}_{\textup{pr}}-KK^{*}$ is injective and has nonnegative spectrum. The range condition $\operatorname{ran}{K}\subset\operatorname{ran}{\mathcal{C}_{\textup{pr}}}$ implies that $\mathcal{C}_{\textup{pr}}^{-1/2}K\in\mathcal{B}(\mathbb{R}^{r},\mathcal{H})$ is a well-defined operator and $\operatorname{ran}{\mathcal{C}_{\textup{pr}}^{-1/2}K}\subset\operatorname{ran}{\mathcal{C}_{\textup{pr}}^{1/2}}$. Because $\mathcal{C}_{\textup{pr}}^{-1/2}$ is injective, we have $\operatorname{rank}\left({\mathcal{C}_{\textup{pr}}^{-1/2}K}\right)=\operatorname{rank}\left({K}\right)$. Since $K^{*},\mathcal{C}_{\textup{pr}}^{-1/2}$ and $K^{*}\mathcal{C}_{\textup{pr}}^{-1/2}$ are all densely defined, and since $\mathcal{C}_{\textup{pr}}^{-1/2}$ is self adjoint, by Lemma A.13(i), with $A\leftarrow\mathcal{C}_{\textup{pr}}^{-1/2}$, and $B\leftarrow K$ that $\mathcal{C}_{\textup{pr}}^{-1/2}KK^{*}\mathcal{C}_{\textup{pr}}^{-1/2}\subset\mathcal{C}_{\textup{pr}}^{-1/2}K(\mathcal{C}_{\textup{pr}}^{-1/2}K)^{*}$. Since $\mathcal{C}_{\textup{pr}}^{-1/2}K(\mathcal{C}_{\textup{pr}}^{-1/2}K)^{*}$ has the same rank as $K$ by Lemma A.8, and is self-adjoint and nonnegative, it follows that $\mathcal{C}_{\textup{pr}}^{-1/2}KK^{*}\mathcal{C}_{\textup{pr}}^{-1/2}$ can be extended by a diagonalisable operator, say, by $\sum_{i=1}^{\operatorname{rank}\left({K}\right)}d_{i}^{2}e_{i}\otimes e_{i}$, for some orthonormal sequence $(e_{i})_{i}\subset\mathcal{H}$ and sequence $(d_{i}^{2})_{i}$ with $d_{i}^{2}\geq d_{i+1}^{2}>0$. Now, $\mathcal{C}_{\textup{pr}}^{-1/2}KK^{*}\mathcal{C}_{\textup{pr}}^{-1/2}(\operatorname{ran}{\mathcal{C}_{\textup{pr}}^{1/2}})\subset\operatorname{ran}{\mathcal{C}_{\textup{pr}}^{1/2}}$, hence $\mathcal{C}_{\textup{pr}}^{-1/2}K(\mathcal{C}_{\textup{pr}}^{-1/2}K)^{*}$ maps $\operatorname{ran}{\mathcal{C}_{\textup{pr}}^{1/2}}$ into itself. Furthermore, $\mathcal{C}_{\textup{pr}}^{-1/2}K(\mathcal{C}_{\textup{pr}}^{-1/2}K)^{*}$ has closed range because it has finite rank. By density of $\operatorname{ran}{\mathcal{C}_{\textup{pr}}^{1/2}}\subset\mathcal{H}$ and Lemma A.1, it follows that $\operatorname{ran}{\sum_{i=1}^{\operatorname{rank}\left({K}\right)}d_{i}^{2}e_{i}\otimes e_{i}}\subset\operatorname{ran}{\mathcal{C}_{\textup{pr}}^{1/2}}.$ Thus, $(e_{i})_{i}\subset\operatorname{ran}{\mathcal{C}_{\textup{pr}}^{1/2}}$ for $i\leq\operatorname{rank}\left({K}\right)$. Since $\mathcal{C}\in\mathscr{C}_{r}$, we then see that
$\mathcal{C}_{\textup{pr}}-KK^{*}=\mathcal{C}_{\textup{pr}}^{1/2}(I-\mathcal{C}_{\textup{pr}}^{-1/2}KK^{*}\mathcal{C}_{\textup{pr}}^{-1/2})\mathcal{C}_{\textup{pr}}^{1/2}=\mathcal{C}_{\textup{pr}}^{1/2}(I-\sum_{i=1}^{\operatorname{rank}\left({K}\right)}d_{i}^{2}e_{i}\otimes e_{i})\mathcal{C}_{\textup{pr}}^{1/2},$ (36)
is injective and positive. If $d_{i}^{2}=1$ for some $i\leq\operatorname{rank}\left({K}\right)$, then $\mathcal{C}_{\textup{pr}}^{-1/2}e_{i}\in\ker{(\mathcal{C}_{\textup{pr}}-KK^{*})}$, contradicting injectivity. If $d_{i}^{2}>1$ for some $i\leq\operatorname{rank}\left({K}\right)$, then $\langle(\mathcal{C}_{\textup{pr}}-KK^{*})\mathcal{C}_{\textup{pr}}^{-1/2}e_{i},\mathcal{C}_{\textup{pr}}^{-1/2}e_{i}\rangle<0$, contradicting positivity. Thus, $d_{i}^{2}\in(0,1)$ for all $i\leq\operatorname{rank}\left({K}\right)$. Hence, $I-\sum_{i=1}^{\operatorname{rank}\left({K}\right)}d_{i}^{2}e_{i}\otimes e_{i}$ maps $\operatorname{ran}{\mathcal{C}_{\textup{pr}}^{1/2}}$ onto itself. It follows from (36) that
$\operatorname{ran}{\mathcal{C}_{\textup{pr}}-KK^{*}}=\mathcal{C}_{\textup{pr}}^{1/2}\left(I-\sum_{i=1}^{\operatorname{rank}\left({K}\right)}d_{i}^{2}e_{i}\otimes e_{i}\right)(\operatorname{ran}{\mathcal{C}_{\textup{pr}}^{1/2}})=\mathcal{C}_{\textup{pr}}^{1/2}(\operatorname{ran}{\mathcal{C}_{\textup{pr}}^{1/2}})=\operatorname{ran}{\mathcal{C}_{\textup{pr}}}.$
Because $(-d_{i}^{2})_{i}\in\ell^{2}((-1,\infty))$, we can then write the inverse that $\mathcal{C}_{\textup{pr}}-KK^{*}$ has on $\operatorname{ran}{\mathcal{C}_{\textup{pr}}}$ as, using (36) and Lemma A.23 with $\delta_{i}\leftarrow(-d_{i}^{2})$
$(\mathcal{C}_{\textup{pr}}-KK^{*})^{-1}$ $=\mathcal{C}_{\textup{pr}}^{-1/2}(I-\sum_{i=1}^{\operatorname{rank}\left({K}\right)}d_{i}^{2}e_{i}\otimes e_{i})^{-1}\mathcal{C}_{\textup{pr}}^{-1/2}$
$=\mathcal{C}_{\textup{pr}}^{-1/2}(I+\sum_{i=1}^{\operatorname{rank}\left({K}\right)}\frac{d_{i}^{2}}{1-d_{i}^{2}}e_{i}\otimes e_{i})\mathcal{C}_{\textup{pr}}^{-1/2}$
$=\mathcal{C}_{\textup{pr}}^{-1}+UU^{*},$
where
$U:=\mathcal{C}_{\textup{pr}}^{-1/2}\sum_{i=1}^{\operatorname{rank}\left({K}\right)}\frac{d_{i}}{\sqrt{1-d_{i}^{2}}}e_{i}\otimes\phi_{i}\in\mathcal{B}(\mathbb{R}^{r},H),$
for any choice of ONB $(\phi_{i})_{i=1}^{r}$ of $\mathbb{R}^{r}$. This proves that $\mathscr{C}_{r}^{-1}\subset\mathscr{P}_{r}$. Since $\mathcal{C}_{\textup{pr}}^{-1/2}$ is injective, $\operatorname{rank}\left({U}\right)=\operatorname{rank}\left({K}\right)$ and this proves the forward correspondence states in item (i).
Next, we prove $\mathscr{C}_{r}^{-1}\supset\mathscr{P}_{r}$. Let $U\in\mathcal{B}(\mathbb{R}^{r},\mathcal{H})$. Then $\mathcal{C}_{\textup{pr}}^{-1}+UU^{*}\in\mathscr{P}_{r}$. The nonnegative, self-ajoint and finite rank operator $\mathcal{C}_{\textup{pr}}^{1/2}UU^{*}\mathcal{C}_{\textup{pr}}^{1/2}$ has the same rank as $\mathcal{C}_{\textup{pr}}^{1/2}UU^{*}$ by Lemma A.1 and injectivity of $\mathcal{C}_{\textup{pr}}^{1/2}$. In turn, $\mathcal{C}_{\textup{pr}}^{1/2}UU^{*}$ has the same rank as $U$. We can therefore diagonalise $\mathcal{C}_{\textup{pr}}^{1/2}UU^{*}\mathcal{C}_{\textup{pr}}^{1/2}$ as, say, $\sum_{i=1}^{\operatorname{rank}\left({U}\right)}c_{i}^{2}f_{i}\otimes f_{i}$ with the $(f_{i})_{i}$ orthonormal, $(f_{i})_{i}\subset\operatorname{ran}{\mathcal{C}_{\textup{pr}}^{1/2}}$ for $i\leq\operatorname{rank}\left({U}\right)$, and $c_{i}^{2}\geq c_{i+i}^{2}>0$. Thus, $I+\sum_{i=1}^{\operatorname{rank}\left({U}\right)}c_{i}^{2}f_{i}\otimes f_{i}\geq I$. In particular, $I+\sum_{i=1}^{\operatorname{rank}\left({U}\right)}c_{i}^{2}f_{i}\otimes f_{i}$ is a bijection, and maps $\operatorname{ran}{\mathcal{C}_{\textup{pr}}^{1/2}}$ onto itself. Hence, as a map $\operatorname{ran}{\mathcal{C}_{\textup{pr}}}\rightarrow\mathcal{H}$, it holds that
$\mathcal{C}_{\textup{pr}}^{-1}+UU^{*}=\mathcal{C}_{\textup{pr}}^{-1/2}(I+\mathcal{C}_{\textup{pr}}^{1/2}UU^{*}\mathcal{C}_{\textup{pr}}^{1/2})\mathcal{C}_{\textup{pr}}^{-1/2}=\mathcal{C}_{\textup{pr}}^{-1/2}\left(I+\sum_{i=1}^{\operatorname{rank}\left({U}\right)}c_{i}^{2}f_{i}\otimes f_{i}\right)\mathcal{C}_{\textup{pr}}^{-1/2},$ (37)
is a bijection, since it is a composition of bijections. Since $UU^{*}\in\mathcal{B}(\mathcal{H})$, $\mathcal{C}_{\textup{pr}}^{-1}+UU^{*}$ is closed by Lemma A.10. As a closed linear bijection, $\mathcal{C}_{\textup{pr}}^{-1}+UU^{*}$ has a closed and everywhere defined inverse by Lemma A.10, which is then bounded by Lemma A.11. By (37), by the fact that $(c_{i})_{i}\in\ell^{2}((-1,\infty)$) and by Lemma A.23 with $\delta_{i}\leftarrow c_{i}^{2}$, this inverse equals,
$\left(\mathcal{C}_{\textup{pr}}^{-1}+UU^{*}\right)^{-1}$ $=\mathcal{C}_{\textup{pr}}^{1/2}(I+\sum_{i=1}^{\operatorname{rank}\left({U}\right)}c_{i}^{2}f_{i}\otimes f_{i})^{-1}\mathcal{C}_{\textup{pr}}^{1/2}$
$=\mathcal{C}_{\textup{pr}}^{1/2}(I-\sum_{i=1}^{\operatorname{rank}\left({U}\right)}\frac{c_{i}^{2}}{1+c_{i}^{2}}f_{i}\otimes f_{i})^{-1}\mathcal{C}_{\textup{pr}}^{1/2}$
$=\mathcal{C}_{\textup{pr}}-KK^{*},$
where
$K\coloneqq\mathcal{C}_{\textup{pr}}^{1/2}\sum_{i=1}^{\operatorname{rank}\left({U}\right)}\frac{c_{i}}{\sqrt{1+c_{i}^{2}}}f_{i}\otimes\phi_{i}\in\mathcal{B}(\mathbb{R}^{r},\mathcal{H}),$
for any choice of ONB $(\phi_{i}^{\prime})_{i=1}^{r}$ of $\mathbb{R}^{r}$. Since $(f_{i})_{i=1}^{\operatorname{rank}\left({U}\right)}\subset\operatorname{ran}{\mathcal{C}_{\textup{pr}}^{1/2}}$, we have $\operatorname{ran}{K}\subset\operatorname{ran}{\mathcal{C}_{\textup{pr}}}$. We now show that $\mathcal{C}_{\textup{pr}}-KK^{*}>0$. Denote by $\sigma_{p}(A)$ the point spectrum of an operator $A$, that is, the set of eigenvalues. Then by Lemma A.18, $\sigma_{p}(\mathcal{C}_{\textup{pr}}^{-1})=\sigma_{p}(\mathcal{C}_{\textup{pr}})^{-1}$ and $\sigma_{p}(\mathcal{C}_{\textup{pr}}-KK^{*})=\sigma_{p}(\mathcal{C}_{\textup{pr}}^{-1}+UU^{*})^{-1}$. Let $\lambda$ be an eigenvalue of $\mathcal{C}_{\textup{pr}}^{-1}+UU^{*}$ with eigenvector $v$. If $v\in\ker{U^{*}}$, then $\lambda\in\sigma_{p}(\mathcal{C}_{\textup{pr}}^{-1})=\sigma_{p}(\mathcal{C}_{\textup{pr}})^{-1}$ and thus $\lambda>0$ because $\mathcal{C}_{\textup{pr}}>0$. If $v\not\in\ker{U^{*}}$, then $\lambda\lVert v\rVert^{2}=\langle(\mathcal{C}_{\textup{pr}}^{-1}+UU^{*})v,v\rangle\geq\lVert U^{*}v\rVert^{2}>0$. Thus $\sigma_{p}(\mathcal{C}_{\textup{pr}}-KK^{*})=\sigma_{p}(\mathcal{C}_{\textup{pr}}^{-1}+UU^{*})^{-1}\subset(0,\infty)$. It follows that $\sigma(\mathcal{C}_{\textup{pr}}-KK^{*})=\left\{0\right\}\cup\sigma_{p}(\mathcal{C}_{\textup{pr}}-KK^{*})\subset[0,\infty)$. We thus conclude that $(\mathcal{C}_{\textup{pr}}^{-1}+UU^{*})^{-1}\in\mathscr{C}_{r}$. This implies that $\mathscr{P}_{r}^{-1}\subset\mathscr{C}_{r}$ and hence $\mathscr{P}_{r}\supset\mathscr{C}_{r}$. Finally, $\operatorname{rank}\left({K}\right)=\operatorname{rank}\left({U}\right)$ by injectivity of $\mathcal{C}_{\textup{pr}}^{1/2}$, thus proving the reverse corresponcence stated in the lemma. ∎
See 4.6
Proof of Lemma 4.6.
The second statement follows from the first statement, since $\mathcal{L}\in\mathscr{L}$ is defined on $\mathcal{E}^{2}$ by definition (15b), and since $\mathscr{P}_{r}^{-1}=\mathscr{C}_{r}$ by Lemma 4.5(i). Let $\mathcal{C}\in\mathscr{C}_{r}$ arbitrary. Thus $\mathcal{C}=\mathcal{C}_{\textup{pr}}-KK^{*}$ for $K\in\mathcal{B}(\mathbb{R}^{r},\mathcal{H})$ and $\mathcal{C}>0$ Since both $\mathcal{C}_{\textup{pr}}$ and $KK^{*}$ belong to $L_{1}(\mathcal{H})_{\mathbb{R}}$, it follows that $\mathcal{C}\in L_{1}(\mathcal{H})_{\mathbb{R}}$. Thus, by (8), it remains to show that $\mathcal{C}^{-1/2}\mathcal{C}_{\textup{pr}}^{1/2}$ satisfies property E, as defined in Definition 3.2. As the set of operators which satisfy Property E is closed under taking adjoints, it suffices to show that $(\mathcal{C}^{-1/2}\mathcal{C}_{\textup{pr}}^{1/2})^{*}$ satisfies property E. By Lemma A.20 and Definition 3.2, we must then show $\operatorname{ran}{\mathcal{C}_{\textup{pr}}^{1/2}}=\operatorname{ran}{\mathcal{C}^{1/2}}$, and $(\mathcal{C}^{-1/2}\mathcal{C}_{\textup{pr}}^{1/2})(\mathcal{C}^{-1/2}\mathcal{C}_{\textup{pr}}^{1/2})^{*}-I\in L_{2}(\mathcal{H})$.
By Lemma 4.5(i), $\mathcal{C}^{-1}=\mathcal{C}_{0}^{-1}+UU^{*}$ for some $U\in\mathcal{B}(\mathbb{R}^{r},\mathcal{H})$ with $\operatorname{rank}\left({U}\right)=\operatorname{rank}\left({K}\right)$. Since $\operatorname{dom}{UU^{*}}=\mathcal{H}$,
$\operatorname{ran}{\mathcal{C}}=\operatorname{dom}{\mathcal{C}^{-1}}=\operatorname{dom}{\mathcal{C}_{\textup{pr}}^{-1}+UU^{*}}=\operatorname{dom}{\mathcal{C}_{\textup{pr}}^{-1}}=\operatorname{ran}{\mathcal{C}_{\textup{pr}}}.$
Let $\sum_{i=1}^{\operatorname{rank}\left({K}\right)}c_{i}f_{i}\otimes f_{i}$ be a diagonalisation of the self-adjoint, $\operatorname{rank}\left({K}\right)$-operator $\mathcal{C}_{\textup{pr}}^{1/2}UU^{*}\mathcal{C}_{\textup{pr}}^{1/2}$, with $(f_{i})_{i}$ an ONB of $\mathcal{H}$, $f_{i}\in\operatorname{ran}{\mathcal{C}_{\textup{pr}}^{1/2}}$ for all $i\leq\operatorname{rank}\left({K}\right)$ and $c_{i}^{2}\geq c_{i+1}^{2}\geq 0$. On $\operatorname{ran}{\mathcal{C}}=\operatorname{ran}{\mathcal{C}_{\textup{pr}}}$ it thus holds that
$\mathcal{C}^{-1}=\mathcal{C}_{\textup{pr}}^{-1}+UU^{*}=\mathcal{C}_{\textup{pr}}^{-1/2}(I+\mathcal{C}_{\textup{pr}}^{1/2}UU^{*}\mathcal{C}_{\textup{pr}}^{1/2})\mathcal{C}_{\textup{pr}}^{-1/2}=\mathcal{C}_{\textup{pr}}^{-1/2}(I+\sum_{i=1}^{\operatorname{rank}\left({K}\right)}c_{i}^{2}f_{i}\otimes f_{i})\mathcal{C}_{\textup{pr}}^{-1/2}.$ (38)
By Lemma A.13(i), we have $(\mathcal{C}^{-1/2}\mathcal{C}_{\textup{pr}}^{1/2})^{*}=\mathcal{C}_{\textup{pr}}^{1/2}\mathcal{C}^{-1/2}$ on $\operatorname{dom}{\mathcal{C}^{-1/2}}=\operatorname{ran}{\mathcal{C}^{1/2}}\subset\operatorname{ran}{\mathcal{C}}$. Therefore, on $\operatorname{ran}{\mathcal{C}}=\operatorname{ran}{\mathcal{C}}_{\textup{pr}}$
$(\mathcal{C}^{-1/2}\mathcal{C}_{\textup{pr}}^{1/2})^{*}(\mathcal{C}^{-1/2}\mathcal{C}_{\textup{pr}}^{1/2})-I=\mathcal{C}_{\textup{pr}}^{1/2}\mathcal{C}^{-1}\mathcal{C}_{\textup{pr}}^{1/2}-I=\sum_{i=1}^{\operatorname{rank}\left({K}\right)}c_{i}^{2}f_{i}\otimes f_{i}.$
Since $\operatorname{ran}{\mathcal{C}_{\textup{pr}}}\subset\mathcal{H}$ is dense, this equality holds on all of $\mathcal{H}$, whence $(\mathcal{C}^{-1/2}\mathcal{C}_{\textup{pr}}^{1/2})^{*}(\mathcal{C}^{-1/2}\mathcal{C}_{\textup{pr}}^{1/2})-I\in L_{2}(\mathcal{H})$. We now show that $\operatorname{ran}{\mathcal{C}^{1/2}}=\operatorname{ran}{\mathcal{C}_{\textup{pr}}^{1/2}}$. Recall the definion $\lVert\cdot\rVert_{\mathcal{C}_{\textup{pr}}^{-1}}$ of the Cameron–Martin norm on $\operatorname{ran}{\mathcal{C}_{\textup{pr}}^{1/2}}$ as $\lVert h\rVert_{\mathcal{C}_{\textup{pr}}^{-1}}=\lVert\mathcal{C}_{\textup{pr}}^{-1/2}\rVert$ from Section 2. Similarly, we define the norm $\lVert\cdot\rVert_{\mathcal{C}^{-1}}$ on $\operatorname{ran}{\mathcal{C}^{1/2}}$. The Cameron–Martin space $(\operatorname{ran}{C}_{\textup{pr}}^{1/2},\lVert\cdot\rVert_{\mathcal{C}_{\textup{pr}}^{-1}})$ can equivalently be defined as a completion of $\operatorname{ran}{\mathcal{C}}_{\textup{pr}}$ with respect to $\lVert\cdot\rVert_{\mathcal{C}_{\textup{pr}}^{-1}}$, see [16, Definition 3.49]. The analogous statement holds for $(\operatorname{ran}{C}^{1/2},\lVert\cdot\rVert_{\mathcal{C}^{-1}})$. We can thus show $\operatorname{ran}{\mathcal{C}_{\textup{pr}}^{1/2}}=\operatorname{ran}{\mathcal{C}_{\textup{pos}}^{1/2}}$ holds by showing that the norms $\lVert\cdot\rVert_{\mathcal{C}^{-1}}$ and $\lVert\cdot\rVert_{\mathcal{C}_{\textup{pr}}^{-1}}$ are equivalent on $\operatorname{ran}{\mathcal{C}}=\operatorname{ran}{\mathcal{C}_{\textup{pr}}}$, as it then follows that also the completions of $\operatorname{ran}{\mathcal{C}}$ and $\operatorname{ran}{\mathcal{C}_{\textup{pr}}}$ in either of the equivalent Cameron–Martin norms agree.
Now, on the one hand, $\langle\mathcal{C}^{-1}h,h\rangle=\langle(\mathcal{C}_{\textup{pr}}^{-1}+UU^{*})h,h\rangle\geq\langle\mathcal{C}_{\textup{pr}}^{-1}h,h\rangle$ for $h\in\operatorname{ran}{\mathcal{C}_{\textup{pr}}}$. On the other hand, for any $h\in\operatorname{ran}{C}_{\textup{pr}}$ we have by (38) and the fact that $\mathcal{C}_{\textup{pr}}^{-1/2}$ is self-adjoint
$\langle\mathcal{C}^{-1}h,h\rangle$ $=\langle\mathcal{C}_{\textup{pr}}^{-1/2}(I+\sum_{i=1}^{\operatorname{rank}\left({K}\right)}c_{i}^{2}f_{i}\otimes f_{i})\mathcal{C}_{\textup{pr}}^{-1/2}h,h\rangle$
$=\langle(I+\sum_{i=1}^{\operatorname{rank}\left({K}\right)}c_{i}^{2}f_{i}\otimes f_{i})\mathcal{C}_{\textup{pr}}^{-1/2}h,\mathcal{C}_{\textup{pr}}^{-1/2}h\rangle$
$=\lVert\mathcal{C}_{\textup{pr}}^{-1/2}h\rVert^{2}+\sum_{i=1}^{\operatorname{rank}\left({K}\right)}c_{i}^{2}\langle f_{i},\mathcal{C}_{\textup{pr}}^{-1/2}h\rangle^{2}.$
Because and $(c_{i}^{2})_{i}$ is nonincreasing, this implies
$\langle\mathcal{C}^{-1}h,h\rangle$ $\leq\lVert\mathcal{C}_{\textup{pr}}^{-1/2}h\rVert^{2}+c_{1}^{2}\sum_{i=1}^{\operatorname{rank}\left({K}\right)}\langle f_{i},\mathcal{C}_{\textup{pr}}^{-1/2}h\rangle^{2}$
$\leq\lVert\mathcal{C}_{\textup{pr}}^{-1/2}h\rVert^{2}+c_{1}^{2}\sum_{i=1}^{\infty}\langle f_{i},\mathcal{C}_{\textup{pr}}^{-1/2}h\rangle^{2}$
$\leq(1+c_{1}^{2})\langle\mathcal{C}_{\textup{pr}}^{-1}h,h\rangle.$
Thus $\operatorname{ran}{\mathcal{C}^{1/2}}=\operatorname{ran}{\mathcal{C}_{\textup{pr}}^{1/2}}$, which concludes the proof.
See 4.7
Proof of Lemma 4.7.
To see that the self-adjoint operator $g(U)$ is a Hilbert–Schmidt operator, notice that by Lemma 4.6, $(\mathcal{C}_{\textup{pr}}^{-1}+UU^{*})^{-1}\in\mathcal{E}$ for $U\in\mathcal{B}(\mathbb{R}^{r},\mathcal{H})$. That is, by Lemma 3.3(ii) applied with $\mathcal{C}_{1}\leftarrow(\mathcal{C}_{\textup{pr}}^{-1}+UU^{*})^{-1}$ and $\mathcal{C}_{2}\leftarrow\mathcal{C}_{\textup{pos}}$, $g(U)$ is Hilbert–Schmidt and
$\mathcal{C}_{\textup{pos}}^{1/2}(\mathcal{C}_{\textup{pr}}^{-1}+UU^{*})\mathcal{C}_{\textup{pos}}^{1/2}-I\subset g(U)=\sum_{i}\gamma_{i}e_{i}\otimes e_{i},$
for some ONB $(e_{i})_{i}$ of $\mathcal{H}$ and eigenvalue sequence $(\gamma_{i})_{i}\in\ell^{2}((-1,\infty))$. By Lemma 3.3(i)-(ii) with $\mathcal{C}_{1}\leftarrow\mathcal{C}_{\textup{pr}}$ and $\mathcal{C}_{2}\leftarrow\mathcal{C}_{\textup{pos}}$, and by Lemma 3.6, we have that $\operatorname{rank}\left({(\mathcal{C}_{\textup{pr}}^{-1/2}\mathcal{C}_{\textup{pos}}^{1/2})^{*}\mathcal{C}_{\textup{pr}}^{-1/2}\mathcal{C}_{\textup{pos}}^{1/2}-I}\right)=\operatorname{rank}\left({H}\right)$ and
$\mathcal{C}_{\textup{pos}}^{1/2}(\mathcal{C}_{\textup{pr}}^{-1}+UU^{*})\mathcal{C}_{\textup{pos}}^{1/2}-I$ $=\mathcal{C}_{\textup{pos}}^{1/2}\mathcal{C}_{\textup{pr}}^{-1}\mathcal{C}_{\textup{pos}}^{1/2}-I+\mathcal{C}_{\textup{pos}}^{1/2}UU^{*}\mathcal{C}_{\textup{pos}}^{1/2}$
$\subset(\mathcal{C}_{\textup{pr}}^{-1/2}\mathcal{C}_{\textup{pos}}^{1/2})^{*}\mathcal{C}_{\textup{pr}}^{-1/2}\mathcal{C}_{\textup{pos}}^{1/2}-I+\mathcal{C}_{\textup{pos}}^{1/2}UU^{*}\mathcal{C}_{\textup{pos}}^{1/2}.$
As $\mathcal{C}_{\textup{pos}}^{1/2}(\mathcal{C}_{\textup{pr}}^{-1}+UU^{*})\mathcal{C}_{\textup{pos}}^{1/2}-I$ is continuous and densely defined, the above extension is unique, which shows that (18) holds. Now, by injectivity of $\mathcal{C}_{\textup{pos}}^{1/2}$, we conclude from (18) that
$\operatorname{rank}\left({g(U)}\right)\leq\operatorname{rank}\left({(\mathcal{C}_{\textup{pr}}^{-1/2}\mathcal{C}_{\textup{pos}}^{1/2})^{*}\mathcal{C}_{\textup{pr}}^{-1/2}\mathcal{C}_{\textup{pos}}^{1/2}-I)}\right)+\operatorname{rank}\left({\mathcal{C}_{\textup{pos}}UU^{*}\mathcal{C}_{\textup{pos}}^{1/2}}\right)=\operatorname{rank}\left({H}\right)+r.$
and that $\operatorname{ran}{g(U)}\subset\operatorname{ran}{\mathcal{C}_{\textup{pos}}^{1/2}}=\operatorname{ran}{\mathcal{C}_{\textup{pr}}^{1/2}}$. For $i\leq\operatorname{rank}\left({H}\right)+r$ we have $\gamma_{i}\not=0$ and it follows that $e_{i}\in\operatorname{ran}{\mathcal{C}_{\textup{pr}}^{1/2}}$ for $i\leq\operatorname{rank}\left({H}\right)+r$. By Lemma A.2, we can extend $(e_{i})_{i=1}^{\operatorname{rank}\left({H}\right)+r}$ to a sequence in $\operatorname{ran}{\mathcal{C}_{\textup{pr}}^{1/2}}$ which is an ONB of $\mathcal{H}$. By replacing $(e_{i})_{i}$ with this sequence, we still have $g(U)=\sum_{i}\gamma_{i}e_{i}\otimes e_{i}$ and now $e_{i}\subset\operatorname{ran}{\mathcal{C}_{\textup{pr}}^{1/2}}$ for all $i$.
In (17), the order of $(\mathcal{C}_{\textup{pr}}^{-1}+UU^{*})^{1/2}\mathcal{C}_{\textup{pos}}^{1/2}$ and $\bigr{(}(\mathcal{C}_{\textup{pr}}^{-1}+UU^{*})^{1/2}\mathcal{C}_{\textup{pos}}^{1/2}\bigr{)}^{*}$ is reversed. Therefore, by comparing Lemma 3.3 (i) and (ii) with the same choices stated above, it follows that the eigenvalues of $R(\mathcal{C}_{\textup{pos}}\|(\mathcal{C}_{\textup{pr}}^{-1}+UU^{*})^{-1})$ and $g(U)$ agree, counting multiplicities. ∎
See 4.9
Proof of Lemma 4.9.
Let $U,V\in\mathcal{B}(\mathbb{R}^{r},\mathcal{H})$, $V\not=0$. We have by (18),
$g(U+V)-g(U)=\mathcal{C}_{\textup{pos}}^{1/2}((U+V)(U+V)^{*}-UU^{*})\mathcal{C}_{\textup{pos}}^{1/2}=\mathcal{C}_{\textup{pos}}^{1/2}(UU^{*}+VV^{*}+UV^{*}+VU^{*}).$
From $\lVert VV^{*}\rVert=\lVert V\rVert^{2}$, c.f. Lemma A.3, it follows that
$\lVert g(U+V)-g(U)-\mathcal{C}_{\textup{pos}}^{1/2}(UV^{*}+VU^{*})\mathcal{C}_{\textup{pos}}^{1/2}\rVert=\lVert\mathcal{C}_{\textup{pos}}^{1/2}(VV^{*})\mathcal{C}_{\textup{pos}}^{1/2}\rVert\leq\lVert\mathcal{C}_{\textup{pos}}^{1/2}\rVert^{2}\lVert VV^{*}\rVert=\lVert\mathcal{C}_{\textup{pos}}^{1/2}\rVert^{2}\lVert V\rVert^{2}$
Dividing by $\lVert V\rVert$ and letting $\lVert V\rVert\rightarrow 0$, this shows that $g$ is differentiable with the correct derivative.
To show differentiability of $F_{f}$, let $x=(x_{i})_{i}\in\ell^{2}((-1,\infty))$, $y=(y_{i})_{i}\in\ell^{2}(\mathbb{R})$ and define $c_{x}\in\ell^{2}(\mathbb{R})$ by $(c_{x})_{i}=f^{\prime}(x_{i})$. By the assumption $f\in\mathscr{F}$, $f^{\prime}$ is Lipschitz continuous in some neighbourhood $(-a,a)$ of 00, with $a>0$ and with Lipschitz constant $M_{0}$. Let us take $N_{x}$ so large that $\lvert x_{i}\rvert<a/2$ for $i>N_{x}$. Let $\varepsilon>0$ arbitrary. By differentiability of $f$, we can choose $\delta_{x,\varepsilon}>0$ such that $x_{i}+z\in(-1,\infty)$ and $\lvert(f(x_{i}+z)-f(x_{i})-f^{\prime}(x_{i})z)\rvert<\varepsilon\lvert z\rvert$ for $\lvert z\rvert<\delta_{x,\varepsilon}$ and $i=1,\ldots,N_{x}$. We then have for $\lVert y\rVert<\min(\delta_{x,\varepsilon},a/2,\varepsilon)$,
$\frac{1}{\lVert y\rVert}\lvert F_{f}(x+y)-F_{f}(x)-\langle c_{x},y\rangle\rvert$ $=\frac{1}{\lVert y\rVert}\lvert\sum_{i}f(x_{i}+y_{i})-f(x_{i})-f^{\prime}(x_{i})y_{i}\rvert$
$\leq\sum_{i=1}^{N_{x}}\frac{1}{\lvert y_{i}\rvert}\lvert f(x_{i}+y_{i})-f(x_{i})-f^{\prime}(x_{i})y_{i}\rvert$
$+\frac{1}{\lVert y\rVert}\sum_{i>N_{x}}\lvert f(x_{i}+y_{i})-f(x_{i})-f^{\prime}(x_{i})y_{i}\rvert.$
As $\lvert y_{i}\rvert<\delta_{x,\varepsilon}$ for $i=1,\ldots,N_{x}$, the first term is bounded from above by $N_{x}\varepsilon$. For the second term, by the mean value theorem, for each $i>N_{x}$ we can find $c_{i}\in[x_{i}-\lvert y_{i}\rvert,x_{i}+\lvert y_{i}\rvert]\subset(-a,a)$ such that,
$\frac{1}{\lVert y\rVert}\sum_{i>N_{x}}\lvert f(x_{i}+y_{i})-f(x_{i})-f^{\prime}(x_{i})y_{i}\rvert$ $=\frac{1}{\lVert y\rVert}\sum_{i>N_{x}}\lvert f^{\prime}(c_{i})y_{i}-f^{\prime}(x_{i})y_{i}\rvert$
$\leq\frac{1}{\lVert y\rVert}\sum_{i>N_{x}}M_{0}\lvert c_{i}-x_{i}\rvert\lvert y_{i}\rvert$
$\leq\frac{1}{\lVert y\rVert}\sum_{i>N_{x}}M_{0}\lvert y_{i}\rvert^{2}\leq M_{0}\lVert y\rVert\leq M_{0}\varepsilon,$
where we used the Lipschitz continuity of $f^{\prime}$ in $(-a,a)$ in the first inequality, and the fact that $c_{i}\in[x_{i}-\lvert y_{i}\rvert,x_{i}+\lvert y_{i}\rvert]$ in the second inequality. Therefore,
$\frac{1}{\lVert y\rVert}\lvert F_{f}(x+y)-F_{f}(x)-\langle c_{x},y\rangle\rvert\leq(N_{x}+M_{0})\varepsilon,$
from which we conclude that $F_{f}^{\prime}(x)$ exists and $F_{f}^{\prime}(x)=c_{x}$. ∎
See 4.12
Proof of Proposition 4.12.
Any statement regarding differentiability in this proof should be understood as Fréchet differentiability. Let us write Sym($m$) for the symmetric matrices on $\mathbb{R}^{m}$ endowed with the trace-inner product: $\langle A,B\rangle_{\text{Sym(}m\text{)}}=\operatorname{tr}\,\left({BA}\right)$ for symmetric matrices $A$ and $B$ on $\mathbb{R}^{m}$. The result of [19, Theorem 1.1] is a statment on Sym($m$), and we therefore identify Sym($m$) with $L_{2}(\mathcal{Z})_{\mathbb{R}}$.
Let $(\varphi_{i})_{i}$ the standard basis of $\mathbb{R}^{m}$. Define $\Phi:L_{2}(\mathcal{Z})_{\mathbb{R}}\rightarrow\text{Sym(}m\text{)}$ by $\Phi(X)=\sum_{i,j=1}^{m}\langle Xe_{j},e_{i}\rangle\varphi_{i}\otimes\varphi_{j}$ and let $X\in L_{2}(\mathcal{Z})_{\mathbb{R}}$. Then $\Phi$ is an isomorphism of Hilbert spaces and by linearity of $\Phi$ we have $\Phi^{\prime}(X)(X_{2})=\sum_{i,j=1}^{m}\langle X_{2}e_{j},e_{i}\rangle\varphi_{i}\otimes\varphi_{j}$ for all $X_{2}\in L_{2}(\mathcal{Z})_{\mathbb{R}}$. Furthermore, $\Lambda^{m}\circ\Phi^{-1}$ is the eigenvalue map on Sym($m$), where the eigenvalues are ordered in a nonincreasing way. We note that $\Phi(X)=\sum_{i=1}^{m}(\Lambda^{m}\circ\Phi^{-1})_{i}(\Phi(X))\varphi_{i}\otimes\varphi_{i}$. By the chain rule, $G\circ\Lambda^{m}$ is differentiable in $X$ if and only $G\circ\Lambda^{m}\circ\Phi^{-1}$ is differentiable in $\Phi(X)$. By [19, Theorem 1.1], $G\circ\Lambda^{m}\circ\Phi^{-1}$ is differentiable in $\Phi(X)$ if and only if $G$ is differentiable in $\Lambda^{m}\circ\Phi(\Phi(X))=\Lambda^{m}(X)$, and the derivative is given by
$(G\circ\Lambda^{m}\circ\Phi^{-1})^{\prime}(\Phi(X))=\sum_{i=1}^{m}G^{\prime}\left(\Lambda^{m}\circ\Phi^{-1}(\Phi(X))\right)_{i}\varphi_{i}\otimes\varphi_{i}=\sum_{i=1}^{m}G^{\prime}(\Lambda^{m}(X))_{i}\varphi_{i}\otimes\varphi_{i}.$
Thus, $G\circ\Lambda^{m}$ is differentiable in $X$ if and only if $G$ is differentiable in $\Lambda^{m}(X)$. Another application of the chain rule, the expression for $\Phi^{\prime}$ and the previous equation then finish the proof by showing that
$\langle(G\circ\Lambda^{m})^{\prime}(X),X_{2}\rangle_{L_{2}(\mathcal{X})}$ $=\langle(G\circ\Lambda^{m}\circ\Phi^{-1}\circ\Phi)^{\prime}(X),X_{2}\rangle_{L_{2}(\mathcal{Z})}$
$=\Big{\langle}(G\circ\Lambda^{m}\circ\Phi^{-1})^{\prime}(\Phi(X)),\Phi^{\prime}(X)(X_{2})\Big{\rangle}_{\text{Sym(}m\text{)}}$
$=\Big{\langle}\sum_{i=1}^{m}G^{\prime}(\Lambda^{m}(X))_{i}\varphi_{i}\otimes\varphi_{i},\sum_{k,j=1}^{m}\langle X_{2}e_{j},e_{k}\rangle_{L_{2}(\mathcal{H})}\varphi_{k}\otimes\varphi_{j}\Big{\rangle}_{\text{Sym(}m\text{)}}$
$=\sum_{i=1}^{m}G^{\prime}(\Lambda^{m}(X))_{i}\langle X_{2}e_{i},e_{i}\rangle_{L_{2}(\mathcal{H})}$
$=\langle\sum_{i=1}^{m}G^{\prime}(\Lambda^{m}(X))e_{i}\otimes e_{i},X_{2}\rangle_{L_{2}(\mathcal{Z})}.$
See 4.14
Proof of Proposition 4.14.
Any statement regarding differentiability in this proof should be understood as Fréchet differentiability. For $X\in\mathcal{W}$, we have $\ker{X}^{\perp}=\ker{X^{*}}^{\perp}=\overline{\operatorname{ran}{X}}=\operatorname{ran}{X}$ for $X\in\mathcal{W}$ by Lemma A.7. Let $m\coloneqq\dim{\mathcal{Z}}$ and extend $(e_{i})_{i}$ to an ONB of $\mathcal{Z}$. Note that the $m^{2}$ operators $e_{i}\otimes e_{j}$, $i,j\leq m$, span the space $\mathcal{W}$. Therefore, $\dim{\mathcal{W}}=m^{2}<\infty$. Because finite-dimensional spaces are closed, we can define $P_{\mathcal{Z}}:\mathcal{H}\rightarrow\mathcal{Z}$, the orthogonal projector onto $\mathcal{Z}$. Furthermore, we let $P_{m}:\ell^{2}(\mathbb{R})\rightarrow\mathbb{R}^{m}$ be the orthogonal projector onto the first $m$-coordinates of an $\ell^{2}$ sequence. Thus, $P_{m}^{*}$ is the natural embedding of $\mathbb{R}^{m}$ into $\ell^{2}(\mathbb{R})$. Since Proposition 4.12 is a statement on $L_{2}$ spaces of finite dimension, we identify $\mathcal{W}$ with $L_{2}(\mathcal{Z})_{\mathbb{R}}$ via $\Phi:L_{2}(\mathcal{H})_{\mathbb{R}}\rightarrow L_{2}(\mathcal{Z})_{\mathbb{R}}$ by $\Phi(X)=P_{\mathcal{Z}}{\left.\kern-1.2ptX\vphantom{\big{|}}\right|_{Z}}$. Now, $\Phi$ is linear and surjective and $\lVert\Phi\rVert\leq 1$. For $X\in\mathcal{W}$, $\operatorname{ran}{X}\subset\mathcal{Z}$ and $\ker{X}^{\perp}=\operatorname{ran}{X}$, hence $\lVert\Phi(X)\rVert=\lVert{\left.\kern-1.2ptX\vphantom{\big{|}}\right|_{\mathcal{Z}}}\rVert=\lVert X\rVert$ and $\Phi$ is an isometry on $\mathcal{W}$.
We first prove the result for the case in which $\Lambda$ assigns a nonincreasing order to eigenvalues of nonnegative operators in $L_{2}(\mathcal{Z})_{\mathbb{R}}$. With $\Lambda^{m}$ defined before Proposition 4.12, we then have $\Lambda(X)_{i}=\Lambda^{m}(\Phi(X))_{i}$ for all $i\leq m$ and $X\in\mathcal{W}$. This implies that
$P_{m}\Lambda(X)$ $=\Lambda^{m}(\Phi(X)),\quad X\in\mathcal{W},$
$\Lambda(X)$ $=P_{m}^{*}\Lambda^{m}(\Phi(X)),\quad X\in\mathcal{W},$ (39)
since any $X\in\mathcal{W}$ has at most $m$ nonzero eigenvalues. With $F^{m}_{f}\coloneqq FP_{m}^{*}$, we also have
$(F\circ\Lambda)(X)=(F^{m}_{f}\circ\Lambda^{m})(\Phi(X)),\quad X\in\mathcal{W}.$ (40)
By the hypothesis on $F$, $F$ is differentiable in $\Lambda(X)$ for $X\in\mathcal{W}$. Furthermore, $P_{m}^{*}$ is linear, hence differentiable, and $(P_{m}^{*})^{\prime}(x)(y)=P_{m}^{*}y$ for $x,y\in\mathbb{R}^{m}$. Then, by (39) and the chain rule, the composition $F^{m}_{f}=F\circ P_{m}^{*}$ is differentiable in $\Lambda^{m}(\Phi(X))$, and by (39) it holds for any $y\in\mathbb{R}^{m}$,
$\langle(F^{m}_{f})^{\prime}(\Lambda^{m}(\Phi(X))),y\rangle=\langle F^{\prime}(P_{m}^{*}\Lambda^{m}(\Phi(X))),P_{m}^{*}y\rangle=\langle P_{m}F^{\prime}(\Lambda(X)),y\rangle,$
i.e.
$(F^{m})^{\prime}(\Lambda^{m}(\Phi(X)))=P_{m}F^{\prime}(\Lambda(X))\in\mathbb{R}^{m}.$ (41)
We may now apply Proposition 4.12 to conclude that $F^{m}\circ\Lambda^{m}$ is differentiable in $\Phi(X)$. By the fact that $e_{i}\in\operatorname{ran}{Z}$ and by the stated diagonalisation of $X$, we have $\Phi(X)=\sum_{i=1}^{m}\Lambda^{m}(\Phi(X))_{i}e_{i}\otimes e_{i}$, where the rank-1 operators $e_{i}\otimes e_{i}$ are now understood to act only on $\mathcal{Z}$. With Proposition 4.12 we thus also obtain the expression for the derivative
$(F^{m}\circ\Lambda^{m})^{\prime}(\Phi(X))=\sum_{i=1}^{m}(F^{m})^{\prime}(\Lambda^{m}(\Phi(X)))_{i}e_{i}\otimes e_{i}=\sum_{i=1}^{m}F^{\prime}(\Lambda(X))_{i}e_{i}\otimes e_{i}\in L_{2}(\mathcal{Z})_{\mathbb{R}},$ (42)
where for second equation we use (41), the definition of $P_{m}$ and the fact that the sums extend only upto $m$. Another application of the chain rule to (40) now shows that $F\circ\Lambda$ is differentiable in $X$ and, because $\Phi^{\prime}(X)(Y)=\Phi(Y)$ for $Y\in\mathcal{W}$, that
$\langle(F\circ\Lambda)^{\prime}(X),Y\rangle_{L_{2}(\mathcal{H})}$ $=\langle(F^{m}\circ\Lambda^{m})^{\prime}(\Phi(X)),\Phi(Y)\rangle_{L_{2}(\mathcal{Z})}=\Big{\langle}\sum_{i=1}^{m}F^{\prime}(\Lambda(X))_{i}e_{i}\otimes e_{i},\Phi(Y)\Big{\rangle}_{L_{2}(\mathcal{Z})}$
$=\Big{\langle}\sum_{i=1}^{m}F^{\prime}(\Lambda(X))_{i}e_{i}\otimes e_{i},Y\Big{\rangle}_{L_{2}(\mathcal{\mathcal{H}})}=\Big{\langle}\sum_{i=1}^{\infty}F^{\prime}(\Lambda(X))_{i}e_{i}\otimes e_{i},Y\Big{\rangle}_{L_{2}(\mathcal{\mathcal{H}})},$
where in the third step we use $Y\in\mathcal{W}$ and again interpret $e_{i}\otimes e_{i}$ as acting on all of $\mathcal{H}$, and in the last step we use that the linear span of $(e_{i})_{i=1}^{m}$ contains $\operatorname{ran}{Y}\subset\mathcal{Z}$. Since $Y\in\mathcal{W}$ is arbitrary, this shows that ${\left.\kern-1.2pt(F\circ\Lambda)\vphantom{\big{|}}\right|_{\mathcal{W}}}$ is differentiable in $X$.
Finally, let us denote by $\tilde{\Lambda}$ an eigenvalue map on $L_{2}(\mathcal{H})_{\mathbb{R}}$ which can assign any fixed but arbitrary ordering on the eigenvalues. Let $X\in\mathcal{W}$, $X=\sum_{i=1}\tilde{\Lambda}(X)_{i}e_{i}\otimes e_{i}$ be given and assume that $F$ is differentiable in $\tilde{\Lambda}(X)$. Given $X$, there exists a permutation $\alpha:\mathbb{N}\rightarrow\mathbb{N}$ such that $\tilde{\Lambda}(X)_{i}=\Lambda(X)_{\alpha(i)}$ for all $i\in\mathbb{N}$. Let $P_{\alpha}:\ell^{2}(\mathbb{R})\rightarrow\ell^{2}(\mathbb{R})$ denote the permutation operator $(P_{\alpha}x)_{i}=(Px)_{\alpha(i)}$, $i\in\mathbb{N}$, so $P_{\alpha}^{*}=P_{\alpha}^{-1}$. Then $\tilde{\Lambda}(X)=(P_{\alpha}\Lambda)(X)$ and $X=\sum_{i}\Lambda_{\alpha(i)}(X)e_{i}\otimes e_{i}=\sum_{i}\Lambda(X)_{i}e_{\alpha^{-1}(i)}\otimes e_{\alpha^{-1}(i)}.$ By the previous part of the proof, ${\left.\kern-1.2pt(F\circ\Lambda)\vphantom{\big{|}}\right|_{\mathcal{W}}}$ is differentiable in $X$. Because $F\circ\tilde{\Lambda}=F\circ\Lambda$ by symmetry of $F$, differentiability of ${\left.\kern-1.2pt(F\circ\tilde{\Lambda})\vphantom{\big{|}}\right|_{\mathcal{W}}}$ in $X$ follows, and
$(F\circ\tilde{\Lambda})^{\prime}(X)=(F\circ\Lambda)^{\prime}(X)=\sum_{i}F_{i}^{\prime}(\Lambda(X))e_{\alpha^{-1}(i)}\otimes e_{\alpha^{-1}(i)}.$ (43)
Since $F$ is symmetric, $F\circ P_{\alpha}=F$. Hence, for $x\in\ell^{2}(\mathbb{R})$,
$F^{\prime}(x)=(F\circ P_{\alpha})^{\prime}(x)=P_{\alpha}^{*}F^{\prime}(P_{\alpha}x)=P_{\alpha}^{-1}F^{\prime}(P_{\alpha}x).$
Thus, $F_{i}^{\prime}(\Lambda(X))=F_{\alpha^{-1}(i)}^{\prime}(P_{\alpha}\Lambda(X))=F_{\alpha^{-1}(i)}^{\prime}(\tilde{\Lambda}(X))$. From (43) it now follows that
$(F\circ\tilde{\Lambda})^{\prime}(X)=\sum_{i}F_{\alpha^{-1}(i)}^{\prime}(\tilde{\Lambda}(X))e_{\alpha^{-1}(i)}\otimes e_{\alpha^{-1}(i)}=\sum_{i}F_{i}^{\prime}(\tilde{\Lambda}(X))e_{i}\otimes e_{i}.$
See 4.15
Proof of Proposition 4.15.
Let $U,V\in\mathcal{B}(\mathbb{R}^{r},\mathcal{H})$. Define $\mathcal{Z}\coloneqq\operatorname{ran}{g(U+V)}$ and $\mathcal{W}\coloneqq\{X\in L_{2}(\mathcal{H})_{\mathbb{R}}:\ \operatorname{ran}{X}\subset\mathcal{Z}\}\subset L_{2}(\mathcal{H})_{\mathbb{R}}$. Then $\dim{\mathcal{Z}}<\infty$ and $\operatorname{ran}{g(U+tV)}\subset\mathcal{Z}$ for all $t\in\mathbb{R}$, hence $g(U+tV)\in\mathcal{W}$ for all $t\in\mathbb{R}$. Thus, $F_{f}\circ\Lambda\circ g(U+tV)={\left.\kern-1.2pt(F\circ\Lambda)\vphantom{\big{|}}\right|_{\mathcal{W}}}\circ g(U+tV)$ for all $t\in\mathbb{R}$. By Lemma 4.9 and Proposition 4.14, ${\left.\kern-1.2pt(F\circ\Lambda)\vphantom{\big{|}}\right|_{\mathcal{W}}}$ is Fréchet differentiable. By Lemma 4.9, $g$ is Fréchet differentiable. In particular, $g$ is Gateaux differentiable in $U$ in direction $V$. Hence, by the chain rule, $J_{f}$ is Gateaux differentiable in $U$ in direction $V$. To compute the derivative, we recall that $(e_{j}\otimes e_{k})_{j,k}$ is an ONB of $L_{2}(\mathcal{H})$. By the expression of the derivative in Proposition 4.14, the Gateaux derivative of $J_{f}$ in $U$ in direction $V$ is
$J_{f}^{\prime}(U)(V)$ $=({\left.\kern-1.2pt(F_{f}\circ\Lambda)\vphantom{\big{|}}\right|_{\mathcal{W}}}\circ g)^{\prime}(U)(V)$
$=\langle{\left.\kern-1.2pt(F_{f}\circ\Lambda)\vphantom{\big{|}}\right|_{\mathcal{W}}}^{\prime}(g(U)),g^{\prime}(U)V\rangle_{L_{2}(\mathcal{H})}$
$=\langle\sum_{i}f^{\prime}\left(\Lambda_{i}(g(U))\right)e_{i}\otimes e_{i},g^{\prime}(U)V\rangle_{L_{2}(\mathcal{H})}$
$=\sum_{j,k}\langle\sum_{i}f^{\prime}\left(\Lambda_{i}(g(U))\right)e_{i}\otimes e_{i},e_{j}\otimes e_{k}\rangle_{L_{2}(\mathcal{H})}\langle g^{\prime}(U)V,e_{j}\otimes e_{k}\rangle_{L_{2}(\mathcal{H})}$
$=\sum_{i}f^{\prime}\left(\Lambda_{i}(g(U))\right)\langle g^{\prime}(U)V,e_{i}\otimes e_{i}\rangle_{L_{2}(\mathcal{H})}$
$=\sum_{i}f^{\prime}\left(\Lambda_{i}(g(U))\right)\langle\mathcal{C}_{\textup{pos}}^{1/2}(UV^{*}+VU^{*})\mathcal{C}_{\textup{pos}}^{1/2},e_{i}\otimes e_{i}\rangle_{L_{2}(\mathcal{H})}$
$=\sum_{i}f^{\prime}\left(\Lambda_{i}(g(U))\right)\langle\mathcal{C}_{\textup{pos}}^{1/2}(UV^{*}+VU^{*})\mathcal{C}_{\textup{pos}}^{1/2}e_{i},e_{i}\rangle$
$=2\sum_{i}f^{\prime}\left(\Lambda_{i}(g(U))\right)\langle\mathcal{C}_{\textup{pos}}^{1/2}e_{i},VU^{*}\mathcal{C}_{\textup{pos}}^{1/2}e_{i}\rangle.$
See 4.16
Proof of Lemma 4.16.
Let $f\in\mathscr{F}$ and let $(U_{n})_{n}$ be a sequence in $\mathcal{B}(\mathbb{R}^{r},\mathcal{V})$ such that $\lVert U_{n}\rVert\rightarrow\infty$. Then, by Lemma A.3, $\lVert U_{n}U_{n}^{*}\rVert=\lVert U_{n}\rVert^{2}\rightarrow\infty$. Since $f(x)\rightarrow\infty$ for $\lVert x\rVert\rightarrow\infty$ and $f$ is bounded from below, it is enough to show that there is an eigenvalue $\alpha_{n}$ of $g(U_{n})$ with $\lvert\alpha_{n}\rvert\rightarrow\infty$. Since $g(U_{n})$ is self-adjoint and compact, $\lVert g(U_{n})\rVert=\max_{i}\lvert\Lambda_{i}(g(U_{n}))\rvert$ by Lemma A.4, and we must therefore show that $\lVert g(U_{n})\rVert\rightarrow\infty$. For this, it is enough to find a bounded sequence $h_{n}\in\mathcal{H}$, such that $\lVert g(U_{n})h_{n}\rVert\rightarrow\infty$.
For any $h\in\mathcal{H}$, we have by the triangle inequality, (18) and Lemma 3.6
$\lVert g(U)h\rVert$ $=\lVert\mathcal{C}_{\textup{pos}}^{1/2}U_{n}U_{n}^{*}\mathcal{C}_{\textup{pos}}^{1/2}h-(I-(\mathcal{C}_{\textup{pr}}^{-1/2}\mathcal{C}_{\textup{pos}}^{1/2})^{*}\mathcal{C}_{\textup{pr}}^{-1/2}\mathcal{C}_{\textup{pos}}^{1/2})h\rVert$
$\geq\lVert\mathcal{C}_{\textup{pos}}^{1/2}U_{n}U_{n}^{*}\mathcal{C}_{\textup{pos}}^{1/2}h\rVert-\lVert(I-(\mathcal{C}_{\textup{pr}}^{-1/2}\mathcal{C}_{\textup{pos}}^{1/2})^{*}\mathcal{C}_{\textup{pr}}^{-1/2}\mathcal{C}_{\textup{pos}}^{1/2})h\rVert$
$\geq\lVert\mathcal{C}_{\textup{pos}}^{1/2}U_{n}U_{n}^{*}\mathcal{C}_{\textup{pos}}^{1/2}h\rVert-1.$ (44)
Let us write $m\coloneqq\dim{\mathcal{V}}$. For each $n$, let us diagonalise $U_{n}U_{n}^{*}=\sum_{j=1}^{m}\beta_{n,j}\psi_{n,j}\otimes\psi_{n,j}$, where $(\psi_{n,j})_{j=1}^{m}$ forms an ONB of $\mathcal{V}$ and $\beta_{n,j}\geq 0$. Define $j_{n}\coloneqq\operatorname*{arg\,max}_{j\leq m}\beta_{n,j}$, hence $\beta_{n,j_{n}}$ is the largest eigenvalue of $U_{n}U_{n}^{*}$. As $U_{n}U_{n}^{*}$ is self-adjoint, $\beta_{n,j_{n}}=\lVert U_{n}U_{n}^{*}\rVert$ by Lemma A.4, showing $\beta_{n,j_{n}}\rightarrow\infty$. Let $\varepsilon>0$. By density of $\operatorname{ran}{\mathcal{C}^{1/2}_{\textup{pos}}}$, for each $j\leq m$ we can choose $k_{j}\in\operatorname{ran}{\mathcal{C}^{1/2}_{\textup{pos}}}$ which satisfies $\lVert\psi_{1,j}-\mathcal{C}_{\textup{pos}}^{1/2}k_{j}\rVert\leq\varepsilon$. Let us decompose $\psi_{n,j_{n}}=\sum_{j=1}^{m}\langle\psi_{n,j_{n}},\psi_{1,j}\rangle\psi_{1,j}$ and define $h_{n}\coloneqq\sum_{j=1}^{m}\langle\psi_{n,j_{n}},\psi_{1,j}\rangle k_{j}$. Note that $\lVert h_{n}\rVert\leq C$ with $C\coloneqq m\max_{j}\lVert k_{j}\rVert$ by the Cauchy–Schwarz inequality. By further application of the Cauchy–Schwarz inequality,
$\lVert\psi_{n,j_{n}}-\mathcal{C}_{\textup{pos}}^{1/2}h_{n}\rVert^{2}$ $=\sum_{j}\langle\psi_{n,j_{n}},\psi_{1,j}\rangle^{2}\lVert\psi_{1,j}-\mathcal{C}_{\textup{pos}}^{1/2}k_{j}\rVert^{2}$
$+2\sum_{i\not=j}\langle\psi_{n,j_{n}},\psi_{1,i}\rangle\langle\psi_{n,j_{n}},\psi_{1,j}\rangle\langle\psi_{1,i}-\mathcal{C}_{\textup{pos}}^{1/2}k_{i},\psi_{1,j}-\mathcal{C}_{\textup{pos}}^{1/2}k_{j}\rangle$
$\leq m\varepsilon^{2}+2m(m-1)\varepsilon^{2}=m(2m-1)\varepsilon^{2}.$
Thus, by the definition of $(k_{j})_{j}$, the triangle inequality and the bound $(a+b)^{2}\leq 2a^{2}+2b^{2}$,
$\lVert\mathcal{C}_{\textup{pos}}^{1/2}k_{n}\rVert^{2}\geq\frac{1}{2}\lVert\psi_{n,j_{n}}\rVert^{2}-\frac{1}{2}\lVert\psi_{n,j_{n}}-\mathcal{C}_{\textup{pos}}^{1/2}h_{n}\rVert^{2}\geq\frac{1}{2}-m(m-\frac{1}{2})\varepsilon^{2}.$
It follows that for $\varepsilon$ small enough, there exists $c>0$ such that
$\langle\psi_{n,j_{n}},\mathcal{C}_{\textup{pos}}^{1/2}h_{n}\rangle=\frac{1}{2}\left(\lVert\psi_{n,j_{n}}\rVert^{2}+\lVert\mathcal{C}_{\textup{pos}}^{1/2}h_{n}\rVert^{2}-\lVert\psi_{n,j_{n}}-\mathcal{C}_{\textup{pos}}^{1/2}h_{n}\rVert^{2}\right)>c.$
By Cauchy–Schwarz, the bound $\lVert h_{n}\rVert\leq C$, the fact that $\mathcal{C}_{\textup{pr}}^{1/2}$ is self-adjoint, the given diagonalisation of $U_{n}U_{n}^{*}$ and the previous lower bound,
$\lVert\mathcal{C}_{\textup{pos}}^{1/2}U_{n}U_{n}^{*}\mathcal{C}_{\textup{pos}}^{1/2}h_{n}\rVert$ $\geq C^{-1}{\langle\mathcal{C}_{\textup{pos}}^{1/2}U_{n}U_{n}^{*}\mathcal{C}_{\textup{pos}}^{1/2}h_{n},h_{n}\rangle}=C^{-1}\langle U_{n}U_{n}^{*}\mathcal{C}_{\textup{pos}}^{1/2}h_{n},\mathcal{C}_{\textup{pos}}^{1/2}h_{n}\rangle$
$=\sum_{j}\beta_{n,j}\lvert\langle\psi_{n,j},\mathcal{C}_{\textup{pos}}^{1/2}h_{n}\rangle\rvert^{2}\geq C^{-1}\beta_{n,j_{n}}\lvert\langle\psi_{n,j_{n}},\mathcal{C}_{\textup{pos}}^{1/2}h_{n}\rangle\rvert^{2}\geq c^{2}C^{-1}\beta_{n,j_{n}}\rightarrow\infty.$
Combining this with (44), we have thus found a bounded sequence $(h_{n})_{n}$ with $\lVert g(U_{n})h_{n}\rVert\rightarrow\infty$. Finally, by Proposition 4.15, $J_{f}$ is differentiable. A coercive, differentiable function has a global minimum, that is attained only among its stationary points. ∎
See 4.18
Proof of Theorem 4.18.
Let $f\in\mathscr{F}$. By Proposition 4.15, $J_{f}$ is Gateaux differentiable. It follows from [5, Theorem 12.4.5 (i)] that local minimisers of $J_{f}$ are stationary points, i.e. have Gateaux derivative equal to 0. The idea of the proof is to find among all stationary points of $J_{f}$ a stationary point that minimises $J_{f}$, and use the coercivity of $J_{f}$ over finite dimensional subspaces of $\mathcal{B}(\mathbb{R}^{r},\mathcal{H})$ to conclude that this stationary point is a global minimiser.
With $(\gamma_{i})_{i}$ and $(e_{i})_{i}$ as in Lemma 4.7, the Gateaux derivative $J_{F}^{\prime}(g(U))\in L_{2}^{*}\simeq L_{2}$ at $U\in\mathcal{B}(\mathbb{R}^{r},\mathcal{H})$ is given by, in direction $V\in\mathcal{B}(\mathbb{R}^{r},\mathcal{H})$,
$J_{f}^{\prime}(U)(V)=2\sum_{i}f^{\prime}(\gamma_{i})\langle\mathcal{C}_{\textup{pos}}^{1/2}e_{i},VU^{*}\mathcal{C}_{\textup{pos}}^{1/2}e_{i}\rangle.$
Setting the derivative to zero, we obtain for all $V\in\mathcal{B}(\mathbb{R}^{r},\mathcal{H})$,
$\sum_{i}f^{\prime}(\gamma_{i})\langle\mathcal{C}_{\textup{pos}}^{1/2}e_{i},VU^{*}\mathcal{C}_{\textup{pos}}^{1/2}e_{i}\rangle=0.$
For an arbitrary fixed $i$, if $f^{\prime}(\gamma_{i})\not=0$, then this implies that $U^{*}\mathcal{C}_{\textup{pos}}^{1/2}e_{i}=0$, as otherwise there exists $\varphi\in\mathbb{R}^{r}$ such that $\langle\varphi,U^{*}\mathcal{C}_{\textup{pos}}^{1/2}e_{i}\rangle_{\mathbb{R}^{r}}\not=0$ and the choice $V=\mathcal{C}_{\textup{pos}}^{1/2}e_{i}\otimes\varphi$ furnishes a contradiction. Since $f\in\mathscr{F}$, it follows from Lemma 4.1(i) that $f^{\prime}(\gamma_{i})=0$ if and only if $\gamma_{i}=0$. Hence, $\gamma_{i}U^{*}\mathcal{C}_{\textup{pos}}^{1/2}e_{i}=0$ for all $i$.
By (17), (3c) and Lemma A.13(i),
$\mathcal{C}_{\textup{pos}}^{1/2}(UU^{*}-H)\mathcal{C}_{\textup{pos}}^{1/2}=\mathcal{C}_{\textup{pos}}^{1/2}(\mathcal{C}_{\textup{pr}}^{-1}+UU^{*}-\mathcal{C}_{\textup{pos}}^{-1})\mathcal{C}_{\textup{pos}}^{1/2}\subset g(U).$
As the domain of $\mathcal{C}_{\textup{pos}}^{1/2}(UU^{*}-H)\mathcal{C}_{\textup{pos}}^{1/2}$ is in fact all of $\mathcal{H}$, by continuity of $g(U)$ it holds that
$\mathcal{C}_{\textup{pos}}^{1/2}(UU^{*}-H)\mathcal{C}_{\textup{pos}}^{1/2}=g(U)=\sum_{i}\gamma e_{i}\otimes e_{i}.$
Multiplying this equation from the right and using $\gamma U^{*}\mathcal{C}_{\textup{pos}}^{1/2}e_{i}=0$ for all $i$,
$\mathcal{C}_{\textup{pos}}^{1/2}(UU^{*}-H)\mathcal{C}_{\textup{pos}}UU^{*}\mathcal{C}_{\textup{pos}}^{1/2}$ $=\left(\sum_{i}\lambda_{i}e_{i}\otimes e_{i}\right)\mathcal{C}_{\textup{pos}}^{1/2}UU^{*}\mathcal{C}_{\textup{pos}}^{1/2}$
$=\left(\sum_{i}e_{i}\otimes(\lambda_{i}U^{*}\mathcal{C}_{\textup{pos}}^{1/2}e_{i})\right)U^{*}\mathcal{C}_{\textup{pos}}^{1/2}=0.$
Thus, if $U\in\mathcal{B}(\mathbb{R}^{r},\mathcal{H})$ is a stationary point of $J_{f}$, then
$(\mathcal{C}_{\textup{pos}}^{1/2}H\mathcal{C}_{\textup{pos}}^{1/2})(\mathcal{C}_{\textup{pos}}^{1/2}UU^{*}\mathcal{C}_{\textup{pos}}^{1/2})=(\mathcal{C}_{\textup{pos}}^{1/2}UU^{*}\mathcal{C}_{\textup{pos}}^{1/2})^{2}.$ (45)
By injectivity of $\mathcal{C}_{\textup{pos}}^{1/2}$ and Lemma A.1, we have $\operatorname{rank}\left({\mathcal{C}_{\textup{pos}}^{1/2}UU^{*}\mathcal{C}_{\textup{pos}}^{1/2}}\right)=\operatorname{rank}\left({UU^{*}}\right)=\operatorname{rank}\left({U}\right)$. Hence $\mathcal{C}_{\textup{pos}}^{1/2}UU^{*}\mathcal{C}_{\textup{pos}}^{1/2}$ is a non-negative and self-adjoint operator of rank at most $r$. In particular, it has $k$ many nonzero eigenvalues for some $k\leq r$. It follows from (45) that these $k$ eigenpairs $(\lambda,v)\in\mathbb{R}\setminus\{0\}\times\mathcal{H}$ of $\mathcal{C}_{\textup{pos}}^{1/2}UU^{*}\mathcal{C}_{\textup{pos}}^{1/2}$ are also eigenpairs of $\mathcal{C}_{\textup{pos}}^{1/2}H\mathcal{C}_{\textup{pos}}^{1/2}$. By Lemma 3.6, there exists an increasing sequence $(\lambda_{i})_{i}\in\ell^{2}((-1,0))$ with exactly $\operatorname{rank}\left({H}\right)$ nonzero entries and an ONB $(w_{i})_{i}$ of $\mathcal{H}$ with $(w_{i})_{i}\subset\operatorname{ran}{\mathcal{C}_{\textup{pr}}^{1/2}}$ such that
$\mathcal{C}_{\textup{pos}}^{1/2}H\mathcal{C}_{\textup{pos}}^{1/2}=\sum_{i}(-\lambda_{i})(\sqrt{1+\lambda_{i}}\mathcal{C}_{\textup{pos}}^{-1/2}\mathcal{C}_{\textup{pr}}^{1/2}w_{i})\otimes(\sqrt{1+\lambda_{i}}\mathcal{C}_{\textup{pos}}^{-1/2}\mathcal{C}_{\textup{pr}}^{1/2}w_{i}).$
It follows that there exists a set of $k$ distinct indices $\mathcal{I}\coloneqq\{i_{1},\ldots,i_{k}\}\subset\{1,\ldots,\operatorname{rank}\left({H}\right)\}$ such that
$\mathcal{C}_{\textup{pos}}^{1/2}UU^{*}\mathcal{C}_{\textup{pos}}^{1/2}$ $=\sum_{j=1}^{k}(-\lambda_{i_{j}})\sqrt{1+\lambda_{i_{j}}}\mathcal{C}_{\textup{pos}}^{-1/2}\mathcal{C}_{\textup{pr}}^{1/2}w_{i_{j}}\otimes\sqrt{1+\lambda_{i_{j}}}\mathcal{C}_{\textup{pos}}^{-1/2}\mathcal{C}_{\textup{pr}}^{1/2}w_{i_{j}}$
$=\sum_{j=1}^{k}(-\lambda_{i_{j}})\sqrt{1+\lambda_{i_{j}}}^{-1}\mathcal{C}_{\textup{pos}}^{1/2}\mathcal{C}_{\textup{pr}}^{-1/2}w_{i_{j}}\otimes\sqrt{1+\lambda_{i_{j}}}^{-1}\mathcal{C}_{\textup{pos}}^{1/2}\mathcal{C}_{\textup{pr}}^{-1/2}w_{i_{j}},$
where we use (10c) for the second equality. If $U$ is a stationary point, then $U\in\mathcal{B}(\mathbb{R}^{r},H)$ satisfies the previous equation, and hence
$U=\sum_{j=1}^{k}\sqrt{\frac{-\lambda_{i_{j}}}{1+\lambda_{i_{j}}}}\mathcal{C}_{\textup{pr}}^{-1/2}w_{i_{j}}\otimes\psi_{j}\in\mathcal{B}(\mathbb{R}^{r},\mathcal{V}),$
for some orthonormal sequence $(\psi_{1},\ldots,\psi_{k})\subset\mathbb{R}^{r}$, with,
$\mathcal{V}:$ $=\operatorname{span}{\left(\mathcal{C}_{\textup{pr}}^{-1/2}w_{i},i=1,\ldots,\operatorname{rank}\left({H}\right)\right)}\subset\mathcal{H}.$
By (10c) and the relation $v_{i}=\sqrt{1+\lambda_{i}}\mathcal{C}_{\textup{pos}}^{-1/2}\mathcal{C}_{\textup{pr}}^{1/2}w_{i}$ given in Lemma 3.6, it holds for all $i$ that $v_{i}=\sqrt{1+\lambda_{i}}^{-1}\mathcal{C}_{\textup{pos}}^{1/2}\mathcal{C}_{\textup{pr}}^{-1/2}$. Thus, by (18), (10b) and the expansion of $\mathcal{C}_{\textup{pos}}^{1/2}UU^{*}\mathcal{C}_{\textup{pos}}^{1/2}$ given above,
$g(U)$ $=(\mathcal{C}_{\textup{pr}}^{-1/2}\mathcal{C}_{\textup{pos}}^{1/2})^{*}\mathcal{C}_{\textup{pr}}^{-1/2}\mathcal{C}_{\textup{pos}}^{1/2}-I+\mathcal{C}_{\textup{pos}}^{1/2}UU^{*}\mathcal{C}_{\textup{pos}}^{1/2}$
$=\sum_{i}\lambda_{i}\frac{\mathcal{C}_{\textup{pos}}^{1/2}\mathcal{C}_{\textup{pr}}^{-1/2}w_{i}}{\sqrt{1+\lambda_{i}}}\otimes\frac{\mathcal{C}_{\textup{pos}}^{1/2}\mathcal{C}_{\textup{pr}}^{-1/2}w_{i}}{\sqrt{1+\lambda_{i}}}+\sum_{j=1}^{k}(-\lambda_{i_{j}})\frac{\mathcal{C}_{\textup{pos}}^{1/2}\mathcal{C}_{\textup{pr}}^{-1/2}w_{i_{j}}}{\sqrt{1+\lambda_{i_{j}}}}\otimes\frac{\mathcal{C}_{\textup{pos}}^{1/2}\mathcal{C}_{\textup{pr}}^{-1/2}w_{i_{j}}}{\sqrt{1+\lambda_{i_{j}}}}.$
Thus, if $U$ is a stationary point of $J_{f}$, then the eigenvalues of $g(U)$ form the set $\{0\}\cup\{\lambda_{i}:\ i\not\in\mathcal{I}\}\subset(-1,0]$. Since $f\in\mathscr{F}$, it holds that $f(0)=0$, and it then follows from (19) that $J_{f}(U)=\sum_{i\not\in\mathcal{I}}f(\lambda_{i})$. Furthermore, $f$ is decreasing on $(-1,0]$, from which we conclude that for $\hat{U}$ to be a minimizer of $J_{f}$ among the stationary points of $J_{f}$, the $(\lambda_{i})_{i\in I}$ should be chosen as the $r$ most negative elements of $(\lambda_{i})_{i}$. That is, we take $\mathcal{I}=\mathcal{I}^{\textup{opt}}$ where $\mathcal{I}^{\textup{opt}}$ is any index set for which $\{\lambda_{i}:\ i\in\mathcal{I}^{\textup{opt}}\}=\{\lambda_{1},\ldots,\lambda_{r}\}$. The set $\mathcal{I}^{\textup{opt}}$ is uniquely defined if and only if $\lambda_{r}<\lambda_{r+1}$, in which case $\mathcal{I}^{\textup{opt}}=\{1,\ldots,r\}$. For such $\hat{U}$ we have
$\hat{U}\hat{U}^{*}=\sum_{i=1}^{r}\frac{-\lambda_{i}}{1+\lambda_{i}}\mathcal{C}_{\textup{pr}}^{-1/2}w_{i}\otimes\mathcal{C}_{\textup{pr}}^{-1/2}w_{i},$ (46)
and $J_{f}(\hat{U})=\sum_{i>r}f(\lambda_{i})$. If $\lambda_{r}=0$, then $\hat{U}$ and $\hat{U}\hat{U}^{*}$ are uniquely defined. If $\lambda_{i}\not=0$, then $\hat{U}$ and $\hat{U}\hat{U}^{*}$ are uniquely defined if and only if $\mathcal{I}^{\textup{opt}}$ is uniquely defined. Hence $\hat{U}\hat{U}^{*}$ is unique if and only if either $\lambda_{r}=0$ or $\lambda_{r}<\lambda_{r+1}$.
By Lemma 4.16, $\hat{U}$ minimizes $J_{f}$ over $\mathcal{B}(\mathbb{R}^{r},\mathcal{V})$. Furthermore, if $\tilde{U}\in\mathcal{B}(\mathbb{R}^{r},\mathcal{H})$ with $\operatorname{ran}{\tilde{U}}\not\subset\mathcal{V}$, then $\tilde{U}$ is not a stationary point of $J_{f}$, because a necessary condition for $\tilde{U}$ to be a stationary point of $J_{f}$ is that $\operatorname{ran}{\tilde{U}}\subset\mathcal{V}$. On the other hand, since $\hat{U}\in\mathcal{B}(\mathbb{R}^{r},\mathcal{V})\subset\mathcal{B}(\mathbb{R}^{r},\mathcal{V}+\operatorname{ran}{\tilde{U}})$, since $\tilde{U}\in\mathcal{B}(\mathbb{R}^{r},\mathcal{V}+\operatorname{ran}{\tilde{U}})$ and since $J_{f}$ is coercive over $\mathcal{B}(\mathbb{R}^{r},\mathcal{V}+\operatorname{ran}{\tilde{U}})$ by Lemma 4.16, it follows that $J_{f}(\tilde{U})>J_{f}(\hat{U})$. Thus, $\hat{U}$ is a global minimiser of $J_{f}$. Since $J_{f}(\hat{U})=\mathcal{L}_{f}(\mathcal{C}_{\textup{pos}},(\mathcal{C}_{\textup{pr}}^{-1}+\hat{U}\hat{U}^{*})^{-1})$ by (19), it follows that the operator $\mathcal{P}^{\textup{opt}}_{r}$ in (20) solves 4.4.
By Lemma 4.5(ii), $(\mathcal{P}^{\textup{opt}}_{r})^{-1}$ solves 4.3. It remains to show that $\mathcal{C}^{\textup{opt}}_{r}$ defined in (21) satisfies $\mathcal{C}^{\textup{opt}}_{r}=(\mathcal{P}^{\textup{opt}}_{r})^{-1}$. Since taking the inverse is a bijective operation, uniqueness of $\mathcal{C}^{\textup{opt}}_{r}$ is then implied by uniqueness of $\mathcal{P}^{\textup{opt}}_{r}$. Now, by Lemma A.23 with $\delta_{i}\leftarrow\lambda_{i}$,
$(\mathcal{P}^{\textup{opt}}_{r})^{-1}$ $=\left(\mathcal{C}_{\textup{pr}}^{-1}+\sum_{i=1}^{r}\frac{-\lambda_{i}}{1+\lambda_{i}}\mathcal{C}_{\textup{pr}}^{-1/2}w_{i}\otimes\mathcal{C}_{\textup{pr}}^{-1/2}w_{i}\right)^{-1}=\left(\mathcal{C}_{\textup{pr}}^{-1/2}(I-\sum_{i=1}^{r}\frac{-\lambda_{i}}{1+\lambda_{i}}w_{i}\otimes w_{i})\mathcal{C}_{\textup{pr}}^{-1/2}\right)^{-1}$
$=\mathcal{C}_{\textup{pr}}^{1/2}(I+\sum_{i=1}^{r}\lambda_{i}w_{i}\otimes w_{i})\mathcal{C}_{\textup{pr}}^{1/2}=\mathcal{C}_{\textup{pr}}-\sum_{i=1}^{r}(-\lambda_{i})\mathcal{C}_{\textup{pr}}^{1/2}w_{i}\otimes\mathcal{C}_{\textup{pr}}^{1/2}w_{i}=\mathcal{C}^{\textup{opt}}_{r}.$
B.3 Proofs of Section 5
See 5.2
Proof of Lemma 5.2.
Since $\mathcal{C}_{\textup{obs}}^{1/2}$ has a bounded inverse, Lemma A.13 and (24) imply
$\mathcal{C}_{\textup{obs}}^{-1/2}G^{*}\mathcal{C}_{\textup{pr}}G\mathcal{C}_{\textup{obs}}^{-1/2}=(\mathcal{C}_{\textup{pr}}^{1/2}G^{*}\mathcal{C}_{\textup{obs}}^{-1/2})^{*}(\mathcal{C}_{\textup{pr}}^{1/2}G^{*}\mathcal{C}_{\textup{obs}}^{-1/2})=\sum_{i=1}^{n}\frac{-\lambda_{i}}{1+\lambda_{i}}\varphi_{i}\otimes\varphi_{i}.$
Therefore, using the definitions of $S_{\textup{y}}$ and $\mathcal{C}_{\textup{y}}$, we have
$\mathcal{C}_{\textup{y}}=\mathcal{C}_{\textup{obs}}+G^{*}\mathcal{C}_{\textup{pr}}G=\mathcal{C}_{\textup{obs}}^{1/2}(I+\mathcal{C}_{\textup{obs}}^{-1/2}G^{*}\mathcal{C}_{\textup{pr}}G\mathcal{C}_{\textup{obs}}^{-1/2})\mathcal{C}_{\textup{obs}}^{1/2}=S_{\textup{y}}S_{\textup{y}}^{*}.$
Because $S_{\textup{y}}$ is a rank-$n$ operator on $\mathbb{R}^{n}$, it has a bounded inverse. By (23), (2) and (3c), it holds on $\operatorname{ran}{\mathcal{C}_{\textup{pos}}}$,
$\mathcal{C}_{\textup{pos}}^{-1}=\mathcal{C}_{\textup{pr}}^{-1}+H=\mathcal{C}_{\textup{pr}}^{-1/2}(I+\mathcal{C}_{\textup{pr}}^{1/2}H\mathcal{C}_{\textup{pr}}^{1/2})\mathcal{C}_{\textup{pr}}^{-1/2}=\mathcal{C}_{\textup{pr}}^{1/2}(I+\sum_{i=1}^{n}\frac{-\lambda_{i}}{1+\lambda_{i}}w_{i}\otimes w_{i})\mathcal{C}_{\textup{pr}}^{1/2}=(S_{\textup{pos}}S_{\textup{pos}}^{*})^{-1}.$
Finally, $I+\sum_{i}\frac{-\lambda_{i}}{1+\lambda_{i}}w_{i}\otimes w_{i}$ is boundedly invertible, hence $\operatorname{ran}{S_{\textup{pos}}}=\operatorname{ran}{\mathcal{C}_{\textup{pr}}^{1/2}}$, Because $S_{\textup{pos}}$ is also an injective operator $\mathcal{H}\rightarrow\operatorname{ran}{\mathcal{C}_{\textup{pr}}^{1/2}}$, its inverse exists. ∎
See 5.4
Proof of Lemma 5.4.
Recall that $\mathcal{C}_{\textup{pos}}=S_{\textup{pos}}S_{\textup{pos}}^{*}$ and from (3a) that $m_{\textup{pos}}(y)=\mathcal{C}_{\textup{pos}}G^{*}\mathcal{C}_{\textup{obs}}^{-1}y$ for $y\in\mathbb{R}^{n}$. Thus if we let $Z\coloneqq\widetilde{A}-S_{\textup{pos}}^{*}G^{*}\mathcal{C}_{\textup{obs}}^{-1}$, then $\widetilde{A}Y-S_{\textup{pos}}^{-1}m_{\textup{pos}}(Y)=ZY$. By Lemma A.22, the covariance of $ZY$ is $Z\mathcal{C}_{\textup{y}}Z^{*}$. Then, by applying Lemma A.21 with $X\leftarrow ZY$, $\mathcal{C}\leftarrow Z\mathcal{C}_{\textup{y}}Z^{*}$, and $S\leftarrow ZS_{\textup{y}}$,
$\mathbb{E}\left\lVert\widetilde{A}Y-S_{\textup{pos}}^{-1}m_{\textup{pos}}(Y)\right\rVert^{2}=\lVert ZS_{\textup{y}}\rVert_{L_{2}(\mathcal{H})}^{2}=\lVert\widetilde{A}S_{\textup{y}}-S_{\textup{pos}}^{*}G^{*}\mathcal{C}_{\textup{obs}}^{-1}S_{\textup{y}}\rVert_{L_{2}(\mathcal{H})}^{2}.$
Thus, to show (28) it remains to show $\mathcal{C}_{\textup{pr}}^{1/2}G^{*}\mathcal{C}_{\textup{obs}}^{-1/2}=S_{\textup{pos}}^{*}G^{*}\mathcal{C}_{\textup{obs}}^{-1}S_{\textup{y}}$. By (25),
$S_{\textup{pos}}^{*}G^{*}\mathcal{C}_{\textup{obs}}^{-1}S_{\textup{y}}$ $=\left(I+\sum_{i=1}^{n}\frac{-\lambda_{i}}{1+\lambda_{i}}w_{i}\otimes w_{i}\right)^{-1/2}\mathcal{C}_{\textup{pr}}^{1/2}G^{*}\mathcal{C}_{\textup{obs}}^{-1/2}\left(I+\sum_{i=1}^{n}\frac{-\lambda_{i}}{1+\lambda_{i}}\varphi_{i}\otimes\varphi_{i}\right)^{1/2}.$
Fix an arbitrary $x\in\mathbb{R}^{n}$. Then,
$S_{\textup{pos}}^{*}G^{*}\mathcal{C}_{\textup{obs}}^{-1}S_{\textup{y}}x$ $=\left(I+\sum_{i=1}^{n}\frac{-\lambda_{i}}{1+\lambda_{i}}w_{i}\otimes w_{i}\right)^{-1/2}\left(\sum_{i}^{n}\sqrt{\frac{-\lambda_{i}}{1+\lambda_{i}}}w_{i}\otimes\varphi_{i}\right)\sum_{i=1}^{n}(1+\lambda_{i})^{-1/2}\langle x,\varphi_{i}\rangle\varphi_{i}$
$=\left(I+\sum_{i=1}^{n}\frac{-\lambda_{i}}{1+\lambda_{i}}w_{i}\otimes w_{i}\right)^{-1/2}\sum_{i=1}^{n}\sqrt{\frac{-\lambda_{i}}{(1+\lambda_{i})^{2}}}\langle x,\varphi_{i}\rangle w_{i}$
$=\sum_{i=1}^{n}\sqrt{\frac{-\lambda_{i}}{1+\lambda_{i}}}\langle x,\varphi_{i}\rangle w_{i}$
$=\left(\sum_{i=1}^{n}\sqrt{\frac{-\lambda_{i}}{1+\lambda_{i}}}w_{i}\otimes\varphi_{i}\right)x=\mathcal{C}_{\textup{pr}}^{1/2}G^{*}\mathcal{C}_{\textup{obs}}^{-1/2}x,$
where we use (24), (25) and (26b) in the first equation, (26a) in the third equation and (24) in the last equation. ∎
See 5.6
Proof of Theorem 5.6.
By Lemma 5.4, it suffices to first find $\widetilde{A}^{\textup{opt},(2)}$ that solves the rank-constrained operator approximation problem
$\min\left\{\left\lVert\widetilde{A}S_{\textup{y}}-\mathcal{C}_{\textup{pr}}^{1/2}G^{*}\mathcal{C}_{\textup{pos}}^{-1/2}\right\rVert^{2}_{L_{2}(\mathcal{H})},\ \widetilde{A}\in\widetilde{\mathscr{M}}^{(2)}_{r}=\mathcal{B}_{00,r}(\mathbb{R}^{r},\mathcal{H})\right\}$
and then set $A_{r}^{\textup{opt},(2)}\coloneqq S_{\textup{pos}}\widetilde{A}_{r}^{\textup{opt},(2)}$. Note that $I^{\dagger}=I$, $S_{y}^{\dagger}=S_{\textup{y}}^{-1}$ and, by (24), $(\mathcal{C}_{\textup{pr}}^{1/2}G^{*}\mathcal{C}_{\textup{pos}}^{-1/2})_{r}=\sum_{i=1}^{r}\sqrt{\frac{-\lambda_{i}}{1+\lambda_{i}}}w_{i}\otimes\varphi_{i}$ is a rank-$r$ truncated SVD of $\mathcal{C}_{\textup{pr}}^{1/2}G^{*}\mathcal{C}_{\textup{pos}}^{-1/2}$. Since $I\in\mathcal{B}(\mathcal{H})$ and $S_{\textup{y}}\in\mathcal{B}(\mathbb{R}^{n})$ have closed range and since $\mathcal{C}_{\textup{pr}}^{1/2}G^{*}\mathcal{C}_{\textup{pos}}^{-1/2}$ has finite rank, hence is Hilbert–Schmidt, we may apply Theorem 5.5 with $\mathcal{H}_{i}\leftarrow\mathbb{R}^{n}$ for $i\in\{1,2\}$, $\mathcal{H}_{i}\leftarrow\mathcal{H}$ for $i\in\{3,4\}$, $T\leftarrow I$, $S\leftarrow S_{\textup{y}}$, $M\leftarrow\mathcal{C}_{\textup{pr}}^{1/2}G^{*}\mathcal{C}_{\textup{pos}}^{-1/2}$ to find
$\widetilde{A}_{r}^{\textup{opt},(2)}$ $=\left(\sum_{i=1}^{r}\sqrt{\frac{-\lambda_{i}}{1+\lambda_{i}}}w_{i}\otimes\varphi_{i}\right)S_{\textup{y}}^{-1},$
and $\operatorname{ran}{{A}_{r}^{\textup{opt},(2)}}=\operatorname{ran}{S_{\textup{pos}}\widetilde{A}_{r}^{\textup{opt},(2)}}\subset\operatorname{span}{\left(S_{\textup{pos}}w_{i},\ i\leq r\right)}\subset\operatorname{ran}{C}_{\textup{pr}}$. Thus,
$A_{r}^{\textup{opt},(2)}$ $=S_{\textup{pos}}\widetilde{A}_{r}^{\textup{opt},(2)}=S_{\textup{pos}}\left(\sum_{i=1}^{r}\sqrt{\frac{-\lambda_{i}}{1+\lambda_{i}}}w_{i}\otimes\varphi_{i}\right)S_{\textup{y}}^{-1}$
$=\mathcal{C}_{\textup{pr}}^{1/2}\left(I+\sum_{i=1}^{n}\frac{-\lambda_{i}}{1+\lambda_{i}}w_{i}\otimes w_{i}\right)^{-1/2}\sum_{i=1}^{r}\sqrt{\frac{-\lambda_{i}}{1+\lambda_{i}}}w_{i}\otimes\varphi_{i}\left(I+\sum_{i=1}^{n}\frac{-\lambda_{i}}{I+\lambda_{i}}\varphi_{i}\otimes\varphi_{i}\right)^{-1/2}\mathcal{C}_{\textup{obs}}^{-1/2}$
$=\mathcal{C}_{\textup{pr}}^{1/2}\left(\sum_{i=1}^{r}\sqrt{-\lambda_{i}(1+\lambda_{i})}w_{i}\otimes\varphi_{i}\right)\mathcal{C}_{\textup{obs}}^{-1/2},$
where we used (25) in the third equation and (26) in the last equation. Using (24), the definition of the Hilbert–Schmidt norm and the definition of $\widetilde{A}_{r}^{\textup{opt},(2)}$, we can compute the corresponding minimal loss:
$\left\lVert\widetilde{A}_{r}^{\textup{opt},(2)}S_{\textup{y}}-\mathcal{C}_{\textup{pr}}^{1/2}G^{*}\mathcal{C}_{\textup{pos}}^{-1/2}\right\rVert_{L_{2}(\mathcal{H})}^{2}$ $=\left\lVert\sum_{i=1}^{r}\sqrt{\frac{-\lambda_{i}}{1+\lambda_{i}}}w_{i}\otimes\varphi_{i}-\sum_{i=1}^{n}\sqrt{\frac{-\lambda_{i}}{1+\lambda_{i}}}w_{i}\otimes\varphi_{i}\right\rVert_{L_{2}(\mathcal{H})}^{2}=\sum_{i>r}\frac{-\lambda_{i}}{1+\lambda_{i}}.$
See 5.7
Proof of Theorem 5.7.
By Lemma 5.4, it suffices to first find $\widetilde{A}^{\textup{opt},(1)}_{r}$ that solves the rank-constrained operator approximation problem
$\min\left\{\left\lVert\widetilde{A}S_{\textup{y}}-\mathcal{C}_{\textup{pr}}^{1/2}G^{*}\mathcal{C}_{\textup{pos}}^{-1/2}\right\rVert^{2}_{L_{2}(\mathcal{H})},\ \widetilde{A}\in\widetilde{\mathscr{M}}^{(1)}_{r}\right\},$
and then set $A_{r}^{\textup{opt},(1)}\coloneqq S_{\textup{pos}}\widetilde{A}^{\textup{opt},(1)}_{r}$. Recall that by definition, $\widetilde{A}\in\widetilde{\mathscr{M}}_{r}^{(1)}$ if and only if $\widetilde{A}=(S_{\textup{pos}}^{-1}\mathcal{C}_{\textup{pr}}-\widetilde{B})G^{*}\mathcal{C}_{\textup{obs}}^{-1}$ for some $\widetilde{B}\in\mathcal{B}_{00,r}(\mathcal{H})$. Notice that for such $\widetilde{A}$,
$\widetilde{A}S_{\textup{y}}-\mathcal{C}_{\textup{pr}}^{1/2}G^{*}\mathcal{C}_{\textup{obs}}^{-1/2}=S_{\textup{pos}}^{-1}\mathcal{C}_{\textup{pr}}G^{*}\mathcal{C}_{\textup{obs}}^{-1}S_{\textup{y}}-\mathcal{C}_{\textup{pr}}^{1/2}G^{*}\mathcal{C}_{\textup{obs}}^{-1/2}-\widetilde{B}G^{*}\mathcal{C}_{\textup{obs}}^{-1}S_{\textup{y}}.$
The above rank-$r$ operator approximation problem is therefore equivalent to the following rank-$r$ operator approximation problem:
$\min\left\{\lVert S_{\textup{pos}}^{-1}\mathcal{C}_{\textup{pr}}G^{*}\mathcal{C}_{\textup{obs}}^{-1}S_{\textup{y}}-\mathcal{C}_{\textup{pr}}^{1/2}G^{*}\mathcal{C}_{\textup{obs}}^{-1/2}-\widetilde{B}G^{*}\mathcal{C}_{\textup{obs}}^{-1}S_{\textup{y}}\rVert_{L_{2}(\mathcal{H})}\ :\ \widetilde{B}\in\mathcal{B}_{00,r}(\mathcal{H})\right\}.$
Since $I\in\mathcal{B}(\mathcal{H})$ and $G^{*}\mathcal{C}_{\textup{obs}}^{-1}S_{\textup{y}}$ have closed range and since $S_{\textup{pos}}^{-1}\mathcal{C}_{\textup{pr}}G^{*}\mathcal{C}_{\textup{obs}}^{-1}S_{\textup{y}}-\mathcal{C}_{\textup{pr}}^{1/2}G^{*}\mathcal{C}_{\textup{obs}}^{-1/2}$ has finite rank and therefore is Hilbert–Schmidt, we may apply Theorem 5.5 with $\mathcal{H}_{1}\leftarrow\mathbb{R}^{n}$ and $\mathcal{H}_{j}\leftarrow\mathcal{H}$ for $j\in\{2,3,4\}$, $T\leftarrow I$, $S\leftarrow G^{*}\mathcal{C}_{\textup{obs}}^{-1}S_{\textup{y}}$ and $M\leftarrow S_{\textup{pos}}^{-1}\mathcal{C}_{\textup{pr}}G^{*}\mathcal{C}_{\textup{obs}}^{-1}S_{\textup{y}}-\mathcal{C}_{\textup{pr}}^{1/2}G^{*}\mathcal{C}_{\textup{obs}}^{-1/2}$ to find a solution $\widetilde{B}^{\textup{opt}}$ to the approximation problem above. For the given choices of $T$ and $S$, we have that $T^{\dagger}=I$, while for the finite rank operator $S$ we have from (24) and (25) that
$S$ $=\mathcal{C}_{\textup{pr}}^{-1/2}\left(\sum_{i=1}^{n}\sqrt{\frac{-\lambda_{i}}{1+\lambda_{i}}}w_{i}\otimes\varphi_{i}\right)\left(I+\sum_{i=1}^{n}\frac{-\lambda_{i}}{1+\lambda_{i}}\varphi_{i}\otimes\varphi_{i}\right)^{1/2}.$
By [13, Theorem 2.8], the Moore–Penrose inverse of $\sum_{i=1}^{n}\sqrt{\frac{-\lambda_{i}}{1+\lambda_{i}}}w_{i}\otimes\varphi_{i}$ is given by $\sum_{i=1}^{n}\sqrt{\frac{1+\lambda_{i}}{-\lambda_{i}}}\varphi_{i}\otimes w_{i}$. Furthermore, the Moore–Penrose inverse of a composition of bounded operators is the composition in reverse order of the Moore–Penrose inverses of these operators, e.g. [18, Equation 3.23]. Since $\mathcal{C}_{\textup{pr}}^{-1/2}$ and $I+\sum_{i=1}^{n}\frac{-\lambda_{i}}{1+\lambda_{i}}\varphi_{i}\otimes\varphi_{i}$ are boundedly invertible, we thus compute
$S^{\dagger}=\left(I+\sum_{i=1}^{n}\frac{-\lambda_{i}}{1+\lambda_{i}}\varphi_{i}\otimes\varphi_{i}\right)^{-1/2}\left(\sum_{i=1}^{n}\sqrt{\frac{1+\lambda_{i}}{-\lambda_{i}}}\varphi_{i}\otimes w_{i}\right)\mathcal{C}_{\textup{pr}}^{1/2}$
Furthermore, by [13, Equation (2.12)], $P_{\ker{S}^{\perp}}=S^{\dagger}S$, showing that $P_{\ker{S}^{\perp}}=\sum_{i=1}^{n}\varphi_{i}\otimes\varphi_{i}.$ Next, we compute for the given choice of $M$,
$M$ $=S_{\textup{pos}}^{-1}\mathcal{C}_{\textup{pr}}^{1/2}\left(\mathcal{C}_{\textup{pr}}^{1/2}G^{*}\mathcal{C}_{\textup{obs}}^{-1/2}\right)\mathcal{C}_{\textup{obs}}^{-1/2}S_{\textup{y}}-\mathcal{C}_{\textup{pr}}^{1/2}G^{*}\mathcal{C}_{\textup{obs}}^{-1/2}$
$=\left(I+\sum_{i=1}^{n}\frac{-\lambda_{i}}{1+\lambda_{i}}w_{i}\otimes w_{i}\right)^{1/2}\left(\sum_{i=1}^{n}\sqrt{\frac{-\lambda_{i}}{1+\lambda_{i}}}w_{i}\otimes\varphi_{i}\right)\left(I+\sum_{i=1}^{n}\frac{-\lambda_{i}}{1+\lambda_{i}}\varphi_{i}\otimes\varphi_{i}\right)^{1/2}$
$-\sum_{i=1}^{r}\sqrt{\frac{-\lambda_{i}}{1+\lambda_{i}}}w_{i}\otimes\varphi_{i}$
$=\sum_{i=1}^{n}\left(\sqrt{\frac{-\lambda_{i}}{(1+\lambda_{i})^{3}}}-\sqrt{\frac{-\lambda_{i}}{1+\lambda_{i}}}\right)w_{i}\otimes\varphi_{i},$
where in the second equation we use (24) and (25), and in the last equation we use (26). Hence, $MP_{\ker{S}^{\perp}}=M$ and Theorem 5.5 yields, with $(M)_{r}$ a rank-$r$ truncated SVD of $M$,
$\widetilde{B}^{\textup{opt}}$ $=B^{\dagger}(M)_{r}C^{\dagger}=\left(\sum_{i=1}^{r}\left(\sqrt{\frac{-\lambda_{i}}{(1+\lambda_{i})^{3}}}-\sqrt{\frac{-\lambda_{i}}{1+\lambda_{i}}}\right)w_{i}\otimes\varphi_{i}\right)S^{\dagger}$
$=\left(\sum_{i=1}^{r}\left(\sqrt{\frac{-\lambda_{i}}{(1+\lambda_{i})^{2}}}-\sqrt{-\lambda_{i}}\right)w_{i}\otimes\varphi_{i}\right)\left(\sum_{i=1}^{n}\sqrt{\frac{1+\lambda_{i}}{-\lambda_{i}}}\varphi_{i}\otimes w_{i}\right)\mathcal{C}_{\textup{pr}}^{1/2}$
$=\left(\sum_{i=1}^{r}\left(\sqrt{\frac{1}{1+\lambda_{i}}}-\sqrt{1+\lambda_{i}}\right)w_{i}\otimes w_{i}\right)\mathcal{C}_{\textup{pr}}^{1/2},$
as can be verified by application to an arbitrary vector $h\in\mathcal{H}$ and direct computation. It follows by (25), (26a) and direct computation, that
$S_{\textup{pos}}\widetilde{B}^{\textup{opt}}$ $=\mathcal{C}_{\textup{pr}}^{1/2}\left(I+\sum_{i\in\mathbb{N}}\frac{-\lambda_{i}}{1+\lambda_{i}}w_{i}\otimes w_{i}\right)^{-1/2}\left(\sum_{i=1}^{r}\left(\sqrt{\frac{1}{1+\lambda_{i}}}-\sqrt{1+\lambda_{i}}\right)w_{i}\otimes w_{i}\right)\mathcal{C}_{\textup{pr}}^{1/2}$
$=\mathcal{C}_{\textup{pr}}^{1/2}\left(\sum_{i=1}^{r}\left(1-(1+\lambda_{i})\right)w_{i}\otimes w_{i}\right)\mathcal{C}_{\textup{pr}}^{1/2}$
$=\sum_{i=1}^{r}(-\lambda_{i})\mathcal{C}_{\textup{pr}}^{1/2}w_{i}\otimes\mathcal{C}_{\textup{pr}}^{1/2}w_{i}.$
Thus, using the expression for $\mathcal{C}^{\textup{opt}}_{r}$ in (21) from Theorem 4.18,
$A_{r}^{\textup{opt},(1)}=S_{\textup{pos}}\widetilde{A}_{r}^{\textup{opt},(1)}=S_{\textup{pos}}(S_{\textup{pos}}^{-1}\mathcal{C}_{\textup{pr}}-\widetilde{B}^{\textup{opt}})G^{*}\mathcal{C}_{\textup{obs}}^{-1}=(\mathcal{C}_{\textup{pr}}-S_{\textup{pos}}\widetilde{B}^{\textup{opt}})G^{*}\mathcal{C}_{\textup{obs}}^{-1}=\mathcal{C}^{\textup{opt}}_{r}G^{*}\mathcal{C}_{\textup{obs}}^{-1}.$
We note that $\operatorname{ran}{A_{r}^{\textup{opt},(1)}}\subset\operatorname{ran}{\mathcal{C}_{r}^{\textup{opt}}}\subset\operatorname{span}{\left(\mathcal{C}_{\textup{pr}}^{1/2}w_{i},\ i\leq n\right)}\subset\operatorname{ran}{\mathcal{C}_{\textup{pr}}}$. Finally, we compute the corresponding loss. By (21) and (24),
$\mathcal{C}_{\textup{pr}}^{-1/2}\mathcal{C}^{\textup{opt}}_{r}G^{*}\mathcal{C}_{\textup{obs}}^{-1/2}$ $=\left(I-\sum_{i=1}^{r}(-\lambda_{i})w_{i}\otimes w_{i}\right)\mathcal{C}_{\textup{pr}}^{1/2}G^{*}\mathcal{C}_{\textup{obs}}^{-1/2}$
$=\left(I-\sum_{i=1}^{r}(-\lambda_{i})w_{i}\otimes w_{i}\right)\sum_{i=1}^{n}\sqrt{\frac{-\lambda_{i}}{1+\lambda_{i}}}w_{i}\otimes\varphi_{i},$
Together with (25), the preceiding equation implies that
$S_{\textup{pos}}^{-1}A_{r}^{\textup{opt},(1)}S_{\textup{y}}$ $=\left(I+\sum_{i=1}^{n}\frac{-\lambda_{i}}{1+\lambda_{i}}w_{i}\otimes w_{i}\right)^{1/2}\mathcal{C}_{\textup{pr}}^{-1/2}\mathcal{C}^{\textup{opt}}_{r}G^{*}\mathcal{C}_{\textup{obs}}^{-1/2}\left(I+\sum_{i=1}^{n}\frac{-\lambda_{i}}{1+\lambda_{i}}\varphi_{i}\otimes\varphi_{i}\right)^{1/2}$
$=\sum_{i=1}^{n}\sqrt{\frac{-\lambda_{i}}{(1+\lambda_{i})^{3}}}w_{i}\otimes\varphi_{i}-\sum_{i=1}^{r}\sqrt{\frac{-\lambda_{i}}{1+\lambda_{i}}}^{3}w_{i}\otimes\varphi_{i}.$
Since $\sqrt{\frac{-\lambda_{i}}{(1+\lambda_{i})^{3}}}=\sqrt{\frac{-\lambda_{i}}{1+\lambda_{i}}}\left(1+\frac{-\lambda_{i}}{1+\lambda_{i}}\right)$,
$\widetilde{A}_{r}^{\textup{opt},(1)}S_{\textup{y}}=S_{\textup{pos}}^{-1}A_{r}^{\textup{opt},(1)}S_{y}$ $=\sum_{i>r}\sqrt{\frac{-\lambda_{i}}{1+\lambda_{i}}}^{3}w_{i}\otimes\varphi_{i}+\sum_{i=1}^{n}\sqrt{\frac{-\lambda_{i}}{1+\lambda_{i}}}w_{i}\otimes\varphi_{i}$
$=\sum_{i>r}\sqrt{\frac{-\lambda_{i}}{1+\lambda_{i}}}^{3}w_{i}\otimes\varphi_{i}+\mathcal{C}_{\textup{pr}}^{1/2}G^{*}\mathcal{C}_{\textup{obs}}^{-1/2},$
where the last equation follows from (24). We conclude, by definition of the Hilbert–Schmidt norm,
$\lVert\widetilde{A}^{\textup{opt},(1)}_{r}S_{\textup{y}}-\mathcal{C}_{\textup{pr}}^{1/2}G^{*}\mathcal{C}_{\textup{obs}}^{-1/2}\rVert_{L_{2}(\mathcal{H})}^{2}=\sum_{i>r}\sqrt{\frac{-\lambda_{i}}{1+\lambda_{i}}}^{6}$
which completes the proof of Theorem 5.7. ∎